A Complete Bibliography of Publications in the SIAM Journal on Applied Mathematics: 2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

13 October 2017 Version 3.16

Title word cross-reference

1 [172]. 1/f [628]. 1: 2 [172]. 2 [544, 881, 685, 367, 557]. 2 + 1 [910]. 3 [561, 267]. _{2} [987]. B_{z} [734]. β [700]. D [718]. hp [669]. k [888, 889]. L^{1} [555]. M/M/1 [679]. n [524]. ω [527]. p [84, 747]. r [632]. R^{d} [528]. z [965, 995].

-Analytic [632]. -Bar [718]. -Cell [700]. -D [172, 544, 557]. -Degree-of-Freedom [524]. -Harmonic [527]. -Harmonically [888, 889]. -System [84]. -Transform [965, 995].

1 [710].

2 [147, 145]. 2-D [147, 145]. 2-Dimensional [782].

8-Moment [963].

above [540]. Absence [713]. Absolute [857]. Absorbing [818, 211, 46]. Absorption [726]. Accelerating [382]. Access [769, 237]. Accuracy [834, 669]. Accurate [610, 22, 703]. Acoustic [712, 620, 166, 937, 873, 177, 320, 425, 464, 727, 334, 392, 897, 760, 597, 363, 501]. Acoustic-Rotational [501]. Acoustical [185]. Acoustics [94]. Action [214]. Activation [802, 970, 764, 593]. Active [377, 113, 217, 334, 749]. Activity [200, 201, 450, 163]. Acute [604]. Adaptive [580, 669]. Adatoms [850]. Added [5].

Addition [501]. Additive [904]. Adhesive [385]. Adiabatic [199]. Adjacent [709]. Adjoint [737, 772]. Adjusting [117]. Admissibility [786]. Admittance [848]. Adult [922]. Advancing [647]. Advection [98]. Advection-Diffusion [98]. Age [558, 21, 161, 261, 681, 710, 545]. Age-Dependent [681]. Age-Distribution [161]. Age-Structured [558, 21, 710, 545]. Aggregation [16, 103, 260]. Aggression [451]. ahead [617]. Algae [783]. Algorithm [839, 118, 278, 953, 819, 734, 291]. Algorithms [93, 152, 648]. Aligning [452]. Alignment [116]. Allee [904]. Allen [903]. Alloy [276]. Alloys [414]. Almost [7]. along [379]. Alternans [880]. Alveoli [294]. Always [143]. American [325, 583, 270, 467, 967]. Amplifications [223]. Amplifiers [596]. Amplitude [101, 221, 880]. Analyses [550, 112]. Analysis [53, 554, 451, 635, 513, 697, 573, 480, 27, 55, 34, 676, 67, 847, 341, 344, 92, 961, 543, 478, 875, 976, 580, 695, 364, 272, 932, 686, 972, 565, 134, 711, 219, 336, 704, 670, 784, 960, 169, 894, 710, 158, 209, 974, 362, 173, 42, 941, 810, 409, 345, 391, 196, 293, 204, 459, 460, 308, 411, 312]. Analytic [632, 888, 889, 942]. Analytical [318, 827, 469, 820]. Analyzing [105, 207]. Anatomically [235]. Angle [347]. Anisotropic [3, 627, 339, 267, 850, 232, 786, 942]. Anisotropy [551, 760, 159, 860]. Annihilation [62]. Annular [437]. Annulus [329]. Anomalous [198]. Anomaly [487]. Antenna [851]. Anticipation [76, 704]. Antiplane [279, 238]. Antiretroviral [710]. Aperture [347]. APEX [310]. Apparently [80]. Application [176, 887, 480, 272, 933, 350, 506, 693, 534, 843, 803]. Applications [532, 629, 207, 660, 664, 940, 703, 364, 544, 420, 559, 722, 991, 720, 500]. Applied [47, 370, 812]. Approach [805, 93, 925, 516, 408, 671, 238, 933, 971, 399, 701, 702, 685, 424, 509, 914, 784, 769, 42, 850, 535, 391]. Approximate [44, 246, 582]. Approximated [371]. Approximation [451, 83, 428, 555, 24, 298, 688, 374, 419, 934, 913, 761, 548, 463, 986, 820, 343]. Approximations [404, 473]. Arbitrarily [129, 137]. Arbitrary [26, 347, 819, 727, 709]. Area [915]. Aris [327]. Arising [480, 222, 465, 447, 436, 23, 941]. Array [699, 733, 169]. Arrays [424, 852, 846]. Arrest [361]. Arrhenius [82, 617]. Arrival [211]. Arterial [682, 407]. Articular [449]. Artificial [654]. Aspects [555, 619]. Assemblies [899]. Assessment [496]. Associated [755, 214]. Associative [855]. Asymmetric [829, 617, 781]. Asymptotic [412, 53, 735, 830, 451, 28, 55, 276, 3, 141, 290, 673, 849, 826, 202, 934, 972, 66, 134, 91, 331, 13, 692, 295, 384, 378, 302, 553, 169, 9, 974, 173, 353, 941, 77, 950, 196, 293, 538, 58, 317, 930]. Asymptotics [6, 226, 273, 944, 918, 259, 124, 362, 360]. Atomic [761, 665]. Attachment [266, 119]. Attenuation [422]. Attractors [752]. Audio [596]. Autocatalysis [859, 100, 75]. Autoimmune [719]. Automata [252]. Automaton [103]. Autosolitons [254]. Average [328]. Averaged [404]. Averaging [246, 316, 579]. Aw [223, 779, 704, 917]. Axial [265]. Axisymmetric [257]. Azimuthal [522].

B [930]. Back [512, 602].

Back-Projections [512]. Back-Scattering [602]. Background [738, 851]. Backprojection [787, 278, 776]. Backprojection-Type [278]. Backscattering [361, 837]. Backward [390, 134]. Backward-Forward [134]. Bacteria [783]. Bacteria-Algae [783]. Bacterial [492, 266, 936]. Bacteriophage [499]. Baking [770]. Balance [178, 928, 416]. Balances [625]. Ball [716]. Balloons [513]. Balls [5, 680]. Band [921]. Bandgap [466]. Bar [48, 718]. Based [844, 309, 961, 402, 787, 397, 493, 779, 442, 769, 235]. Bases [468]. Basic [635]. Bautin [357, 25]. Bayesian [453]. Beam [28, 520, 776, 251, 469]. Beavers [50]. Bed [17]. Beddington [624]. Beds [751]. Beetle [841]. Before [584]. Behavior [412, 922, 758, 655, 682, 642, 492, 342, 908, 588, 91, 90, 553, 95, 796]. Beker [255]. Beltrami [546]. Belyakov [213]. Bénard [142, 153]. Beneath [973]. Berry [578, 649]. Between [490, 518, 124, 808, 625, 907, 827, 709]. Beyond [463]. BIAcore [664]. Bifurcation [53, 554, 102, 390, 635, 639, 323, 395, 676, 344, 357, 976, 561, 544, 650, 214, 743, 90, 136, 209, 123, 852, 810, 754, 312]. Bifurcations [503, 481, 182, 213, 329, 508, 612]. Bilaterally [438]. Bilinear [282]. Binary [724, 268]. Binding [552]. Bingham [86]. Binomial [856]. Biochemical [63]. Biodegradation [115]. Biofilm [228]. Biofilms [877]. Biological [246, 932, 814, 974, 482]. Biology [926]. Biomathematical [135]. Biomedical [991]. Bioremediation [793]. Biphasic [595, 449]. Bipolar [956]. Birth [500]. Bistability [601]. Bistable [222, 167]. Blind [180, 981]. Bloch [444]. Block [108]. Blocking [577]. Blood [682, 666, 407, 235]. Blooming [901]. Blow [226, 457, 549, 870]. Blow-Up [457, 226, 549, 870]. Blown [709]. Bodies [744]. Body [192, 975, 371, 661, 816, 781]. Bogy [685]. Boltzmann [444, 629, 178, 724, 387]. Bones [570]. Booming [505]. Boring [841]. Borne [694]. Bose [610]. Bound [416, 394, 938]. Boundaries [497, 952, 15, 81]. Boundary [195, 240, 929, 140, 696, 73, 854, 233, 543, 774, 403, 544, 891, 270, 36, 155, 484, 284, 265, 436, 50, 420, 672, 591, 742, 983, 396, 654, 822, 75, 211, 46, 548, 351, 110, 42, 852, 343, 74, 346, 459, 460, 680]. Boundary-Integral [284]. Boundary-Layer [544]. Boundary-Roughness [696]. Boundary-Value [265]. Bounded [178, 438]. Bounds [989, 683, 792, 920, 495, 489]. Boussinesq [324]. Branched [117, 789]. Branched-Chain [117, 789]. Branching [970]. Bread [770]. Breakdowns [762]. Breaking [853, 568]. Breakup [971]. Breathers [567]. Bremmer [58]. Brownian [374, 33]. Brusselator [858]. Bubble [581, 516]. Bubbling [360]. Bubbly [366]. Buckling [272, 656, 972, 251]. Budaev [685]. Buffered [173, 585]. Bulk [683]. Bumps [289]. Burgers [823, 446, 486]. Burning [219]. Burning-Rate [219]. Burst [700]. Bursting [662, 909, 25]. BV [780, 368, 453].

C [885, 670, 894]. Cable [113]. Cahn [925, 903, 863, 95, 607]. Calcium [451, 519, 173]. Calculation [210, 96]. Calculations [284, 74]. Calculus [377, 780]. Calibration [608]. Calorimetry [242]. Can [478, 882, 206, 281]. Canal [740]. Canal-Neck [740]. Canard [478, 372]. Cancer [47]. Capacity [536, 384, 378]. Capillaries [817, 77]. Capillarity [84]. Capillary [273, 262]. Car [480, 417]. Carbonate [451, 958]. Cardiac [655, 880]. Caricature [305]. Carlo [924, 772]. Carrier [99]. Cart [782]. Cartilage [449]. Cascade [688, 746]. Case [639, 668, 963, 203]. Cases [133]. Catalyst [73]. Catalytic [73, 331]. Catastrophe [933]. Catastrophe-Theoretic [933]. Catenoid [939]. Cathode [345]. Cathodes [952]. Cauchy [516, 462]. Caused [501]. Caustics [122]. Cavitating [348]. Cavitation [753, 751, 609]. Cavities [409]. Cell [27, 958, 666, 47, 370, 449, 953, 814, 693, 700, 199, 899, 957, 345]. Cell-Cycle-Specific [47]. Cell-Kill [370].

Cell-Matrix [449]. Cells [956, 311, 949, 719, 948, 950, 951]. Cellular [667, 252, 128, 40, 359, 537, 750, 103]. Center [639, 561, 908, 656, 131]. Cercignani [672]. Chain [117, 789, 213, 745, 947, 970]. Chain-Branching [970]. Chains [616, 762, 441, 171, 130, 500]. Challenge [958]. Change [214, 651]. Changes [958]. Changing [234]. Channel [829, 721, 395, 556, 68, 724, 284]. Channels [723, 128, 882, 974]. Chaos [40, 110]. Chaotic [80]. Chapman [824, 785]. Characteristic [968, 981, 212]. Characterization [262, 419, 798, 851]. Characterizations [662]. Charged [853, 761]. Checkerboard [174]. Chemical [830, 747, 542, 634, 327, 976, 387, 699, 893, 112]. Chemistry [970]. Chemostat [91, 1, 674, 119, 485]. Chemostat-Type [1, 674]. Chemotactic [359, 537, 260]. Chemotaxis [492, 590, 491, 244, 511, 104, 189]. Chemotherapeutic [370]. Chemotherapy [47]. Chiral [885, 670]. Choose [945]. Chronic [532]. Chute [861]. CIR [608]. Circle [215, 205, 918]. Circuits [181]. Circular [240]. Circulating [719]. Circulation [4, 405, 601]. Clamped [305]. Clarifier [515]. Clarifier-Thickener [515]. Class [830, 845, 596, 233, 857, 299, 839, 776, 603, 146, 810, 942]. Class-D [596]. Classes [59]. Classical [575, 540, 920]. Classification [902, 149]. Cleansing [294]. Clear [263]. Clocks [246]. Clog [882]. Clogging [980]. Clogs [584]. Close [976]. Closed [59, 467, 80, 856]. Closed-Form [467, 856]. Closely [151, 398, 709]. Closure [824, 963, 9]. Closures [353]. Clouds [528]. Cluster [786]. CO [987]. Coated [410, 510, 514, 489, 661, 816]. Coatings [62]. Coaxial [827]. Cochlea [111]. Code [237]. Coefficient [820]. Coefficients [26, 798, 52]. Coevolution [344]. Coexistence [114, 603, 876]. Cognition [594]. Coiled [111]. Coinfection [614]. Collapse [361, 441]. Collective [492]. Collision [178]. Collocation [769]. Collocation-Kernel [769]. Color [546, 780]. Column [581, 488, 638]. Combination [710]. Combing [445]. Combustion [697, 984, 641, 526, 550, 338, 9, 785, 269, 286, 680]. Communication [819]. Communications [660]. Community [639]. Comparison [922, 202, 112]. Compartments [941]. Competing [370]. Competition [375, 70, 1, 674, 119, 485, 754]. Competitive [838, 686]. Complementarity [56]. Complementary [91]. Complete [857, 891, 419, 902, 476, 812]. Complex [542, 634, 683, 397, 42, 809, 832]. Complex-Valued [683]. Component [683, 792, 866]. Components [414]. Composite [749]. Composites [94, 174, 683, 792, 920, 496]. Compositional [34]. Compounds [644]. Compressible [418, 753, 881, 522, 227, 898, 293]. Compression [380]. Computation [610, 729, 924, 479, 145]. Computational [290, 949, 684]. Compute [741]. Computing [10, 239]. Concentration [6, 724, 433, 434]. Concept [15]. Concrete [927]. Condensates [610]. Condition [240, 315, 961, 36, 484, 50, 672]. Conditions [176, 543, 301, 544, 155, 654, 822, 46, 548, 691, 351, 110, 919]. Conducting [151, 92, 169]. Conductive [326]. Conductivities [718]. Conductivity [510, 527, 591, 742]. Conductor [250]. Conductors [709]. Cone [240, 26, 776]. Configuration [239]. Configurational [606]. Configurations [85, 489, 45]. Confined [900, 219, 752]. Congestion [319, 431]. Conical [988]. Conjecture [590, 491]. Connected [439]. Connecting [321]. Connection [808]. Connectivity [503, 777]. Conservation [188, 854, 69]. Conserved [416]. Consolidation [190].

Constant [798, 52, 508]. Constrained [958, 438]. Constrictions [474]. Constructing [439]. Construction [141, 120, 773]. Constructive [839]. Contact [385, 627, 118, 124]. Containing [971]. Continuation [535]. Continuous [715, 515, 717, 967, 441, 800, 442]. Continuous-Installment [967]. Continuum [297, 885, 861, 400, 654, 211, 321, 296, 910]. Contours [377]. Contrast [523]. Control [245, 958, 465, 47, 370, 841, 217, 334, 750, 749]. Convection [8, 301, 664, 931, 153, 205, 579, 960, 905]. Convection-Diffusion [8, 205, 579, 960]. Convective [774, 199, 607]. Convective-Diffusive [199]. Convergence [657, 928, 150, 819, 734, 579, 601, 609]. Convergent [795]. Converters [911]. Convex [26, 602]. Convexity [325]. Convolution [39]. Convolution-Generated [39]. Cooperative [29]. Coordinate [547]. Coordinates [512, 811, 46]. Copolymer [623]. Copolymers [925]. Corner [144]. Coronary [235]. Correction [871]. Corrections [578]. Correctors [477]. Correlation [127]. Correlations [890]. Corrigendum [816]. Corrosion [956]. Costs [928, 216]. Couette [504, 630, 318, 232]. Counter [621]. Counterexamples [992]. Counterflow [643]. Counterpropagating [470]. Coupled [600, 618, 576, 311, 38, 486, 421, 772, 369, 605, 171, 336, 200, 201, 372, 123, 752, 901]. Coupling [490, 357, 478, 162, 426, 699, 564, 326]. Couplings [143]. Covered [472]. Covering [803]. Cowan [354]. Crack [964, 907, 382]. Cracks [238, 118]. Crashing [505]. Creep [763, 184]. Creeping [781]. Criteria [601, 659]. Criterion [992, 381]. Critical [639, 301, 11, 949, 367, 800, 473]. Critically [378]. Cross [744, 192, 517, 915]. Cross-Immunity [517]. Cross-Section [192]. Cross-Sectional [915]. Crystal [668, 85, 284, 394, 74, 860, 684]. Crystalline [239, 54]. Crystals [523, 696, 885, 188, 179, 933, 670, 145]. CT [520, 278]. Cubic [988]. Cucker [794]. Culling [659]. Current [539, 540, 911, 128, 955, 831]. Current-Voltage [539]. Currents [79]. Curvature [309, 159, 871, 427]. Curvature-Based [309]. Curvature-Induced [427]. Curved [68, 903]. Curves [80, 776, 159, 943]. Customers [59]. Customizable [896]. Cutoff [679]. Cycle [47, 561]. Cycles [904, 876]. Cyclic [719]. Cylinder [656, 727]. Cylindrical [818, 272, 547, 46, 398]. Czochralski [644].

D [596, 561, 172, 544, 881, 685, 367, 147, 145, 267, 557]. Damped [415, 168, 361, 765]. Damped-Driven [415]. Darcy [642]. Data [60, 945, 410, 134, 727, 559, 535, 895]. Data-Handling [134]. DeAngelis [624]. Death [500]. Deblurring [456]. Decay [73, 495, 75, 322, 440]. Decomposition [967]. Deconvolution [180, 981]. Defect [466, 961]. Defects [284, 846]. Deflagrations [219]. Deformable [97, 751]. Deformation [836, 581, 251, 196, 293]. Deformations [279, 522]. Deformed [611]. Degeneracy [460]. Degenerate [380, 357]. Degree [524]. Dejittering [453]. Delay [299, 499, 628, 693, 91, 136, 959, 985, 754]. Delayed [490]. Delays [622, 819, 711, 123, 796]. Dendritic [818, 113, 519]. Denoising [368, 648, 788]. Density [55, 541, 553, 557]. Dependent [373, 94, 221, 207, 940, 638, 604, 743, 495, 777, 681, 745, 947, 508]. Depinning [330]. Deposition [387]. Depressed [108]. Depression [222]. Depth [790]. Derivation [297, 87, 677, 76, 49, 104, 189, 296, 910, 946]. Derivative [484]. Derivatives [325, 282]. Derived [256, 126]. Describe [673]. Described [724]. Describing [791]. Description [191, 117, 177, 5]. Design [176, 513, 621, 480, 383, 530, 514, 62, 772, 765, 896, 965, 995, 391]. Desorption [290]. Destinations [807]. Detachment [992]. Detecting [923]. Detection [487, 248, 12, 185, 339, 268, 788]. Determination [410, 510]. Detonation [148, 592]. Develop [943]. Development [665]. Deviations [237]. Device [588]. Devices [383, 955]. Diabetes [719]. Diagonalizable [356]. Diagonalization [176]. Diagram [925, 626]. Diagrams [389, 400]. Diameter [195]. Diblock [925, 623]. Dielectric [510, 588, 975, 230, 895]. Diffeomorphic [799]. Difference [167]. Differential [167, 299, 927, 798, 693, 242, 165, 754, 96]. Differential-Difference [167]. Differentially [329]. Diffraction [240, 26, 865, 961, 320, 685, 975, 476, 907, 986]. Diffuse [414]. Diffusers [211]. Diffusion [600, 83, 73, 818, 330, 8, 859, 944, 828, 298, 167, 886, 38, 374, 368, 728, 713, 447, 499, 98, 483, 229, 126, 287, 35, 934, 420, 66, 205, 913, 918, 71, 935, 926, 244, 681, 745, 947, 784, 579, 960, 267, 548, 351, 801, 173, 445, 752, 247, 951, 459, 460, 51]. Diffusional [443]. Diffusional-Thermal [443]. Diffusions [157]. Diffusive [263, 231, 906, 770, 768, 537, 199, 970, 905, 455]. Diffusivity [587, 479]. Dilatation [964]. Dilute [748]. Dimension [11, 864]. Dimensional [53, 569, 333, 729, 782, 148, 673, 14, 155, 284, 275, 903, 541, 902, 964, 164, 205, 913, 271, 476, 502, 183, 227, 898, 243, 669, 291, 283, 974, 957, 938, 482, 878, 965, 995, 74, 803, 888, 889, 877, 252]. Dimensionality [93]. Dimensions [19, 620, 185, 778, 916, 742, 910]. Diodes [285]. Dioxide [451]. Dipole [198, 653]. Direct [180, 957]. Direction [639, 445].

Directional [860]. Directions [26]. Dirichlet [544, 661, 816, 769, 351]. Dirichlet-Integral [769]. Discharge [833]. Disclinations [571]. Discontinuities [809]. Discontinuous [718, 359, 537, 915]. Discrete [618, 311, 330, 525, 14, 172, 447, 2, 109, 654, 321, 132, 645, 130]. Discrete-Time [645, 130]. Discretization [442]. Discs [827]. Disease [137]. Diseases [694]. Dislocation [471, 188, 30]. Dispersal [636, 531, 640]. Dispersed [5]. Dispersion [337, 327, 253, 879, 316, 440, 872, 878, 132, 427]. Dispersion-Managed [316, 440]. Dispersions [177, 320, 597]. Dispersive [446, 403, 458, 572, 422]. Dissipative [845, 502]. Distance [528, 978]. Distributed [129, 137, 91, 796, 306]. Distribution [161, 59, 856, 855, 264, 534]. Disturbances [859, 232]. Disynaptic [740]. Divergence [381]. Divider [829]. Division [237]. DNA [803]. Domain [962, 906, 124, 461, 589, 632]. Domains [373, 178, 439, 138, 803]. Dominance [132]. Dominated [125]. Dopant [862]. Döring [255]. Dose [514]. Double [551, 329, 905, 912]. Double-Diffusive [905]. Doubly [313]. Down [791, 755]. Downward [535]. Dpp [552]. Dpp-Receptor [552]. Drag [759]. Drainage [133]. Draw [631]. Drawing [862]. Drift [600, 178, 828, 447, 918, 187]. Drift-Collision [178]. Drift-Diffusion [600, 828, 447, 918]. Drilling [669, 274]. Driven [606, 415, 663, 233, 4, 641, 931, 724, 863, 628, 159, 280, 647, 208, 646, 362, 852, 459, 460, 936, 598]. Drivers [796]. Drop [379]. Droplet [160, 88, 982]. Drops [729, 853, 259]. Dual [819, 619, 769]. Dual-Access [769]. Dual-Phase-Lag [619]. Duct [418, 465, 526]. Ducted [533]. Due [622, 422, 108, 867]. Duffing [586]. During [770, 242, 604]. Dynamic [655, 85, 642, 807, 303]. Dynamical [37, 84]. Dynamically [382]. Dynamics [616, 610, 161, 210, 313, 341, 762, 182, 57, 137, 695, 739, 690, 550, 949, 519, 701, 702, 275, 369, 902, 672, 193, 568, 49, 652, 759, 614, 71, 171, 304, 517, 681, 498, 710, 899, 89, 104, 189, 175, 736, 232, 433, 434, 783, 247, 959, 757, 840, 617].

Ear [332]. Early [270]. Easter [505]. Eddies [164]. Eddy [79, 287, 831]. Eddy-Current [831]. Edge [788]. Edges [224]. Effect [904, 994, 514, 627, 5, 672, 531, 700, 106, 842]. Effective [6, 368, 837, 155, 920, 921, 479, 814, 649, 613, 42, 366, 258]. Effects [586, 696, 664, 86, 767, 702, 552, 111, 777, 670, 796, 186, 985]. Efficient [610, 118, 953]. Egg [892]. Egg-Limited [892]. Eigenfunctions [413, 31, 32]. Eigenvalue [944, 219, 960]. Eigenvalues [661, 816, 31, 32]. Eigenvalues/Eigenfunctions [32]. Eikonal [300]. Einstein [610]. Elastic [195, 836, 257, 432, 497, 3, 937, 737, 580, 916, 36, 790, 518, 265, 964, 685, 522, 507, 7, 45, 878, 965, 995, 860, 910]. Elastica [309]. Elasticity [238, 654, 348, 609, 296]. Electric [181, 833, 939, 212, 709]. Electrical [487, 836, 804, 923, 419, 699, 927, 718, 734, 186]. Electrical-Optical [186]. Electrically [171]. Electrified [736, 969]. Electro [427]. Electro-Osmotic [427]. Electrochemical [539, 540]. Electrode [419, 812, 199]. Electrodeposition [896]. Electrodiffusion [128]. Electrolyte [345, 950, 951]. Electromagnetic [887, 621, 566, 826, 284, 66, 975, 281, 169, 748, 409, 74]. Electromagnetism [176]. Electron [869, 530, 703, 821, 463, 7]. Electron-Phonon [463, 7]. Electrons [900]. Electrophoretic [398]. Electrorheological [547]. Electroseismic [538, 678]. Electrostatic [695, 670, 230]. Electrowetting [914]. Element [669, 991, 300]. Elementary [251, 917]. Elements [402]. Elliptic [811, 660, 477, 25].

Elliptical [744]. Embedded [263, 80, 972, 592]. Embedding [986]. Emerging [280]. Emulators [872]. Encapsulation [20]. Encounter [892]. Endemic [129, 137]. Energetic [219, 307, 308]. Energy [194, 717, 627, 421, 605, 568, 326, 970, 593]. Engineering [562]. Enhanced [837]. Ensembles [630]. Entropy [101, 677, 963]. Enumeration [968]. Envelope [105]. Environment [545, 645]. Enzymatic [688]. Epidemic [554, 735, 558, 390, 193, 545, 640, 645]. Epidemiological [475]. Epilimnion [783]. Epistemic [594]. Epitaxial [296, 910]. Epitaxially [239]. Equation [806, 705, 415, 758, 28, 791, 198, 653, 19, 629, 226, 823, 575, 880, 543, 238, 342, 446, 11, 168, 778, 703, 148, 724, 387, 755, 462, 934, 975, 798, 183, 693, 765, 623, 563, 95, 986, 165, 820, 300, 343, 607, 757, 658]. Equations [454, 53, 102, 79, 249, 830, 595, 444, 428, 194, 618, 161, 457, 330, 845, 221, 256, 64, 477, 167, 525, 403, 299, 908, 837, 486, 868, 881, 98, 400, 421, 483, 126, 78, 677, 768, 30, 921, 863, 404, 227, 367, 146, 244, 722, 46, 371, 203, 407, 613, 104, 189, 366, 386, 754, 43, 946, 258, 96]. Equilibria [542, 634, 348]. Equilibrium [239, 114, 928, 89]. Eradication [841]. Ericksen [48]. Erlang [384, 378, 930]. Erratum [590, 947, 674, 189, 995]. Error [157]. Eshelby [874]. Essential [605, 206, 485]. Estimate [287]. Estimates [6, 504, 709]. Estimating [890]. Estimation [120, 788, 184, 435]. Estimator [121]. Euclidean [452]. Euler [968, 309, 881, 227, 367]. Eulerian [729]. Evaluated [744]. Evaluating [133]. Evaluation [26, 927]. Evans [338]. Evaporation [51]. Even [778]. Events [473]. Everted [272]. Evolution [606, 125, 516, 636, 75, 159, 461, 850, 43, 799]. Exact [705, 516, 36, 421, 100, 287, 278, 654, 963, 31].

Exchange [86, 949, 426, 153]. Excitability [108]. Excitable [311, 109, 108, 601, 306]. Excitation [733, 831, 450, 300]. Excitatory [481, 478, 567]. Excited [488, 942]. Exclusion [838, 686, 617]. Exclusive [866]. Exercise [270]. Exhibits [390]. Existence [782, 839, 229, 65, 316, 623, 283, 708, 511, 585, 919, 69]. Exit [628, 264]. Exotic [436]. Expansion [227, 58, 317]. Expansions [13, 9]. Experiment [112]. Experimental [956]. Experiments [638]. Exploiting [520, 777]. Explosion [301]. Explosions [117, 789]. Exponent [310, 637, 225]. Exponential [273, 886]. Exponents [198]. Expressions [856]. Extended [99, 990, 608, 411]. Extensional [729]. Extensions [223]. Exterior [891, 632]. External [766]. Extinction [694, 615, 659]. Extracellular [116]. Extremal [489]. Extreme [496].

Fabry [74]. Factorization [990, 826, 804, 812]. Faddeev [462]. Fall [729]. False [981]. Fan [520]. Far [887, 410, 473, 895]. Faraday [725]. Fast [512, 22, 383, 137, 703, 787]. Faster [281]. Feasibility [848]. Features [152]. Feed [957]. Feedback [596, 884, 106]. Fekete [944]. FENE [822]. Fermi [463]. Ferromagnetic [155]. Ferromagnets [426, 461]. Ferronematic [842]. Fiber [966, 791, 655, 880, 763, 755, 862, 565, 489, 652, 759]. Fiber-Reinforced [763]. Fibers [54]. Field [454, 887, 571, 834, 410, 276, 890, 560, 976, 15, 447, 202, 496, 939, 473, 788, 203, 769, 895, 877]. Field-Field [890]. Field-Induced [571]. Fields [503, 857, 364, 464, 559, 709]. Film [245, 140, 758, 198, 653, 315, 179, 908, 868, 472, 280, 133, 646, 969]. Films [539, 239, 376, 540, 797, 647, 736, 684]. Filter [218]. Filtered [278, 776]. Filtering [18]. Filters [368, 980, 882]. Filtration [984]. Finding [648]. Fine [471, 518].

Finger [228]. Fingering [813]. Finite [503, 402, 839, 790, 702, 138, 692, 669, 991, 251, 970, 300, 196, 293, 943]. Fire [994, 37, 143, 810]. Fires [673]. Firing [478]. First [715, 944]. First-Order [715]. Fisher [19]. Fitting [398]. FitzHugh [525, 171, 305]. Five [866]. Five-Component [866]. Fixed [157, 386]. Fixed-Point [157]. Flame [716, 498, 680]. Flames [443, 331, 615, 970, 31, 32, 593]. Flexibility [127, 38]. Flexible [472]. Floating [916, 790]. Flocculated [515]. Flocking [794]. Floods [304]. Flow [829, 418, 10, 41, 297, 598, 663, 355, 581, 218, 504, 85, 682, 744, 34, 395, 67, 314, 262, 516, 642, 546, 252, 465, 861, 698, 162, 82, 364, 724, 891, 405, 400, 5, 547, 70, 266, 164, 479, 139, 318, 495, 507, 331, 779, 898, 827, 704, 225, 691, 883, 407, 626, 796, 235, 617, 501, 969, 232, 786, 888, 889, 943, 258, 360]. Flow-Referenced [262]. Flowpath [67]. Flows [373, 125, 140, 533, 494, 630, 215, 753, 231, 374, 86, 544, 584, 652, 759, 208, 291, 427, 799]. Fluid [705, 161, 423, 598, 34, 807, 256, 949, 687, 495, 111, 611, 898, 725, 915, 781]. Fluidized [17]. Fluids [86, 518, 547, 271, 366]. Focus [561, 612]. Focus-Center-Limit [561]. Focused [469]. Focusing [887, 342, 11, 168, 253, 425, 835]. Fokker [791, 703, 755, 216]. Fold [122]. Folding [44]. Follow [292, 297]. Follow-the-Leader [292, 297]. Following [324]. Food [213, 745, 947]. Force [665, 557]. Forced [234, 526, 354]. Forces [606]. Forcing [806, 422]. Forest [994]. Form [467, 977, 65, 856]. Formal [353, 360]. Formation [715, 966, 556, 105, 228, 723, 881, 858, 590, 706, 921, 699, 491, 982, 110, 144, 434, 112, 684]. Formed [647]. Formula [649, 986, 930]. Formulae [195]. Formulas [632]. Formulation [675, 529, 444, 845, 851, 960]. Formulations [891]. Forward [772, 134, 473].

Forward-Adjoint [772]. Four [174, 879]. Four-Phase [174]. Four-Wave [879]. Fourier [52]. Fourth [253, 863]. Fourth- [863]. Fourth-Order [253]. Fractal [363]. Fraction [668]. Fractional [374, 229, 75]. Fractures [692]. Framework [701, 848, 652, 692, 81, 888, 889]. Fréchet [282, 484]. Frederiks [842]. Free [973, 705, 929, 962, 140, 73, 854, 233, 717, 891, 685, 133, 365, 850, 175, 912, 680]. Free-Moving [73]. Free-Radical [365]. Free-Space [962]. Freedom [524]. Frequencies [478]. Frequency [94, 191, 55, 207, 940, 342, 177, 927, 404, 170, 362, 386]. Frequency-Averaged [404]. Frequency-Dependent [94]. Frequent [216]. Frictionless [118]. Front [481, 71, 269]. Fronts [375, 813, 641, 760, 647, 200, 115, 286]. Frozen [374]. Fuel [27, 956, 958, 949, 953, 199, 948, 957, 345, 950, 951]. Fully [148]. Function [555, 536, 890, 587, 38, 713, 338, 462, 52, 159, 820, 557]. Functional [925, 158, 985, 612, 312]. Functionals [637]. Functions [981, 432, 527, 780, 541, 528, 632, 888, 889]. Fundamental [129, 400, 626]. Funnel [974].

Gale [389]. Gas [657, 73, 114, 672, 227, 199, 326]. Gas-Solid [73]. Gasdynamics [825]. Gaseous [593]. Gating [435]. Gaussian [18, 886]. Gels [556, 991, 840]. General [675, 529, 503, 767, 416, 776, 798, 652, 731, 810]. Generalization [823]. Generalized [554, 161, 215, 327, 565, 522, 49, 579, 39, 785, 820, 43, 58, 51]. Generated [191, 521, 124, 39, 209]. Generation [125, 574, 172, 730, 520]. Generation-Recombination [730]. Generators [277]. Generic [381]. Genesis [529, 595]. Genetic [493, 187]. Geochemical [389]. Geodesics [528]. Geological [44]. Geomagnetic [559].

Geometric [182, 671, 320, 507, 509, 267]. Geometrical [428, 809]. Geometry [470, 233, 506, 111]. Geometry-Driven [233]. geostrophic [395]. Gierer [275, 866]. Ginzburg [53, 102, 183, 149, 474, 757]. Glacier [751]. Glass [350, 706]. Gliding [103]. Global [830, 390, 341, 739, 690, 686, 91, 139, 193, 648, 1, 158, 674, 832]. Globally [372]. Glucose [711]. Glucose-Insulin [711]. Go [400]. Goal [669]. Goal-Oriented [669]. Golden [463]. Governed [486, 607]. Gower [904]. Gradient [238, 920, 502, 647]. Gradient/Skew [502]. Gradient/Skew-Gradient [502]. Gradients [377, 724, 941, 867]. Grain [459, 460]. Grains [861]. Granular [252, 405, 557]. Graph [634]. Graphs [954]. Grating [402, 476]. Gravitational [133]. Gravity [769, 643, 646]. Gravity-Field [769]. Gray [141, 182, 100, 254]. Green [432, 890, 462, 912]. Greitzer [131]. Grounded [250]. Group [210, 120]. Group-Theoretic [120]. Groups [482]. Growing [529, 595]. Growth [245, 644, 808, 637, 931, 469, 899, 843, 860, 296]. Guided [466, 708].

Hagen [691]. Half [432, 804, 907, 196, 293]. Half-Space [432, 804, 196, 293]. Hall [730]. Hamiltonian [524, 214, 154]. Handling [134]. Hannay [578]. Hard [461]. Hardening [265]. Harmonic [574, 575, 527, 172, 425, 734, 831, 360]. Harmonically [888, 889]. Harmonics [586]. Hartree [575]. Harvesting [985, 508]. Having [347]. Heart [235]. Heat [350, 404, 827, 651, 589, 269, 360]. Heat-Induced [350]. Heated [160, 329]. Heating [326]. Hedging [216, 106]. Hele [208]. Helical [63]. Helmholtz [317]. Hematopoiesis [532]. Hematopoietic [693]. Hepatitis [894]. Hepatocyte [894]. Heteroclinic [100, 743]. Heterogeneity [938]. Heterogeneous [87, 772, 832].

Hexagonal [988]. Hidden [271]. Hierarchical [787, 29, 794]. Hierarchy [376, 768]. High [454, 191, 697, 8, 342, 164, 170, 669]. High-Accuracy [669]. High-Frequency [191, 170]. High-Péclet-Number [8]. Higher [457, 543, 768, 786]. Higher-Order [457, 768, 786]. Highly [795, 169, 593]. Highways [319]. Hilbert [632]. Hill [806]. Hilliard [925, 863, 95, 607]. Hilliard-Type [863]. History [777]. History-Dependent [777]. Hitting [33]. HIV [710]. HIV-1 [710]. Hodgkin [435]. Holes [328]. Holling [323, 985]. Homeostasis [894]. Homoclinic [214, 100, 213, 876]. Homogeneity [121]. Homogeneous [117, 789, 4, 87, 393]. Homogenization [668, 64, 477, 387, 518, 496, 42, 386]. Homogenized [337, 423]. Hopf [395, 347, 357, 650, 329, 136, 612, 754]. Hoppers [405]. Hormone [135]. Horn [563]. Host [694, 892]. HPCVD [245]. Human [506]. Huxley [435]. Huygens [39]. Hybrid [779]. Hydraulic [692]. Hydrodynamic [791, 677, 294]. Hydrodynamical [99, 963, 455]. Hyperbolic [176, 675, 337, 380, 314, 107, 731, 203, 407, 69]. Hyperelastic [196, 293]. Hyperelasticity [468]. Hypersingular [238, 30]. Hysteresis [268, 493, 643]. Hysteresis-Based [493].

Ice [313, 701, 702]. Ice-Shelf [720]. Ice/Ice [720]. Ideal [371]. Idealized [994]. Identification [285, 638, 417, 798, 895]. Identifying [773]. Ignition [124]. II [595, 12, 634, 182, 137, 851, 75, 244, 201, 157, 32, 460, 985, 889]. III [644, 673]. III-V [644]. Ill [549]. Ill-posed [549]. Illiquid [106]. Illustration [746]. Image [377, 310, 456, 152, 637, 207, 408, 671, 448, 368, 399, 771, 732, 648, 267, 788, 453, 445]. Images [945, 151, 780, 81]. Imaging [487, 844, 285, 220, 420]. Immersed [497, 396]. Immune [754]. Immunity [517]. Impact [241]. Impact-Induced [241]. Impacted [964]. Impedance [487, 836, 240, 410, 873, 484, 804, 923, 419, 927, 718, 812, 734, 727]. Impedance-Acoustic [873]. Impenetrable [347]. Imperfect [750]. Imperfections [195]. Imperfectly [591, 742]. Implications [32]. Impulsive [659]. Incidence [554, 690, 767, 876]. Incipient [840]. Including [692]. Inclusion [569, 339, 874]. Inclusions [923, 419, 169]. Incompressible [194, 257, 432, 265, 479, 495, 725, 196]. Indentation [257]. Independent [211]. Index [562]. Individual [492]. Induced [571, 401, 909, 350, 241, 741, 469, 427]. Induction [124]. Infection [129, 137, 261, 690, 499, 894, 640]. Infer [777]. Infestation [841]. Infinite [284, 907, 424, 733, 938, 959]. Infinite-Dimensional [938]. Inflated [513]. Influence [681, 110]. Influenza [517]. Information [560]. Inhibition [201]. Inhomogeneities [3, 342]. Inhomogeneous [413, 197, 565, 392, 156]. Initial [598, 403, 75, 835, 346]. Initial-Boundary [403, 346]. Initiation [365]. Injection [97]. Injectivity [747, 542]. Injured [966]. Injury [506]. Inner [332]. Inpainting [309]. Inpaintings [236]. Instabilities [853, 486, 229]. Instability [573, 470, 992, 627, 401, 770, 970]. Installment [967]. Instance [990]. Instantaneous [143]. Insulin [711]. Integral [195, 891, 284, 30, 975, 146, 769, 343, 74]. Integrals [968, 456]. Integrate [37, 143, 810]. Integrate-and-Fire [37, 143, 810]. Integrodifferential [845, 238]. Interacting [423, 576, 2, 104, 189]. Interaction [101, 682, 188, 488, 301, 647, 356, 7, 846, 910, 757]. Interactions [333, 323, 622, 892, 917, 783]. Interest [325]. Interface [191, 210, 931, 625, 414, 202, 50, 379, 156, 553, 39]. Interfaces [94, 83, 518, 903]. Interfacial [907]. Interlayer [426]. Intermittency [335]. Intermittent [884]. Internal [98, 838, 977, 553, 501]. Interpolation [945]. Interrogation [621]. Interval [364]. Intervention [604]. Intracellular [941]. Intrahost [686]. Introducing [639]. Invariances [271]. Invariant [430, 468, 993]. Invasion [23]. Inverse [667, 249, 566, 629, 721, 527, 738, 282, 937, 683, 792, 826, 220, 591, 742, 340, 851, 914, 295, 775]. Inversion [60, 512, 22, 778, 541, 278, 776, 435, 930]. Inverted [782]. Investigation [753, 975, 318]. Investor [815]. Inviscid [215, 883]. Involving [2]. Ion [721, 974]. Ionic [128, 321]. Ionization [813]. Ionospheric [573, 559]. Irreversible [605]. Isentropic [455]. Island [505]. Isolation [517]. Isotropic [432, 265, 522, 907, 424, 760, 196, 293]. Isotropy [393]. Issues [956]. Iterated [464].

Jet [971]. Jets [271]. Joining [767]. Joint [788]. Joseph [50]. Josephson [714, 65]. Jouguet [824, 785]. Jump [126, 672]. Junction [65]. Junctions [714, 15]. Justification [79]. Juvenile [922, 622]. Juvenile-Adult [922]. Juxtacrine [204].

KdV [322]. Keldysh [441]. Keller [587, 433, 434]. Kelvin [800]. Kernel [769, 589]. Kernels [96]. Kill [370]. Kinematic [482]. Kinetic [454, 256, 774, 5, 549, 76]. Kinetics [87, 931, 713, 599]. Kirchhoff [954, 507, 835]. Known [591, 742]. Kramers [264]. Kuramoto [343].

Lacunae [748]. Laden [797]. Lag [619]. Lagrangian [701, 779, 291]. Lamellar [3]. Lampis [672]. Landau [757, 53, 102, 183, 149, 474]. Landscapes [636]. Landslide [884]. Laplace [820]. Laplacian [774]. Large [101, 59, 741, 136, 815, 269, 346, 593, 237, 616]. Laser [136, 469, 186, 274]. Lasers [847]. Late [210]. Lateral [201]. Latitudinal [388]. Lattice [110, 599]. Law [82, 416, 150, 593]. Laws [310, 188, 69]. Lay [791, 755]. Lay-Down [791, 755]. Layer [257, 544, 983, 800, 898, 905, 343, 912]. Layered [826, 352, 623, 965, 995, 678]. Layers [61, 675, 263, 228, 551, 155, 442, 46]. LBIC [285]. Leader [292, 297]. Leadership [794]. Leading [224]. Leak [692]. Leak-Off [692]. Length [234, 679]. Lens [391]. Leslie [904]. Leukemia [532, 12]. Level [671, 267, 850]. Level-Set [671]. Lévy [18, 310]. Lewis [697]. Lifshitz [255]. Lift [744]. Light [16]. Lights [292]. Limit [904, 140, 791, 653, 575, 255, 561, 447, 126, 66, 139, 13, 244, 597, 104, 189, 876, 433, 434, 455]. Limited [728, 892]. Limiting [540, 133]. Limits [231, 169]. Line [466, 714, 49, 871, 946]. Linear [140, 279, 653, 637, 12, 561, 56, 382, 336, 564, 548, 691, 95, 534, 269, 987]. Linear-Nonlinear [336, 564]. Linearized [880, 724]. Lines [406, 92, 627]. Lineshapes [650]. Link [577]. Liquid [696, 581, 85, 885, 114, 57, 54, 933, 971, 379, 472, 280, 647, 670, 371, 646, 957, 736, 684]. Liquid-Feed [957]. Liquid-Gas [114]. Liquid/Vapor [57]. Liquidation [815]. Liquids [421]. Living [29]. Lloyd [649]. Load [958, 196, 293]. Loaded [27, 378]. Loading [111]. Local [989, 845, 236, 859, 449, 121, 821, 393]. Localization [413, 82, 921, 372]. Localized [944, 283, 326]. Locating [726]. Location [250, 156]. Locked [567, 628]. Locus [592]. Log [512]. Log-polar [512]. Logging [669]. Logging-While-Drilling [669]. Logistic [587]. Long [922, 762, 897, 597, 910, 658]. Long-Range [910]. Long-Time [922]. Look [430, 617]. look-ahead [617].

Lookback [583]. Loop [628]. Loops [490, 474, 876]. Lorentzian [650]. Loss [913]. Lossy [942]. Lost [412]. Lotka [375, 959, 919]. Low [523, 373, 6, 310, 55, 478, 551, 387, 447, 177, 183, 13, 651]. Low-Contrast [523]. Low-Dimensional [183]. Low-Field [447]. Low-Temperature [651]. Lower [989]. Lump [125]. Lung [170]. LWD [669]. Lyapunov [225].

Mach [373, 13]. Macroscopic [76, 7]. Magnetic [581, 734, 955]. Magnetoencephalography [249]. Magnetohydrodynamics [346]. Magnetostrictive [707]. Make [281]. Malaria [676, 686]. Male [135]. Malignant [23]. Malliavin [924]. Managed [316, 440]. Manifold [639, 420, 131]. Manifolds [246]. Manufacture [417]. Many [206, 371, 104, 189]. Many-Particle [104, 189]. Map [662, 979, 360]. Marangoni [280]. Marangoni-Driven [280]. Marble [854]. Markets [106]. Markov [839, 130, 500]. Martensitic [599]. Mass [5, 613, 501, 345]. Master [934]. Matched [675, 442, 9]. Material [621, 713, 241, 242, 184]. Materials [279, 717, 940, 265, 522, 219, 281, 307, 707, 308, 842]. Mathematical [532, 451, 218, 956, 236, 676, 116, 92, 961, 246, 802, 841, 562, 565, 848, 833, 230, 948, 710, 20, 186, 485, 594]. Mathematics [854]. Matrix [747, 911, 116, 298, 449, 603, 411]. Matter [906]. Maturation [622, 666]. Maximizers [568]. Maximum [546, 963]. Maxwell [79, 249, 46, 386]. McKean [305]. McKendrick [221]. Mean [944, 976]. Mean-Field [976]. Means [778]. Measurements [166, 923, 669]. Measures [604]. Mechanics [449, 170]. Mechanism [478, 29, 372]. Mechanisms [665, 260]. Media [989, 349, 844, 576, 34, 67, 738, 458, 817, 826, 197, 352, 108, 851, 897, 760, 988, 601, 175, 878, 965, 995, 951, 678, 306, 557, 594, 286, 942]. Mediated [172, 98]. Medical [420]. Medium [195, 705, 566, 984, 890, 413, 14, 172, 109, 66, 870]. Meinhardt [275, 866]. Melnikov [697]. Melt [469]. Melting [160, 242]. Membrane [385, 6, 818, 949]. Membranes [513, 128]. Memory [855, 289, 203, 796]. MEMS [695, 588, 230]. Meniscus [280, 647]. Menten [743]. Menten-Type [743]. Mesenchymal [864]. Messages [412]. Metal [581]. Metallic [956]. Metastability [494]. Metastable [35]. Methanol [957]. Method [990, 729, 310, 981, 151, 12, 105, 937, 737, 703, 272, 826, 804, 927, 718, 975, 396, 812, 382, 340, 52, 741, 582, 302, 669, 722, 579, 267, 788, 300, 799]. Methods [811, 448, 954, 906, 397, 991, 965, 995]. Michaelis [743]. Micro [258]. Microbial [838, 70]. Microelectrodes [350, 706]. Microfluid [397]. Microfluidic [327]. Microlocal [176]. Micromixers [437]. Microphase [925]. Microscope [665]. Microscopic [297, 822]. Microscopy [821]. Microstrip [406, 714, 946]. Microstructure [496]. Microwaves [12]. Middle [978]. Middle-Distance [978]. Midlatitude [224]. Mie [559]. Migration [359, 537]. Mildly [620]. Milne [454]. Minimal [468]. Minimization [677]. Minimizers [648, 609]. Minimum [717]. Mirrors [712, 425]. Mix [761]. Mixed [417, 960, 991]. Mixing [879, 281]. Mixture [529, 595, 471, 724, 593]. Mobile [798]. Mobility [398]. Modal [961, 42]. Mode [162, 396, 397, 784, 872]. Mode-Coupling [162]. Model [16, 922, 904, 793, 606, 554, 735, 79, 160, 558, 99, 451, 292, 21, 390, 616, 966, 929, 600, 635, 27, 644, 956, 323, 808, 515, 834, 744, 34, 676, 276, 113, 642, 828, 807, 116, 298, 344, 4, 141, 182, 587, 666, 875, 861, 886, 976, 688, 87, 57, 446, 129, 261, 580, 695, 802, 414, 128, 673, 739, 690, 713, 858, 679, 755, 447, 730, 499, 841, 223, 767, 662, 550, 716, 155, 449, 100, 202, 932, 419, 562, 35, 275, 902, 135, 303, 913, 833, 213, 289, 812, 91, 139, 193, 711, 624, 23, 219, 331]. Model [90, 779, 254, 242, 704, 498, 745, 947, 978, 963, 474, 307, 469, 884, 970, 453, 785, 20, 165, 974, 511, 235, 665, 353, 917, 957, 876, 300, 305, 482, 433, 434, 345, 435, 545, 640, 485, 860, 296, 131, 910, 754, 645, 608, 308, 901, 51, 680, 599]. Modeled [866]. Modeling [523, 97, 887, 245, 380, 573, 530, 48, 682, 859, 514, 519, 604, 701, 702, 719, 652, 935, 914, 230, 225, 769, 796, 617, 103, 535, 896, 950, 951, 684]. Modelling [218, 313, 188, 437, 304, 991, 186]. Models [375, 715, 41, 297, 337, 452, 44, 152, 236, 376, 854, 256, 17, 560, 698, 857, 129, 137, 246, 980, 949, 686, 76, 603, 183, 475, 822, 926, 648, 894, 1, 674, 7, 187, 659, 274, 843, 353, 184, 864, 786, 936, 455, 877]. Moderately [104, 189]. Modes [565, 800, 708]. Modica [732]. Modified [192]. Modulated [141, 182, 242]. Modulational [68]. Modulations [224]. Moisture [770]. Molding [97]. Molecular [38]. Molten [958]. Moment [703, 839, 768, 693, 963]. Moments [629, 856, 722]. Monotone [691]. Monte [924, 772]. Monument [854]. Moore [131]. Morphological [408]. Mortola [732]. Mosaic [40]. Motion [973, 471, 430, 571, 374, 625, 5, 33, 379, 814, 506, 120, 611, 291, 788, 371, 39, 326, 884, 646, 433, 864, 781]. Motions [308, 942]. Motor [38]. Mountain [656]. Moving [73, 774, 15, 687]. Multibump [226]. Multicell [595]. Multicomponent [345]. Multidimensional [753, 455]. Multidisease [614]. MultiFrequency [848, 483]. Multigraphs [980]. Multilane [319]. Multilateral [847]. Multilayer [955]. Multilevel [22]. Multimodal [448]. Multimode [775]. Multiparametric [635]. Multiparasitoid [892]. Multiparasitoid-Host [892]. MultiPhase [15, 67, 949, 732]. Multiple [59, 847, 542, 634, 15, 414, 164, 289, 649, 13, 750, 212, 123, 941, 77, 803, 612, 237]. Multiple-Scales [847]. Multiplicative [805]. Multiply [439]. Multiplying [548]. Multipole [208]. Multiresolution [42]. Multiscale [24, 936]. Multiservice [384, 378]. Multispike [275]. Multistability [633, 123]. Multisymplectic [470]. Multivalued [400]. Mumford [671]. Muscle [802, 77]. MUSIC [990]. Mutating [475]. Mutation [928]. Mutation-Selection [928]. Mutually [866]. Myelogenous [532]. Myocardial [300]. Myriad [348]. Myxobacteria [103].

Nagai [590, 491]. Nagumo

[618, 525, 171, 305]. Nanoparticle [771]. Nanowire [900]. Narrow [474, 764, 974]. Natural [513, 301, 565]. Nature [403]. Navier [194, 315, 421]. Near [443, 355, 114, 650, 895, 191, 357, 173]. Near-Stoichiometric [443]. Nearly [993]. Neck [740]. Necrotic [808]. Néel [551]. Negative [596, 568]. Nematic [571, 696, 48]. Nernst [509, 974]. Nerve [966]. Network [481, 490, 252, 567, 882, 777, 243, 407]. Networks [616, 37, 633, 527, 542, 634, 807, 660, 327, 954, 40, 564, 200, 201, 893, 163, 832]. Neural [503, 481, 576, 490, 633, 857, 567, 40, 855, 795]. Neuronal [289, 200, 201, 450, 89, 163]. Neurons [478, 243, 143, 123, 810]. Newtonian [423, 218, 271, 124]. Nile [902]. Nine [902]. Nine-Dimensional [902]. No [340]. Node [852]. Noise [18, 805, 185, 465, 909, 138, 628, 217, 334, 741, 336, 700]. Noise-Induced [741]. Noisy [52]. Non [423, 218, 650, 271]. Non-Lorentzian [650]. Non-Newtonian [423, 218, 271]. Nonadiabatic [578, 331, 615]. Nonautonomous [922]. Nonaxisymmetric [405]. Noncommuting [169]. Nonconservative [441]. Nonconvex [865, 785]. Noncrystallographic [471]. Nondecaying [60]. Nondestructive [927]. Nondiffusive [263, 231]. Nondivergent [544]. Nonexistence [603, 708]. Nonflat [152]. Nonhomogeneous [738, 851]. Nonhyperbolic [214]. Nonisentropic [915]. NonLeaky [143]. Nonlinear [16, 830, 415, 428, 388, 28, 869, 470, 226, 313, 574, 234, 763, 357, 477, 342, 446, 11, 168, 361, 368, 148, 690, 14, 172, 572, 447, 216, 550, 416, 964, 30, 628, 352, 800, 159, 741, 302, 926, 336, 564, 681, 704, 722, 251, 7, 348, 609, 613, 752, 876, 689, 810, 736, 232, 204, 43, 96]. Nonlinearity [605]. Nonlinearly [61, 265, 45]. Nonlocal [925, 35, 670, 901]. Nonmixing [231]. Nonmonotonic [158, 612, 312]. Nonnecrotic [808]. Nonparaxiality [361]. Nonreciprocal [181]. Nonreflecting [36]. Nonreflective [62]. Nonrigid [408, 448]. Nonrotating [388, 568]. Nonscattering [773]. Nonsingular [26]. Nonsteady [421]. Nontexture [236]. Nonuniformly [606]. Nonuniqueness [65]. Nonwoven [755]. Normalizing [524]. Novel [799]. Nozzle [915]. Nucleated [767]. Number [412, 373, 697, 8, 277, 882, 164, 13, 938]. Numbers [277]. Numerical [712, 451, 610, 218, 956, 834, 845, 753, 477, 15, 849, 550, 220, 118, 462, 30, 318, 914, 827, 46, 960, 957, 950, 943, 979]. Nutrient [838, 70, 1, 674, 901]. Nutrients [570].

Objects [263, 452, 895]. Obstacle [140, 410, 313, 937, 484, 33]. Obstacle-Type [140]. Obstacles [773, 392]. Occupancy [277]. Ocean [4, 225]. Off [692]. One [53, 333, 729, 148, 673, 14, 155, 275, 964, 502, 243, 878, 965, 995, 864, 803, 877, 317]. One-Dimensional [53, 333, 729, 148, 673, 14, 155, 275, 964, 502, 243, 878, 965, 995, 803, 877]. One-Way [317]. Open [569, 961]. Opening [764].

Operation [186]. Operator

[194, 629, 940, 317]. Operators [534].

Opinion [715]. Optical

[10, 772, 862, 726, 565, 147, 784, 186, 684]. Optics [428, 809]. Optimal [18, 989, 383, 166, 958, 413, 886, 47, 370, 978, 157, 815, 965, 995]. Optimization [530, 958, 402, 937, 399, 438, 701, 702, 334]. Optimize [514]. Option [929, 358, 270]. Options [583, 467, 436, 967]. Orbit [214]. Orbits [80, 603]. Order [715, 41, 245, 457, 834, 543, 253, 399, 768, 863, 75, 786]. Organizing [656]. Oriented [558, 669]. Orifice [584]. Orographic [422]. Orthonormal [364]. Oscillating [55, 301, 518]. Oscillation [601, 115]. Oscillations [61, 222, 234, 526, 666, 688, 299, 909]. Oscillators [586, 37, 78, 369, 699, 605, 493, 171, 356, 372]. Oscillatory [355, 138, 691, 351]. Oseen [744]. Osmotic [427]. Ostwald [868]. Other [981, 720]. Outbreak [604]. Oval [982]. Overcoming [520]. Overfilled [409]. Overloaded [577]. Overshoot [290]. Overview [949, 948]. Oxide [27, 952]. Oxygen [77].

Paced [655]. Padé [52]. Pairs [874, 328]. Pairwise [564]. Palm [855]. Parabolic [606, 380, 457, 313, 148, 134, 798, 156, 986]. Parabolic-Hyperbolic [380]. Paradox [689, 825]. Parallel [953]. Parameter [285, 638, 417, 149, 435]. Parameters [184]. Parametric [497, 440]. Parametrically [488]. Part [701, 702, 851, 459, 460, 888, 889]. Partial [512, 517]. Partially [818, 410, 510, 801]. Particle [797, 687, 397, 761, 104, 189]. Particle-Based [397]. Particle-Laden [797]. Particles [423, 2, 971]. Partly [118]. Pass [656]. Passage [944, 572, 214]. Passive [605, 479]. Passivity [707]. Past [205, 918]. Patch [735]. Patches [24].

Patchy [636, 545]. Path [374]. Pathogens [475]. Pathologies [981]. Paths [471]. Pattern [966, 105, 858, 40, 699, 745, 947, 982, 110, 725, 866, 112]. patterned [684]. Patterning [936]. Patterns [486, 138, 502, 531, 450, 511, 163, 482, 204, 607]. Payoff [583]. PDE [480, 958, 660, 216, 131]. PDEs [811]. Péclet [8]. Peierls [444, 30]. PEM [956, 948]. Pencil [940]. Pendant [259]. Pendulum [782]. Penetrant [290]. Percussion [274]. Perfectly [675, 91, 442]. Performance [851, 684]. Period [700, 640]. Periodic [102, 248, 782, 327, 167, 38, 940, 181, 14, 172, 964, 138, 733, 71, 351, 708, 878, 919, 607]. Periodically [521, 354]. Periodicity [248]. Periods [129, 137]. Permanent [137, 977]. Permeability [6, 643]. Permittivity [212]. Perot [74]. Perpetual [467]. Perspective [579]. Perturbation [21, 28, 509, 219, 146]. Perturbations [358, 436, 983, 741, 766]. Perturbative [697]. Perturbed [660, 11, 903, 784, 43, 130]. Phage [843]. Phase [606, 578, 834, 48, 114, 925, 276, 314, 174, 560, 389, 698, 57, 952, 15, 202, 78, 369, 732, 356, 788, 619, 149, 203, 307, 651, 599, 308, 877, 84, 258]. Phase-Field [276, 202]. Phases [808, 414]. Phenomena [847, 932, 842]. Phonon [463, 7]. Photonic [523, 466, 668, 179, 849, 284, 145, 394, 74]. Phytoplankton [16, 901]. Phytoplankton-Nutrient [901]. Piecewise [561, 52]. Piezoelectric [279]. Pile [188]. Pile-Up [188]. Pinching [88]. Pinning [247]. Pipe [86, 495]. Pitchfork [214]. Planar [27, 381, 31, 32, 346, 943]. Planck [791, 703, 755, 216, 509, 974]. Plane [102, 240, 504, 630, 318, 159, 232, 269]. Planes [907]. Plasma [573, 362, 748]. Plasmas [813]. Plastic [368, 920]. Plate [916, 790, 328, 835]. Platelet [630]. Plates [956]. Plug [70]. Point [191, 430, 355, 388, 714, 265, 424, 331, 528, 769, 173, 157, 689, 825, 433, 434, 196, 293, 942]. Point-Source [942]. Point-Vortices [388]. Points [944, 933]. Poiseuille [630, 691]. Poisson [900, 600, 178, 446, 367, 509, 974]. Poissonian [211]. polar [512]. Polarization [72, 339, 872]. Polarization-Mode [872]. Pole [31, 32]. Polydisperse [380]. Polygon [865]. Polymer [262, 290, 54, 767, 345, 950, 951]. Polymer-Electrolyte [951]. Polymer-Penetrant [290]. Polymerization [767, 365]. Polymers [256]. Polynomial [928, 364, 524]. Population [21, 635, 639, 452, 603, 493, 766, 659, 938, 640, 959, 919]. Populations [505, 694, 932, 531, 450, 89, 936]. Porcupine [445]. Pore [884, 398]. Pore-Pressure [884]. Porous [97, 705, 984, 34, 67, 817, 219, 898, 199, 905, 175, 594, 286]. posed [549]. Posedness [922, 448, 675]. Postcritical [473]. Potential [575, 38, 177, 5, 951]. Potentials [60]. Powder [190]. Power [911, 750]. Predator [904, 323, 622, 344, 875, 739, 743, 624, 158, 985, 508, 612, 132, 312]. Predator-Prey [904, 323, 739, 743, 158, 508, 612, 132, 312, 622, 624, 985]. Prediction [291]. Preemptive [679]. Preforms [97]. Premium [967]. Premixed [443, 970]. Prenatal [420]. Presence [18, 122, 905]. Present [292]. Pressure [642, 641, 724, 387, 884]. Pressure-Driven [641]. Prey [904, 323, 344, 875, 739, 743, 158, 508, 612, 132, 312, 622, 624, 985]. Prey-Predator [344, 875]. Price [325]. Pricing [358]. Principle [922, 279, 546, 153, 686]. Principles [197, 39]. Prion [767]. Priority [679, 303]. Probabilistic [72]. Probabilities [577]. Probability [541, 33, 856, 855]. Probe [772]. Problem [454, 667, 249, 418, 929, 142, 140, 73, 653, 313, 494, 465, 738, 774, 673, 881, 190, 220, 2, 438, 265, 462, 972, 549, 672, 134, 591, 742, 417, 628, 765, 851, 75, 295, 156, 824, 579, 9, 264, 986, 785, 915, 144, 346, 987, 51]. Problems [83, 566, 721, 527, 233, 282, 839, 547, 436, 271, 382, 334, 340, 960, 982, 831, 609, 42, 752, 888, 889]. Procedure [524, 435]. Process [532, 755, 211, 982, 617, 896]. Processes [97, 72, 791, 755, 126, 187]. Processing [267, 788, 445]. Processor [59]. Processor-Sharing [59]. Production [666, 755]. Productivity [562]. Profile [402, 800]. Profiles [735, 771, 647, 69]. Progressive [977, 391]. Projection [740]. Projections [512]. Projective [506]. Proliferation [767]. Proof [649]. Propagating [521, 163, 31, 32]. Propagation [503, 994, 311, 859, 940, 756, 14, 177, 320, 197, 109, 147, 784, 597, 651, 165, 300, 366, 306]. Propagator [317]. Properties [747, 161, 536, 683, 129, 179, 588, 419, 230, 251, 32]. Property [542, 874]. Prospecting [538, 678]. Protein [818, 127, 38]. Proton [949]. Pseudodifferential [176]. Public [604]. Pull [588]. Pull-In [588]. Pulmonary [294]. Pulsatile [682]. Pulse [311, 141, 182, 756, 601, 363]. Pulses [521, 576, 128, 200, 201]. Pumpkin [513]. Pure [625, 242]. Put [270, 467].

Quadratic [28, 859]. Quadrature [439]. Qualitative [875, 619]. Quality [207]. Quantities [416]. Quantized [757]. Quantum [600, 677]. Quasi [395, 688, 993, 951, 500]. Quasi-Birth-Death [500]. Quasi-geostrophic [395]. Quasi-Potential [951]. Quasi-Steady [993]. Quasi-Steady-State [688]. Qubit [530]. Quenching [148]. Queue [679, 303]. Queue-Length [679]. Queues [500]. Queuing [616].

Rabies [681]. Radial [332, 449, 348]. Radially [687]. Radiation [176, 961, 401, 483]. Radiative [83, 162, 716, 772, 404, 440, 593, 680].

Radical [117, 365]. Radio [362]. Radio-Frequency [362]. Radiotherapy [703]. Radon [22, 512, 534]. Random [586, 806, 989, 844, 67, 890, 762, 162, 374, 342, 277, 458, 756, 837, 98, 879, 66, 496, 897, 187, 175, 803, 349, 363]. Randomly [784]. Range [897, 910]. Rapid [262]. Rapidly [55, 518]. Rarefaction [881]. Rarefied [672]. Rascle [223, 779, 704, 917]. Ratchet [127]. Rate [325, 150, 219, 876, 508]. Rates [892]. Ratio [743, 745, 947, 508]. Ratio-Dependent [743, 745, 947, 508]. Ray [191, 8, 660]. Rayleigh [142, 627, 153]. Reaction [747, 330, 859, 542, 634, 664, 87, 954, 713, 499, 229, 35, 934, 71, 935, 351, 893, 269, 247]. Reaction-Diffusion [330, 499, 229, 35, 934, 71, 351, 247]. Reactions [327, 63, 728]. Reactive [142, 494]. Reactor [245, 266]. Read [730]. Real [929]. Realizable [328]. Receptor [552]. Receptors [818]. Recombination [117, 928, 730]. Reconciling [246]. Reconstructing [809]. Reconstruction [195, 569, 402, 737, 787, 771, 727, 52, 760, 81]. Reconstructions [955]. Recovering [589]. Recovery [456, 981, 763, 137, 212]. Rectangular [128, 46]. Rectilinear [471]. Recurrent [633, 450]. Recursive [878]. Recycling [1, 674]. Reduced [245]. Reduction [93, 396, 397, 950]. Redux [431]. Reference [923]. Referenced [262]. Reflected [801]. Reflection [574, 992, 66, 288]. Reflectionless [46]. Reflectivity [60]. Refocusing [349]. Refraction [988]. Regime [886, 467]. Regime-Switching [886]. Regimes [301]. Region [918, 384, 378, 469]. Regions [263, 726, 983, 108, 749, 433, 434]. Registration [207, 408, 671, 448]. Regular [992, 243, 943]. Regularization [438, 549, 534]. Regularized [555, 121, 609]. Regulated [884]. Regulation [693, 29]. Regulatory [711]. Reinforced [763, 489].

Related [721, 417, 987, 842]. Relation [642]. Relations [194, 539, 878]. Relationship [625]. Relationships [557]. Relaxation [299, 78, 139, 199, 115, 69, 455]. Relaxations [107]. Release [109]. Removal [690]. Removing [805]. Renormalization [210, 287, 302]. Repeated [272]. Replicator [187]. Representations [559]. Reproduction [938]. Reproductive [135]. Rescue [688]. Reservation [577, 913]. Reservoir [562]. Residual [529, 595]. Resilience [766]. Resistivity [669]. Resolvent [504, 318]. Resonance [488, 497, 14, 172, 572, 734, 440, 631, 394, 154]. Resonances [523, 406, 849, 284]. Resonant [101, 416]. Resonantly [234]. Resource [295, 384, 378]. Resources [91, 206, 485]. Respiratory [604]. Response [604, 340, 158, 985, 612, 411, 312]. Restoration [152, 637, 780]. Restricted [367]. Results [390, 290]. Resurrection [41]. Retarded [165]. Retrieving [560]. Retrofocusing [464]. Return [662]. Revealing [564]. Reversal [712, 887, 349, 458, 756, 425, 464, 429, 835]. Reversible [494]. Revisited [681, 631, 563]. Reynolds [164]. Rheology [423]. Riemann [418, 494, 881, 824, 785, 915]. Rigid [630, 724, 781]. Rigidity [489]. Rigorous [292, 49]. Ring [388, 601, 165]. Ripening [868]. Risk [325, 506]. Risk-Sensitivity [325]. River [723]. Robust [568]. Robustness [893]. Rod [630, 580, 498]. Rods [507, 45]. Role [38, 614, 871]. Roll [107, 731]. Roll-Waves [731]. Root [317]. Roots [272]. Rossby [553]. Rotating [610, 329, 283]. Rotation [787]. Rotation-Based [787]. Rotational [501]. Rotationally [468]. Rough [620]. Roughness [696]. Route [615]. Rubberlike [241]. Rule [463]. Rules [886]. Runner [978]. Running [978]. Rupture Saddle [852]. Saddle-Node [852]. Saffman [50]. Saint [279]. Saint-Venant [279]. Sampling [12, 737, 340]. Sandwiched [907]. Satisfying [874]. Saturation [642, 982]. Scalar [962, 479, 785, 157]. Scale [13, 500, 258]. Scales [117, 847]. Scaling [194]. Scan [487]. Scanner [848]. Scanning [927, 242]. Scatterer [737]. Scatterers [887, 990, 424]. Scattering [962, 566, 629, 620, 298, 937, 162, 916, 826, 484, 602, 284, 177, 726, 649, 733, 727, 340, 773, 392, 582, 473, 122, 169, 463, 846, 409, 74, 775, 720, 895]. Scheme [534]. Schemes [845]. School [558]. School-Oriented [558]. Schrödinger [900, 600, 415, 226, 342, 11, 168, 462, 741, 722, 613, 43, 757, 979]. Scott [141, 182, 100, 254]. Sea [701, 702, 720]. Sea-Ice [720]. Sea-Ice/Ice-Shelf [720]. Search [250]. Searching [132]. Seasonal [645]. Second [586, 41, 834, 574, 172, 399]. Second-Harmonic [574]. Secondary [53, 405]. Section [192, 744]. Sectional [915]. Sections [872]. Sedimentation [380, 515]. Segel [587, 433, 434]. Segmentation [377, 399, 732, 648, 866]. Segregation [643]. SEIR [193]. Seismic [473, 535]. Selection [928, 368, 164, 351, 725]. Selective [887, 425]. Selectivity [721]. Self [61, 117, 452, 44, 226, 457, 11, 168, 253, 29, 669, 283, 288, 825, 979]. Self-Adaptive [669]. Self-Adjusting [117]. Self-Aligning [452]. Self-Focusing [11, 168, 253]. Self-Regulation [29]. Self-Similar [226, 457, 283, 288, 825, 979]. Self-Similarity [44]. Self-Sustained [61]. Selling [886]. Semi [575, 907, 424, 733]. Semi-Infinite [907, 424, 733]. Semianalytical [644]. Semiclassical [463]. Semiconductor [600, 869, 383, 847, 828, 298, 285, 447, 186]. Semiconductors [99, 730, 768, 963, 7, 455]. Semicontinuous [444]. Semilinear [457]. Sensitivities [924]. Sensitivity [325, 587, 893]. Separation [925, 544, 145]. Separatrix [264]. Sequence [259, 788]. Sequencing [803]. Sequential [105]. Series [473, 58]. Serpentine [427]. Service [303]. Services [750]. Set [671, 267, 850]. Sets [993]. Severe [604]. Severity [604]. Severity-Dependent [604]. Shading [16]. Shadow [918]. Shafts [489]. Shah [671]. Shape [377, 580, 853, 399, 727, 914, 384, 378, 589, 709]. Shaped [513, 982, 781]. Shaping [581]. Shared [295, 384, 378]. Sharing [59]. Sharp [202, 553]. Sharp-Interface [202]. Sharpening [310]. Shaw [208]. Shear [279, 238, 82, 265, 658, 921, 522, 883]. Shearable [45]. Sheath [362]. Sheet [313]. Shelf [720]. Shell [972]. Shielding [217]. Shift [452]. Shift-Twist [452]. Shock [657, 797, 881, 23, 124, 288, 69]. Shock-Generated [124]. Shockley [730]. Shocks [823]. Short [883]. Shrinkage [207]. Si [963]. Sidewall [501]. Signal [18]. Signaling [941]. Signals [281]. Similar [226, 457, 283, 143, 288, 825, 979]. Similarity [44, 421]. Simple [746, 291, 970, 308, 946]. Simulation [712, 333, 218, 420, 397, 953, 321, 669, 617, 896]. Simulations [600, 15, 772, 877]. Simultaneous [303]. Single [838, 13, 65, 617]. Single-Wave-Form [65]. Singular [21, 456, 358, 436, 975, 509, 219, 146, 433, 434]. Singular-Perturbation [219]. Singularities [88, 208, 943]. Singularity [891, 590, 971, 491]. Singularity-Free [891]. Singularly [660, 130]. Sintering [833]. SIRS [554, 876]. SIS [735, 645]. Sites [70, 109]. Sivashinsky [343]. Sixth [863]. Sixth-Order [863]. Skeletal [802, 77]. Skew-Gradient [502]. Slabs [284, 394, 708]. Slender [663, 729, 744, 192]. Slightly [611]. Slip [315]. Slipping [884]. Slow [117, 572, 662, 214]. Slowly [234, 141, 182, 687, 369]. Slyozov [255]. Smale [794]. Small [195, 569, 887, 758, 668, 782, 642, 587, 584, 909, 934, 983, 149, 941]. Small- [758, 642]. Smectic [48, 885, 670]. Smooth [26, 359, 537]. Smoother [157]. Smoothing [157]. Smoothness [456]. SNIPER [976]. Soft [906]. Sog [552]. Soil [638]. Solenoidal [559]. Solid [27, 73, 530, 763, 952, 853, 550, 971, 498, 615, 269]. Solid-Oxide [952]. Solidification [276, 274]. Solids [3, 371]. Solitary [125, 113, 828, 352, 43]. Solitons [415, 741, 322, 316, 440]. Solutal [337]. Solute [817]. Solution [418, 660, 839, 220, 658, 462, 35, 30, 295, 302, 46, 89, 919, 786]. Solutions [606, 102, 595, 618, 198, 19, 226, 808, 924, 188, 221, 276, 782, 797, 516, 3, 141, 182, 167, 403, 467, 421, 483, 100, 202, 416, 275, 863, 359, 537, 139, 65, 146, 156, 827, 623, 982, 469, 20, 288, 825, 809, 31, 32, 346, 942, 979, 96]. Solving [811, 382]. Some [161, 188, 536, 282, 694, 982, 260]. Sonic [592]. Sorption [638]. Sound [101, 533, 162, 749, 366]. Source [249, 191, 185, 738, 798, 334, 851, 731, 769, 173, 752, 942]. Sources [807, 326]. Space [962, 432, 954, 804, 305, 196, 293]. Space-Clamped [305]. Spaced [151]. Spacers [426]. Spaces [534]. Spatial [19, 694, 40, 138, 495, 681, 450, 110]. Spatially [618, 116, 327, 525, 200, 201, 306]. Spatio [765]. Spatio-temporal [765]. Spatiotemporal [740, 120]. SPDEs [924]. Species [634, 229, 838, 206]. Species-Reaction [634]. Specific [47, 759]. Spectra [911, 181, 442, 145]. Spectral [179, 650, 790, 746, 579, 986, 938]. Spectrum [880, 717, 940]. Speed [503, 593]. Speeds [197]. Sphere [388, 944, 215, 568, 611, 398]. Spheres [962, 151, 724, 649]. Spherical [811, 778, 687, 972, 46, 534]. Spherically [623]. Spheroid [595]. Spike [37, 35, 254, 511, 247]. Spiky [866]. Spin [869, 530]. Spindle [632, 781].

Spindle-Shaped [781]. Spines [113, 519]. Spiral [278]. Splashing [274]. Splitting [182]. Sponge [46]. Spread [499, 694, 681]. Spreading [160, 363, 832]. Square [317]. Square-Root [317]. Stabilities [153]. Stability [375, 102, 830, 675, 503, 747, 388, 828, 68, 141, 857, 690, 338, 483, 903, 441, 693, 926, 254, 156, 429, 316, 691, 623, 883, 1, 674, 123, 796, 511, 585, 269, 31, 32, 260, 832, 987, 919, 115, 43, 84]. Stabilization [526, 774]. Stabilizes [138]. Stabilizing [871]. Stable [310, 845, 891, 150]. Stack [953]. Stage [375, 635, 210, 875, 739, 624, 659, 665]. Stage-Structured [659, 624]. Stagnation [355, 331]. Stagnation-Point [331]. Standing [470, 234, 201]. State [530, 688, 98, 509, 993, 835]. State-Mediated [98]. States [735, 394]. Static [85, 254]. Statics [714]. Stationary [808, 483, 547, 462, 863, 211, 745, 947, 607]. Statistical [771, 429]. Statistics [741, 872]. Steady [735, 639, 688, 802, 164, 379, 509, 993, 691, 884]. Steady-State [509]. Stefan [549, 144, 51]. Stem [693]. Stents [514]. Step [713, 820, 969]. Step-Function [713]. Stepped [910]. Stereographic [445]. Stimulus [567]. Stimulus-Locked [567]. Stochastic [67, 762, 17, 976, 755, 135, 396, 397, 693, 225, 264, 617, 104, 103, 189, 959]. Stoichiometric [443, 783]. Stokes [829, 194, 516, 891, 421, 888, 889]. Stones [451]. Stop [400]. Stopping [368]. Storage [838, 987]. Storm [224]. Straight [45]. Strain [238, 261, 920, 328, 517]. Strained [239]. Straining [162]. Strategies [370, 978]. Stratified [393]. Stream [531]. Strength [702, 334]. Stress [989, 644, 86, 920, 328, 459, 460]. Stress-Driven [459, 460]. Stresses [529, 595]. Stretching [350]. Strictly [176]. String [55]. Strong [657, 992, 121, 867]. Strongly [900, 380]. Structural [544]. Structure [375, 466, 430, 323, 332, 875, 698, 192, 739, 938]. Structured [558, 21, 694, 200, 201, 710, 659, 545, 624]. Structures [63, 849, 518, 582, 149, 283, 74, 775]. Studies [445]. Study [532, 772, 77, 943, 979]. Subcritical [214, 25]. Subdiffusion [935]. Subdiffusive [870]. Subject [136]. Subjected [196, 293]. Submanifolds [528]. Subspace [908]. Substrate [160, 851]. Subsurface [220]. Sufficient [691]. Sulfation [854]. Sulphur [451]. Supercavitating [973]. Supercharacteristic [107]. Superconducting [376, 92]. Superconductivity [53]. Superconductors [49]. Superdiffusion [858]. Superfluids [49]. Superlattice [298, 90]. Superlattices [869, 447]. Supersonic [983]. Supplemental [625]. Supply [616, 762]. Support [602, 206]. Suppression [879, 661, 816]. Supratransmission [852]. Surface [973, 705, 621, 125, 598, 663, 410, 510, 432, 931, 853, 916, 627, 733, 727, 647, 124, 209, 850, 912, 296, 799]. Surface-Tension-Dominated [125]. Surface-Tension-Gradient-Driven [647]. Surfaces [620, 175, 910, 684]. Surfactant [472, 646]. Surfactant-Covered [472]. Surfactants [133]. Suspension [423, 584]. Suspensions [515]. Sustained [61, 450]. Swarming [482]. Swelling [840]. Switching [886, 467]. Symmetric [561, 551]. Symmetries [740]. Symmetry [480, 452, 853, 520, 568, 154]. Symmetry-Breaking [853]. Synaptic [222, 478]. Synaptically [576, 200, 201]. Synchronization [857, 369, 171]. Synchronize [478, 143]. Synchronous [603]. Synchrony [493]. Syndrome [604]. System [900, 487, 618, 59, 73, 178, 255, 976, 679, 229, 547, 903, 135, 628, 693, 711, 743, 71, 750, 324, 356, 354, 136, 158, 407, 123, 689, 305, 247, 866, 919, 508, 612, 754, 312, 84]. Systems [176, 675, 747, 380, 188, 311, 859, 151, 958, 290, 561, 524, 214, 56, 107, 138, 441, 134, 397, 502, 509, 29, 65, 795, 935, 302, 146, 731, 336, 381, 371, 372, 351, 110, 104, 189, 585, 154, 985, 132, 237].

T [487, 719]. T-Scan [487]. Tailored [230]. Tanner [323]. Target [968, 934, 411]. Targets [621]. Taxis [936]. Taxis-Driven [936]. Taylor [68, 327, 759]. TE [784]. Technique [347, 387, 62]. Techniques [10, 402, 849, 250, 957]. Telecommunication [807]. Temperature [724, 672, 983, 219, 242, 651, 867]. Temperature-Jump [672]. Temperatures [568]. temporal [765]. Tensile [241, 196, 293]. Tension [125, 663, 350, 647, 871, 598]. Tension-Driven [663]. Tensor [479, 912]. Tensors [339]. Terms [806, 606, 731, 594]. Terrain [324]. Terrain-Following [324]. Test [340]. Testing [277]. Tests [834]. Texture [456, 684]. Thalamic [163]. Their [682]. Theorems [215, 613]. Theoretic [933, 120]. Theoretically [278]. Theories [625]. Theory [443, 529, 595, 72, 430, 28, 37, 93, 629, 885, 8, 182, 148, 192, 447, 320, 790, 30, 496, 898, 722, 251, 592, 846, 538, 112, 877]. Therapy [710, 843]. Thermal [443, 337, 644, 301, 870, 970, 860]. Thermal-Diffusive [970]. Thermoelasticity [619]. Thermomechanical [307, 308]. Thermophoresis [867]. Thermophoretic [611]. Thermotropic [54]. Theta [243]. Thickener [515]. Thickness [280]. Thin [245, 539, 140, 758, 218, 198, 653, 315, 257, 376, 540, 797, 664, 179, 908, 868, 849, 155, 790, 472, 582, 280, 898, 149, 474, 646, 165, 736, 895]. Thin-Film [758, 179]. Third [520]. Three [904, 620, 541, 742, 271, 974, 957, 888, 889]. Three-Dimensional [541, 271, 974, 957, 888, 889]. Thresholds [367]. Throughflow [142]. Tilted [38]. Time [922, 373, 712, 887, 962, 349, 758, 117, 845, 221, 944, 642, 638, 458, 906, 368, 756, 425, 464, 495, 429, 835, 831, 123, 938, 985, 645, 130, 500, 943]. Time-Dependent [373, 221]. Time-Domain [906]. Time-Harmonic [425, 831]. Time-Local [845]. Timing [323, 795]. Tissue [170, 77]. Tissues [529, 595, 772]. Tomographic [787, 541]. Tomography [836, 873, 804, 923, 419, 726, 718, 812, 734, 955, 821]. Tongue [488]. Top [225]. Topography [969]. Topological [560]. Torsional [489, 827]. Total [555, 438, 416]. Touchdown [695, 588]. Toy [980]. Trace [589]. Tracking [120]. Tracks [224]. Traction [667, 685, 912]. Traction-Free [685]. Traffic [292, 41, 297, 314, 223, 400, 76, 303, 139, 779, 704, 626, 796, 353, 617, 786]. Trafficking [814]. Train [37]. Trans [848]. Trans-Admittance [848]. Transaction [216]. Transfer [83, 162, 716, 932, 605, 404, 827, 345, 680]. Transform [512, 22, 516, 776, 534, 965, 995]. Transformation [951]. Transformations [307, 308]. Transforms [248]. Transient [673, 171, 235, 501, 409, 942]. Transition [48, 114, 335, 732, 720]. Transitions [37, 330, 314, 57, 149, 651, 599, 84]. Transmission [406, 574, 676, 92, 476, 193, 946]. Transonic [881]. Transparent [726]. Transport [900, 99, 72, 263, 844, 869, 17, 64, 231, 906, 837, 570, 817, 772, 126, 770, 862, 479, 761, 244, 199, 548, 77]. Transport-Based [844]. Transversely [432, 760]. Trapping [258]. Traps [818, 944]. Travel [281]. Traveling [793, 375, 415, 618, 19, 315, 332, 823, 641, 167, 567, 537, 23, 243, 200, 20, 585]. Travelling [359, 75, 351]. Treatment [292]. Tricritical [933]. Triple [952, 689, 825]. Tritrophic [213]. Trunk [577, 913]. Tube [234, 272, 350]. Tuberculosis [261]. Tumor [20, 754]. Tumors [808]. Tunnels [673].

Turbulence [333, 335, 225, 393]. Turbulent [162, 374, 364, 287, 479, 652, 759, 304, 291]. Turing [858, 229, 502, 112]. Twist [452]. Two [569, 962, 635, 323, 19, 490, 185, 252, 141, 182, 683, 792, 698, 448, 261, 625, 86, 916, 679, 662, 284, 229, 265, 903, 164, 205, 303, 913, 271, 476, 907, 91, 206, 227, 711, 219, 898, 517, 669, 978, 291, 283, 123, 482, 74, 485, 500, 258]. Two-Component [683, 792]. Two-Dimensional [569, 284, 903, 164, 205, 913, 271, 476, 227, 898, 669, 291, 283, 482, 74, 252]. Two-Phase [698, 258]. Two-Point [265]. Two-Priority [679]. Two-Pulse [141, 182]. Two-Runner [978]. Two-Species [229]. Two-Stage [635]. Two-Strain [261, 517]. Two-Temperature [219]. Two-Time-Scale [500]. Two-Timing [323]. Type [140, 25, 30, 278, 863, 417, 743, 624, 1, 674, 651, 985, 644]. Types [303].

Ultrasonic [420]. Unambiguous [262]. Unbounded [533, 620]. Unbuffered [295]. Undercooled [144]. Undercooling [774, 549]. Underground [987]. Underloaded [577]. Undersampling [520]. Unequal [872]. Unexpectedly [95]. Uniaxial [86]. Unified [424]. Uniform [3, 58, 317]. Uniformity [874]. Uniqueness [839, 65, 955]. Units [515]. Universal [524]. Unperturbed [11]. Unsteady [898]. Unstirred [485]. Uplink [750]. Upper [225]. Upscaling [594]. Urn [187]. Use [310]. Used [368]. Using [512, 377, 600, 323, 744, 12, 425, 462, 677, 975, 935, 669, 809, 951].

V [644]. Vaccination [390]. Valuation [967]. Value [536, 403, 265, 436, 75, 42, 346]. Valued [683, 780]. Values [496]. Vanishing [88]. Vapor [57, 387]. Vaporization [824]. Variable [637, 192]. Variables [662, 145]. Variation [555, 86, 438, 416]. Variational [10, 805, 981, 152, 408, 780, 2, 197, 391]. Variations [377]. Varying [116, 588, 687]. Vector [954, 780, 694, 559, 302, 851]. Vector-Space [954]. Vehicular [76, 353]. Velocimetry [771]. Velocity [126, 771, 800]. Velocity-Jump [126]. Venant [279]. Ventilation [170]. Versatile [252]. Vertical [193, 133]. via [922, 10, 968, 925, 737, 787, 541, 732]. Vibration [905]. Vicinal [296]. Video [453]. View [430]. Virulent [843]. Virus [341, 690, 902, 814, 894, 832]. Visco [368]. Visco-Plastic [368]. Viscoelastic [61, 423, 823, 682, 256, 763]. Viscoelasticity [382]. Viscometry [262]. Viscoplastic [184]. Viscosity [88, 687, 84]. Viscosity-Capillarity [84]. Viscous [423, 598, 218, 729, 215, 495, 611, 725, 969, 131, 781]. Void [366]. Voids [817, 927]. Voltage [539, 911, 588, 831]. Volterra [375, 959, 919, 96]. Volume [668, 584, 975, 579]. Vortex [581, 533, 24, 568, 49, 800]. Vortical [283]. Vortices [388, 68, 757]. Vorticity [164].

Waals [657]. Waiting [758, 642]. Waiting-Time [758, 642]. Walks [98]. Wall [266, 983, 271, 472, 119, 70]. Walls [682, 551, 461]. Water [333, 916, 790]. Wave [523, 101, 240, 618, 333, 19, 845, 543, 167, 940, 698, 778, 14, 572, 422, 62, 879, 464, 964, 138, 66, 359, 537, 108, 765, 147, 760, 65, 883, 20, 351, 165, 689, 366, 720, 306, 317, 942]. Wave-Block [108]. Wavefields [191]. Wavefront [300]. Wavefronts [697]. Waveguide [569, 574, 961, 975, 147, 852]. Waveguides [756, 837, 784]. Wavelength [597]. Wavelet [248, 207, 559]. Wavelets [364, 120]. Wavenumber [649]. Waves [793, 101, 466, 962, 566, 125, 333, 470, 315, 332, 823, 273, 113, 828, 763, 162, 361, 567, 458, 916, 36, 338, 400, 107, 109, 241, 352, 733, 773, 719, 23, 897, 75, 977, 731, 212, 243, 553, 498, 365, 9, 209, 725, 917, 585, 748, 43].

Wavy [486]. Way [317]. Weak [630, 746, 288, 748, 860]. Weakly [28, 576, 572, 550, 302, 898]. Webster [563]. Wedge [973, 598, 663, 685]. Wedges [347]. Weinstein [986]. Well [922, 675, 448, 562]. Well-Posedness [922, 448, 675]. West [902]. Wet [984]. Which [478]. While [669]. White [336]. Wide [786]. Wiener [347]. Willshaw [855]. Wilson [354]. Wind [994, 4, 209]. Wind-Driven [4]. Windshield [480]. Windshields [417]. Wire [668]. Without [923, 450]. WKB [272]. Wood [841]. Working [289].

Yawed [973]. Yield [86].

Zero [139]. Zones [664].

References

Ruan:1998:GSC

[1] Shigui Ruan and Xue-Zhong He. Global stability in chemostat-type competition models with nutrient recycling. SIAM Journal on Applied Mathematics, 58(1):170–192, February 1998. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/29924. See erra-

tum [674].

Gueron:1999:DVP

[2] Shay Gueron and Itai Shafrir. On a discrete variational problem involving interacting particles. SIAM Journal on Applied Mathematics, 60(1):1– 17, 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/31525.

Dascalu:1999:UAS

[3] Cristian Dascalu and Dorel Homentcovschi. Uniform asymptotic solutions for lamellar inhomogeneities in anisotropic elastic solids. SIAM Journal on Applied Mathematics, 60(1):18– 42, 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/33720.

Desjardins:1999:HMW

[4] B. Desjardins and E. Grenier. On the homogeneous model of wind-driven ocean circulation. SIAM Journal on Applied Mathematics, 60(1):43–60, 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/32426.

Herrero:1999:MDB

[5] Henar Herrero, Brigitte Lucquin-Desreux, and Benoît Perthame. On the motion of dispersed balls in a potential flow: a kinetic description of the added mass effect. SIAM Journal on Applied Mathematics, 60(1):61– 83, 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/32971.

Belyaev:1999:EMP

[6] Alexander G. Belyaev, Gregory A. Chechkin, and Rustem R. Gadyl’shin. Effective membrane permeability: Estimates and low concentration asymptotics. SIAM Journal on Applied Mathematics, 60(1):84–108, 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/31288.

Schmeiser:1999:MMA

[7] Christian Schmeiser. Macroscopic models for almost elastic nonlinear electron-phonon interaction in semiconductors. SIAM Journal on Applied Mathematics, 60(1):109–120, 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/33757.

Chapman:1999:RTH

[8] S. J. Chapman, J. M. H. Lawry, and J. R. Ockendon. Ray theory for high-Péclet-number convectiondiffusion. SIAM Journal on Applied Mathematics, 60(1):121–135, 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/34408.

Schult:1999:MAE

[9] Daniel A. Schult. Matched asymptotic expansions and the closure problem for combustion waves. SIAM Journal on Applied Mathematics, 60(1):136– 155, 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/33715.

Aubert:1999:COF

[10] G. Aubert, R. Deriche, and P. Kornprobst. Computing optical flow via variational techniques. SIAM Journal on Applied Mathematics, 60(1):156– 182, 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/34017.

Fibich:1999:SFP

[11] Gadi Fibich and George Papanicolaou. Self-focusing in the perturbed and unperturbed nonlinear Schrödinger equation in critical dimension. SIAM Journal on Applied Mathematics, 60(1):183– 240, 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/32240.

Colton:1999:LSM

[12] David Colton and Peter Monk. A linear sampling method for the detection of leukemia using microwaves II. SIAM Journal on Applied Mathematics, 60 (1):241–255, 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/34426.

Meister:1999:ASM

[13] Andreas Meister. Asymptotic single and multiple scale expansions in the low Mach number limit. SIAM Journal on Applied Mathematics, 60(1):256– 271, 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/34319.

Georgieva:1999:WPR

[14] Anna Georgieva, Thomas Kriecherbauer, and Stephanos Venakides. Wave propagation and resonance in a onedimensional nonlinear discrete periodic medium. SIAM Journal on Applied Mathematics, 60(1):272–294, 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/34031.

Garcke:1999:MFC

[15] Harald Garcke, Britta Nestler, and Barbara Stoth. A multiphase field concept: Numerical simulations of moving phase boundaries and multiple junctions. SIAM Journal on Applied Mathematics, 60(1):295–315, 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/33489.

Ackleh:1999:NPA

[16] Azmy S. Ackleh and Robert R. Ferdinand. A nonlinear phytoplankton aggregation model with light shading. SIAM Journal on Applied Mathematics, 60 (1):316–336, 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/33343.

Dehling:1999:SMT

[17] H. G. Dehling, A. C. Hoffmann, and H. W. Stuut. Stochastic models for transport in a fluidized bed. SIAM Journal on Applied Mathematics, 60(1):337– 358, 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/30631.

Ahn:1999:OFG

[18] Hyungsok Ahn and Raisa E. Feldman. Optimal filtering of a Gaussian signal in the presence of Lévy noise. SIAM Journal on Applied Mathematics, 60(1):359– 369, 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/31200.

Brazhnik:1999:TWS

[19] Pavel K. Brazhnik and John J. Tyson. On traveling wave solutions of Fisher’s equation in two spatial dimensions. SIAM Journal on Applied Mathematics, 60(2):371–391, 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/32549.

Sherratt:1999:TWS

[20] Jonathan A. Sherratt. Traveling wave solutions of a mathematical model for tumor encapsulation. SIAM Journal on Applied Mathematics, 60(2):392– 407, 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/34535.

Arino:1999:SPA

[21] Ovide Arino, Eva Sánchez, Rafael Bravo de la Parra, and Pierre Auger. A singular perturbation in an age-structured population model. SIAM Journal on Applied Mathematics, 60(2):408–436, 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/33796.

Brandt:1999:FAM

[22] Achi Brandt, Jordan Mann, Matvei Brodski, and Meirav Galun. A fast and accurate multilevel inversion of the radon transform. SIAM Journal on Applied Mathematics, 60(2):437– 462, 1999. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/32425.

Marchant:2000:TSW

[23] B. P. Marchant, J. Norbury, and A. J. Perumpanani. Traveling shock waves arising in a model of malignant invasion. SIAM Journal on Applied Mathematics, 60(2):463–476, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/32803.

Cohen:2000:MAV

[24] Albert Cohen and Raphael Danchin. Multiscale approximation of vortex patches. SIAM Journal on Applied Mathematics, 60(2):477–502, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/31978.

Izhikevich:2000:SEB

[25] Eugene M. Izhikevich. Subcritical elliptic bursting of Bautin type. SIAM Journal on Applied Mathematics, 60(2):503– 535, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/33263.

Babich:2000:EDC

[26] Vassilii M. Babich, Daniel B. Dement’ev, Boris A. Samokish, and Valery P. Smyshlyaev. On evaluation of the diffraction coefficients for arbitrary ‘nonsingular’ directions of a smooth convex cone. SIAM Journal on Applied Mathematics, 60(2):536– 573, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/33366.

Billingham:2000:AML

[27] J. Billingham, A. C. King, R. C. Copcutt, and K. Kendall. Analysis of a model for a loaded, planar, solid oxide fuel cell. SIAM Journal on Applied Mathematics, 60(2):574– 601, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/34064.

Boertjens:2000:ATW

[28] G. J. Boertjens and W. T. van Horssen. An asymptotic theory for a weakly nonlinear beam equation with a quadratic perturbation. SIAM Journal on Applied Mathematics, 60(2):602– 632, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/33676.

Lubashevsky:2000:CMS

[29] I. A. Lubashevsky and V. V. Gafiychuk. Cooperative mechanism of self-regulation in hierarchical living systems. SIAM Journal on Applied Mathematics, 60(2):633–663, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/27917.

Karlin:2000:NSN

[30] V. Karlin, V. G. Maz’ya, A. B. Movchan, J. R. Willis, and R. Bullough. Numerical solution of nonlinear hypersingular integral equations of the Peierls type in dislocation theory. SIAM Journal on Applied Mathematics, 60(2):664– 678, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/31162.

Vaynblat:2000:SPSa

[31] Dimitri Vaynblat and Moshe Matalon. Stability of pole solutions for planar propagating flames: I. exact eigenvalues and eigenfunctions. SIAM Journal on Applied Mathematics, 60(2):679– 702, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/34643.

Vaynblat:2000:SPSb

[32] Dimitri Vaynblat and Moshe Matalon. Stability of pole solutions for planar propagating flames: II. properties of eigenvalues/eigenfunctions and implications to stability. SIAM Journal on Applied Mathematics, 60(2):703– 728, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/34644.

Knessl:2000:PBM

[33] Charles Knessl and Joseph B. Keller. Probability of Brownian motion hitting an obstacle. SIAM Journal on Applied Mathematics, 60(2):729–745, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/34627.

Chen:2000:ACM

[34] Zhangxin Chen, Guan Qin, and Richard E. Ewing. Analysis of a compositional model for fluid flow in porous media. SIAM Journal on Applied Mathematics, 60(3):747–777, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/33342.

Iron:2000:MSS

[35] David Iron and Michael J. Ward. A metastable spike solution for a nonlocal reaction-diffusion model. SIAM Journal on Applied Mathematics, 60(3):778– 802, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/33834.

Grote:2000:ENB

[36] Marcus J. Grote and Joseph B. Keller. Exact nonreflecting boundary condition for elastic waves. SIAM Journal on Applied Mathematics, 60(3):803– 819, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/34422.

Bressloff:2000:DTS

[37] P. C. Bressloff and S. Coombes. A dynamical theory of spike train transitions in networks of integrate-and-fire oscillators. SIAM Journal on Applied Mathematics, 60(3):820–841, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/33964.

Elston:2000:RPF

[38] Timothy C. Elston and Charles S. Peskin. The role of protein flexibility in molecular motor function:

Coupled diffusion in a tilted periodic potential. SIAM Journal on Applied Mathematics, 60(3):842–867, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/34534.

Ruuth:2000:CGM

[39] Steven J. Ruuth and Barry Merriman. Convolution-generated motion and generalized Huygens’ principles for interface motion. SIAM Journal on Applied Mathematics, 60(3):868– 890, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/33397.

Juang:2000:CNN

[40] Jonq Juang and Song-Sun Lin. Cellular neural networks: Mosaic pattern and spatial chaos. SIAM Journal on Applied Mathematics, 60(3):891– 915, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/32360.

Aw:2000:RSO

[41] A. Aw and M. Rascle. Resurrection of ‘second order” models of traffic flow. SIAM Journal on Applied Mathematics, 60(3):916–938, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/33209.

Steinberg:2000:MAH

[42] Ben Zion Steinberg, John J. McCoy, and Mark Mirotznik. A multiresolution approach to homogenization and effective modal analysis of complex boundary value problems. SIAM Journal on Applied Mathematics, 60(3):939–966, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/32369.

Yang:2000:SES

[43] Jianke Yang and D. J. Kaup. Stability and evolution of solitary waves in perturbed generalized nonlinear Schrödinger equations. SIAM Journal on Applied Mathematics, 60(3):967– 989, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/34189.

Budd:2000:ASS

[44] Chris J. Budd and Mark A. Peletier. Approximate self-similarity in models of geological folding. SIAM Journal on Applied Mathematics, 60(3):990– 1016, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/33714.

Schuricht:2000:SCS

[45] Friedemann Schuricht. Straight configurations of shearable nonlinearly elastic rods. SIAM Journal on Applied Mathematics, 60(3):1017–1036, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/33701.

Petropoulos:2000:RSL

[46] Peter G. Petropoulos. Reflectionless sponge layers as absorbing boundary conditions for the numerical solution of Maxwell equations in rectangular, cylindrical, and spherical coordinates. SIAM Journal on Applied Mathematics, 60(3): 1037–1058, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/33468.

Fister:2000:OCA

[47] K. Renee Fister and John Carl Panetta. Optimal control applied to cell-cycle-specific cancer chemotherapy. SIAM Journal on Applied Mathematics, 60(3): 1059–1072, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/33850.

Calderer:2000:EBM

[48] M. Carme Calderer and Peter Palffy-Muhoray. Ericksen’s bar and modeling of the smectic A–nematic phase transition. SIAM Journal on Applied Mathematics, 60(3):1073–1098, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/32756.

Lin:2000:RGD

[49] Tai-Chia Lin. Rigorous and generalized derivation of vortex line dynamics in superfluids and superconductors. SIAM Journal on Applied Mathematics, 60(3): 1099–1110, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/34188.

Jager:2000:IBC

[50] Willi Jäger and Andro Mikelic. On the interface boundary condition of Beavers, Joseph, and Saffman. SIAM Journal on Applied Mathematics, 60(4):1111– 1127, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/33678.

vandeFliert:2000:GSP

[51] B. W. van de Fliert and R. van der Hout. A generalized Stefan problem in a diffusion model with evaporation. SIAM Journal on Applied Mathematics, 60(4): 1128–1136, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/32759.

March:2000:RPC

[52] Riccardo March and Piero Barone. Reconstruction of a piecewise constant function from noisy Fourier coefficients by Padé method. SIAM Journal on Applied Mathematics, 60(4):1137–1156, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/33384.

Aftalion:2000:AAS

[53] A. Aftalion and S. J. Chapman. Asymptotic analysis of a secondary bifurcation of the one-dimensional Ginzburg–Landau equations of superconductivity. SIAM Journal on Applied Mathematics, 60(4):1157–1176, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/34479.

Forest:2000:TLC

[54] M. Gregory Forest, Hong Zhou, and Qi Wang. Thermotropic liquid crystalline polymer fibers. SIAM Journal on Applied Mathematics, 60(4):1177– 1204, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/33677.

Castro:2000:LFA

[55] Carlos Castro and Enrique Zuazua. Low frequency asymptotic analysis of a string with rapidly oscillating density. SIAM Journal on Applied Mathematics, 60(4): 1205–1233, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/33063.

Heemels:2000:LCS

[56] W. P. M. H. Heemels, J. M. Schumacher, and S. Weiland. Linear complementarity systems. SIAM Journal on Applied Mathematics, 60(4):1234– 1269, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/32519.

Fan:2000:MDL

[57] Haitao Fan. On a model of the dynamics of liquid/vapor phase transitions. SIAM Journal on Applied Mathematics, 60(4): 1270–1301, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/34320.

deHoop:2000:UAE

[58] Maarten V. de Hoop and A. K. Gautesen. Uniform asymptotic expansion of the generalized Bremmer series. SIAM Journal on Applied Mathematics, 60(4): 1302–1329, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/32361.

Berger:2000:DPS

[59] Arthur Berger and Yaakov Kogan. Distribution of processor-sharing customers for a large closed system with multiple classes. SIAM Journal on Applied Mathematics, 60(4):1330–1339, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/34606.

Aktosun:2000:IRD

[60] Tuncay Aktosun and Paul E. Sacks. Inversion of reflectivity data for nondecaying potentials. SIAM Journal on Applied Mathematics, 60(4):1340– 1356, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/35558.

Antman:2000:SSO

[61] Stuart S. Antman and Herbert Koch. Self-sustained oscillations of nonlinearly viscoelastic layers. SIAM Journal on Applied Mathematics, 60(4):1357–1387, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/33795.

Hager:2000:WAT

[62] William W. Hager, Rouben Rostamian, and Dongxing Wang. The wave annihilation technique and the design of nonreflective coatings. SIAM Journal on Applied Mathematics, 60(4):1388– 1424, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/32409.

Edwards:2000:BRH

[63] D. A. Edwards. Biochemical reactions on helical structures. SIAM Journal on Applied Mathematics, 60(4):1425– 1446, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/34376.

Dumas:2000:HTE

[64] Laurent Dumas and François Golse. Homogenization of transport equations. SIAM Journal on Applied Mathematics, 60(4):1447–1470, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/33208.

Mirollo:2000:EUN

[65] Rennie Mirollo and Ned Rosen. Existence, uniqueness, and nonuniqueness of single-wave-form solutions to Josephson junction systems. SIAM Journal on Applied Mathematics, 60(5):1471– 1501, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/34385.

Kim:2000:ADL

[66] Jeong-Hoon Kim and So Young Sohn. An asymptotic diffusion limit for electromagnetic wave reflection from a random medium. SIAM Journal on Applied Mathematics, 60(5):1502–1519, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/34641.

Christakos:2000:SFA

[67] George Christakos, Dionissios T. Hristopulos, and Alexander Kolovos. Stochastic flowpath analysis of multiphase flow in random porous media. SIAM Journal on Applied Mathematics, 60(5):1520–1542, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/32054.

Dando:2000:MST

[68] A. H. Dando, P. Hall, and D. T. Papageorgiou. The modulational stability of Taylor vortices in a curved channel. SIAM Journal on Applied Mathematics, 60(5):1543–1564, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/31557.

Yong:2000:ERS

[69] Wen-An Yong and Kevin Zumbrun. Existence of relaxation shock profiles for hyperbolic conservation laws. SIAM Journal on Applied Mathematics, 60(5): 1565–1575, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/35270.

Jones:2000:MCN

[70] Don A. Jones and Hal Smith. Microbial competition for nutrient and Wall sites in plug flow. SIAM Journal on Applied Mathematics, 60(5):1576– 1600, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/34158.

Medvedev:2000:RDS

[71] Georgiy S. Medvedev, Tasso J. Kaper, and Nancy Kopell. A reactiondiffusion system with periodic front dynamics. SIAM Journal on Applied Mathematics, 60(5):1601–1638, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/34463.

Bal:2000:PTT

[72] Guillaume Bal, George Papanicolaou, and Leonid Ryzhik. Probabilistic theory of transport processes with polarization. SIAM Journal on Applied Mathematics, 60(5):1639–1666, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35696.

Bobula:2000:FMB

[73] E. Bobula, D. A. Tarzia, K. Twardowska, and L. T. Villa. On a free-moving boundary diffusion problem in a catalytic gas-solid system with catalyst decay. SIAM Journal on Applied Mathematics, 60(5):1667– 1685, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/30043.

Venakides:2000:BIC

[74] Stephanos Venakides, Mansoor A. Haider, and Vassilis Papanicolaou. Boundary integral calculations of twodimensional electromagnetic scattering by photonic crystal Fabry–Perot structures. SIAM Journal on Applied Mathematics, 60(5):1686–1706, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/35077.

McCabe:2000:ETW

[75] P. M. McCabe, J. A. Leach, and D. J. Needham. The evolution of travelling waves in fractional order autocatalysis with decay. II. the initial boundary value problem. SIAM Journal on Applied Mathematics, 60(5):1707– 1748, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/34477.

Klar:2000:KDM

[76] Axel Klar and Raimund Wegener. Kinetic derivation of macroscopic anticipation models for vehicular traffic. SIAM Journal on Applied Mathematics, 60(5): 1749–1766, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/35618.

Titcombe:2000:ASO

[77] Michèle S. Titcombe and Michael J. Ward. An asymptotic study of oxygen transport from multiple capillaries to skeletal muscle tissue. SIAM Journal on Applied Mathematics, 60(5):1767– 1788, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/34335.

Izhikevich:2000:PER

[78] Eugene M. Izhikevich. Phase equations for relaxation oscillators. SIAM Journal on Applied Mathematics, 60(5):1789– 1804, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35100.

Ammari:2000:JEC

[79] H. Ammari, A. Buffa, and J.-C. Nédélec. A justification of eddy currents model for the Maxwell equations. SIAM Journal on Applied Mathematics, 60(5): 1805–1823, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/34897.

Hueter:2000:ACO

[80] Irene Hueter and Sonya Bahar. Apparently chaotic orbits embedded in closed curves. SIAM Journal on Applied Mathematics, 60(5):1824–1840, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/32167.

Ravi:2000:NFR

[81] D. Ravi and Y. V. Venkatesh. A new framework for reconstruction of images from boundaries. SIAM Journal on Applied Mathematics, 60(6):1841– 1866, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/34386.

Flemming:2000:SLA

[82] R. P. Flemming, W. E. Olmstead, and S. H. Davis. Shear localization with an Arrhenius flow law. SIAM Journal on Applied Mathematics, 60(6):1867– 1886, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35054.

Bal:2000:DAR

[83] Guillaume Bal and Leonid Ryzhik. Diffusion approximation of radiative transfer problems with interfaces. SIAM Journal on Applied Mathematics, 60(6): 1887–1912, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/35208.

Zumbrun:2000:DSP

[84] Kevin Zumbrun. Dynamical stability of phase transitions in the p-system with viscosity-capillarity. SIAM Journal on Applied Mathematics, 60(6):1913– 1924, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/35269.

Calderer:2000:LCF

[85] M. Carme Calderer and Chun Liu. Liquid crystal flow: Dynamic and static configurations. SIAM Journal on Applied Mathematics, 60(6):1925– 1949, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/33624.

Frigaard:2000:EYS

[86] I. A. Frigaard and O. Scherzer. The effects of yield stress variation on uniaxial exchange flows of two Bingham fluids in a pipe. SIAM Journal on Applied Mathematics, 60(6):1950–1976, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/33516.

Evans:2000:DHR

[87] J. D. Evans and J. R. King. On the derivation of heterogeneous reaction kinetics from a homogeneous reaction model. SIAM Journal on Applied Mathematics, 60(6):1977–1996, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/34649.

Eggers:2000:SDP

[88] Jens Eggers. Singularities in droplet pinching with vanishing viscosity. SIAM Journal on Applied Mathematics, 60(6): 1997–2008, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/33488.

Sirovich:2000:DNP

[89] L. Sirovich, A. Omurtag, and B. W. Knight. Dynamics of neuronal populations: The equilibrium solution. SIAM Journal on Applied Mathematics, 60(6): 2009–2028, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/34492.

Moscoso:2000:BBS

[90] M. Moscoso, J. Galán, and L. L. Bonilla. Bifurcation behavior of a superlattice model. SIAM Journal on Applied Mathematics, 60(6):2029–2057, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/33675.

Li:2000:GAB

[91] Bingtuan Li, Gail S. K. Wolkowicz, and Yang Kuang. Global asymptotic behavior of a chemostat model with two perfectly complementary resources and distributed delay. SIAM Journal on Applied Mathematics, 60(6):2058– 2086, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35975.

Dhia:2000:MAC

[92] Anne-Sophie Bonnet-Ben Dhia and Karim Ramdani. Mathematical analysis of conducting and superconducting transmission lines. SIAM Journal on Applied Mathematics, 60(6):2087– 2113, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35242.

Broomhead:2000:NAD

[93] D. S. Broomhead and M. Kirby. A new approach to dimensionality reduction: Theory and algorithms. SIAM Journal on Applied Mathematics, 60(6):2114– 2142, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/33858.

Avellaneda:2000:FDA

[94] M. Avellaneda, L. Berlyand, and J.-F. Clouet. Frequency-dependent acoustics of composites with interfaces. SIAM Journal on Applied Mathematics, 60(6): 2143–2181, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/34281.

Sander:2000:ULB

[95] Evelyn Sander and Thomas Wanner. Unexpectedly linear behavior for the Cahn–Hilliard equation. SIAM Journal on Applied Mathematics, 60(6):2182– 2202, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35222.

vanHemmen:2000:CVK

[96] J. Leo van Hemmen, Werner M. Kistler, and Erik G. F. Thomas. Calculation of Volterra kernels for solutions of nonlinear differential equations. SIAM Journal on Applied Mathematics, 61(1):1– 21, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/33603.

Ambrosi:2000:MIM

[97] D. Ambrosi and L. Preziosi. Modeling injection molding processes with deformable porous preforms. SIAM Journal on Applied Mathematics, 61(1):22– 42, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/33367.

Grunbaum:2000:ADE

[98] Daniel Grünbaum. Advection-diffusion equations for internal state-mediated random walks. SIAM Journal on Applied Mathematics, 61(1):43–73, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/33207.

Anile:2000:EHM

[99] Angelo Marcello Anile, Vittorio Romano, and Giovanni Russo. Extended hydrodynamical model of carrier transport in semiconductors. SIAM Journal on Applied Mathematics, 61(1):74– 101, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/33294.

Hale:2000:EHH

[100] J. K. Hale, L. A. Peletier, and W. C. Troy. Exact homoclinic and heteroclinic solutions of the Gray–Scott model for autocatalysis. SIAM Journal on Applied Mathematics, 61(1):102– 130, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/33491.

Ali:2000:RIS

[101] Giuseppe Alì and John K. Hunter. The resonant interaction of sound waves with a large amplitude entropy wave. SIAM Journal on Applied Mathematics, 61 (1):131–148, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/34404.

Almog:2000:BSP

[102] Yaniv Almog. On the bifurcation and stability of periodic solutions of the Ginzburg–Landau equations in the plane. SIAM Journal on Applied Mathematics, 61(1):149– 171, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/35369.

Stevens:2000:SCA

[103] Angela Stevens. A stochastic cellular automaton modeling gliding and aggregation of myxobacteria. SIAM Journal on Applied Mathematics, 61(1):172– 182, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/34205.

Stevens:2000:DCE

[104] Angela Stevens. The derivation of chemotaxis equations as limit dynamics of moderately interacting stochastic many-particle systems. SIAM Journal on Applied Mathematics, 61(1):183– 212, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/34206. See erra-

tum [189].

Cruywagen:2000:EMA

[105] Gerhard C. Cruywagen, Philip K. Maini, and James D. Murray. An envelope method for analyzing sequential pattern formation. SIAM Journal on Applied Mathematics, 61(1):213– 231, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/32857.

Schonbucher:2000:FEH

[106] Philipp J. Schönbucher and Paul Wilmott. The feedback effect of hedging in illiquid markets. SIAM Journal on Applied Mathematics, 61(1):232– 272, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/30853.

Jin:2000:HSS

[107] Shi Jin and Markos A. Katsoulakis. Hyperbolic systems with supercharacteristic relaxations and roll waves. SIAM Journal on Applied Mathematics, 61 (1):273–292, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/35078.

Lewis:2000:WBE

[108] Timothy J. Lewis and James P. Keener. Wave-block in excitable media due to regions of depressed excitability. SIAM Journal on Applied Mathematics, 61 (1):293–316, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/34929.

Keener:2000:PWE

[109] J. P. Keener. Propagation of waves in an excitable medium with discrete release sites. SIAM Journal on Applied Mathematics, 61(1):317–334, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/35081.

Shih:2000:IBC

[110] Chih-Wen Shih. Influence of boundary conditions on pattern formation and spatial chaos in lattice systems. SIAM Journal on Applied Mathematics, 61 (1):335–368, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/34065.

Manoussaki:2000:EGF

[111] Daphne Manoussaki and Richard S. Chadwick. Effects of geometry on fluid loading in a coiled cochlea. SIAM Journal on Applied Mathematics, 61(2):369– 386, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35840.

Wollkind:2000:CTP

[112] David J. Wollkind and Laura E.

Stephenson. Chemical Turing pattern formation analyses: Comparison of theory with experiment. SIAM Journal on Applied Mathematics, 61(2):387– 431, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/32621.

Coombes:2000:SWM

[113] S. Coombes and P. C. Bressloff. Solitary waves in a model of dendritic cable with active spines. SIAM Journal on Applied Mathematics, 61(2):432– 453, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35660.

Chen:2000:PTN

[114] Xinfu Chen and Xiao-Ping Wang. Phase transition near a liquid-gas coexistence equilibrium. SIAM Journal on Applied Mathematics, 61(2):454–471, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35428.

Xin:2000:SRO

[115] Jack X. Xin and James M. Hyman.

Stability, relaxation, and oscillation of biodegradation fronts. SIAM Journal on Applied Mathematics, 61(2):472– 505, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/32696.

Dallon:2000:MMS

[116] John C. Dallon and Jonathan A. Sherratt. A mathematical model for spatially varying extracellular matrix alignment. SIAM Journal on Applied Mathematics, 61(2):506–527, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/35934.

Bonilla:2000:DHB

[117] Luis L. Bonilla, Antonio L. Sanchez, and Manuel Carretero. The description of homogeneous branched-chain explosions with slow radical recombination by self-adjusting time scales. SIAM Journal on Applied Mathematics, 61(2):528– 550, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/35943.

Helsing:2000:ENA

[118] Johan Helsing and Gunnar Peters. An efficient numerical algorithm for cracks partly in frictionless contact. SIAM Journal on Applied Mathematics, 61 (2):551–566, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/35693.

Stemmons:2000:CCW

[119] Eric D. Stemmons and Hal L. Smith. Competition in a chemostat with wall attachment. SIAM Journal on Applied Mathematics, 61(2):567–595, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/35813.

Leduc:2000:SWG

[120] J.-P. Leduc, F. Mujica, R. Murenzi, and M. J. T. Smith. SpatioTemporal wavelets: a group-theoretic construction for motion estimation and tracking. SIAM Journal on Applied Mathematics, 61(2):596–632, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/35186.

Nikolova:2000:LSH

[121] Mila Nikolova. Local strong homogeneity of a regularized estimator. SIAM Journal on Applied Mathematics, 61 (2):633–658, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/32779.

Nolan:2000:SPF

[122] Clifford J. Nolan. Scattering in the presence of fold caustics. SIAM Journal on Applied Mathematics, 61(2):659– 672, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35610.

Shayer:2000:SBM

[123] Leslie P. Shayer and Sue Ann Campbell. Stability, bifurcation, and multistability in a system of two coupled neurons with multiple time delays. SIAM Journal on Applied Mathematics, 61(2):673– 700, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/34401.

Parkins:2000:SGI

[124] C. J. Parkins. Shock-generated ignition. Newtonian asymptotics for the induction domain between a contact surface and shock. SIAM Journal on Applied Mathematics, 61(2):701–729, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/33156.

Berger:2000:GEL

[125] Kurt M. Berger and Paul A. Milewski. The generation and evolution of lump solitary waves in surface-tension-dominated flows. SIAM Journal on Applied Mathematics, 61(3):731–750, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35697.

Hillen:2000:DLT

[126] Thomas Hillen and Hans G. Othmer. The diffusion limit of transport equations derived from velocityjump processes. SIAM Journal on Applied Mathematics, 61(3):751–775, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35816.

Elston:2000:PFC

[127] Timothy C. Elston, Dwight You, and Charles S. Peskin. Protein flexibility and the correlation ratchet. SIAM Journal on Applied Mathematics, 61(3):776– 791, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35394.

Gardner:2000:EMR

[128] Carl L. Gardner, Joseph W. Jerome, and Robert S. Eisenberg. Electrodiffusion model of rectangular current pulses in ionic channels of cellular membranes. SIAM Journal on Applied Mathematics, 61(3):792–802, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35164.

Feng:2000:EMAa

[129] Zhilan Feng and Horst R. Thieme.

Endemic models with arbitrarily distributed periods of infection I: Fundamental properties of the model. SIAM Journal on Applied Mathematics, 61 (3):803–833, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/34783.

Yin:2000:SPD

[130] G. Yin and Q. Zhang. Singularly perturbed discrete-time Markov chains. SIAM Journal on Applied Mathematics, 61(3):834–854, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/35429.

Xiao:2000:CMV

[131] MingQing Xiao and Tamer Basar. Center manifold of the viscous Moore– Greitzer PDE model. SIAM Journal on Applied Mathematics, 61(3):855– 869, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35426.

Yakubu:2000:SPP

[132] Abdul-Aziz Yakubu. Searching predator and prey dominance in discrete predator-prey systems with dispersion. SIAM Journal on Applied Mathematics, 61(3):870–888, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/34642.

Naire:2000:LCG

[133] S. Naire, R. J. Braun, and S. A. Snow. Limiting cases of gravitational drainage of a vertical free film for evaluating surfactants. SIAM Journal on Applied Mathematics, 61(3):889– 913, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/35676.

Knessl:2000:AAB

[134] Charles Knessl. Asymptotic analysis of A backward-forward parabolic problem for data-handling systems. SIAM Journal on Applied Mathematics, 61(3):914– 933, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/35241.

Keenan:2000:SBM

[135] Daniel M. Keenan, Weimin Sun, and Johannes D. Veldhuis. A stochastic biomathematical model of the male reproductive hormone system. SIAM Journal on Applied Mathematics, 61 (3):934–965, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/33471.

Pieroux:2000:HBS

[136] Didier Pieroux, Thomas Erneux, Athanasios Gavrielides, and Vassilios Kovanis. Hopf bifurcation subject to a large delay in a laser system. SIAM Journal on Applied Mathematics, 61 (3):966–982, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/36013.

Feng:2000:EMAb

[137] Zhilan Feng and Horst R. Thieme.

Endemic models with arbitrarily distributed periods of infection II: Fast disease dynamics and permanent recovery. SIAM Journal on Applied Mathematics, 61(3):983–1012, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/34784.

Kay:2000:SNS

[138] Alison L. Kay and Jonathan A. Sherratt. Spatial noise stabilizes periodic wave patterns in oscillatory systems on finite domains. SIAM Journal on Applied Mathematics, 61(3):1013– 1041, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36069.

Li:2000:GSZ

[139] Tong Li. Global solutions and zero relaxation limit for a traffic flow model. SIAM Journal on Applied Mathematics, 61(3): 1042–1061, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/35678.

Bernis:2000:LLT

[140] Francisco Bernis, Josephus Hulshof, and Fernando Quirós. The “linear” limit of thin film flows as an obstacle-type free boundary problem. SIAM Journal on Applied Mathematics, 61(3):1062– 1079, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/35827.

Doelman:2000:SMT

[141] Arjen Doelman, Wiktor Eckhaus, and Tasso J. Kaper. Slowly modulated two-pulse solutions in the Gray–Scott model I: Asymptotic construction and stability. SIAM Journal on Applied Mathematics, 61(3):1080–1102, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/35492.

Bayliss:2000:RRB

[142] A. Bayliss, T.-K. Ma, B. J. Matkowsky, and C. W. Wahle. The reactive Rayleigh–Bénard problem with throughflow. SIAM Journal on Applied Mathematics, 61(4):1103–1142, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/34785.

Senn:2000:SNI

[143] Walter Senn and Robert Urbanczik. Similar nonleaky integrate-and-fire neurons with instantaneous couplings always synchronize. SIAM Journal on Applied Mathematics, 61(4):1143–1155, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/34603.

Velazquez:2000:CFU

[144] J. J. L. Velázquez. Corner formation for the undercooled Stefan problem. SIAM Journal on Applied Mathematics, 61(4): 1156–1201, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/33979.

Ponomarev:2000:SVC

[145] Igor Ponomarev. Separation of variables in the computation of spectra in 2-D photonic crystals. SIAM Journal on Applied Mathematics, 61(4):1202– 1218, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/32089.

Ni:2000:SPS

[146] Luqun Ni and Sia Nemat-Nasser.

Singular perturbation solutions of a class of systems of singular integral equations. SIAM Journal on Applied Mathematics, 61(4):1219–1236, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36574.

Magnanini:2000:WPD

[147] Rolando Magnanini and Fadil Santosa. Wave propagation in a 2-D optical waveguide. SIAM Journal on Applied Mathematics, 61(4):1237–1252, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36048.

Galaktionov:2000:QOD

[148] Victor A. Galaktionov, Stéphane Gerbi, and Juan L. Vazquez. Quenching for a one-dimensional fully nonlinear parabolic equation in detonation theory. SIAM Journal on Applied Mathematics, 61(4):1253–1285, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/34192.

Richardson:2000:CPT

[149] G. Richardson and J. Rubinstein. Classification of phase transitions in thin structures with small Ginzburg–Landau parameter. SIAM Journal on Applied Mathematics, 61(4):1286–1307, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36184.

Kuske:2000:RCS

[150] Rachel Kuske and Joseph B. Keller. Rate of convergence to a stable law. SIAM Journal on Applied Mathematics, 61(4):1308–1323, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/34271.

Cheng:2000:MIS

[151] Hongwei Cheng. On the method of images for systems of closely spaced conducting spheres. SIAM Journal on Applied Mathematics, 61(4):1324– 1337, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36499.

Chan:2000:VRN

[152] Tony Chan and Jianhong Shen. Variational restoration of nonflat image features: Models and algorithms. SIAM Journal on Applied Mathematics, 61(4): 1338–1361, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/35799.

Herron:2000:PES

[153] Isom H. Herron. On the principle of exchange of stabilities in Rayleigh– Bénard convection. SIAM Journal on Applied Mathematics, 61(4):1362–1368, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/37038.

Tuwankotta:2000:SRH

[154] J. M. Tuwankotta and F. Verhulst.

Symmetry and resonance in Hamiltonian systems. SIAM Journal on Applied Mathematics, 61(4):1369–1385, 2000. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36532.

Haddar:2001:EBC

[155] H. Haddar and P. Joly. Effective boundary conditions for thin ferromagnetic layers: The one-dimensional model. SIAM Journal on Applied Mathematics, 61(4):1386–1417, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/34636.

Norbury:2001:LSI

[156] John Norbury and Li-Chin Yeh. The location and stability of interface solutions of an inhomogeneous parabolic problem. SIAM Journal on Applied Mathematics, 61(4):1418–1430, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/32649.

Steinberg:2001:FPS

[157] Y. Steinberg, B. Z. Bobrovsky, and Z. Schuss. Fixed-point smoothing of scalar diffusions II: The error of the optimal smoother. SIAM Journal on Applied Mathematics, 61(4):1431–1444, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/32122.

Ruan:2001:GAP

[158] Shigui Ruan and Dongmei Xiao. Global analysis in a predator-prey system with nonmonotonic functional response. SIAM Journal on Applied Mathematics, 61(4):1445–1472, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36189.

Mikula:2001:EPC

[159] Karol Mikula and Daniel Sevcovic. Evolution of plane curves driven by a nonlinear function of curvature and anisotropy. SIAM Journal on Applied Mathematics, 61(5):1473–1501, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35928.

Anderson:2001:MSM

[160] D. M. Anderson, M. G. Forest, and R. Superfine. A model for a spreading and melting droplet on a heated substrate. SIAM Journal on Applied Mathematics, 61(5):1502–1525, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36718.

Beckers:2001:SPG

[161] J. M. Beckers, E. Delhez, and E. Deleersnijder. Some properties of generalized age-distribution equations in fluid dynamics. SIAM Journal on Applied Mathematics, 61(5):1526–1544, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36381.

Fannjiang:2001:RTS

[162] Albert Fannjiang and Leonid Ryzhik.

Radiative transfer of sound waves in a random flow: Turbulent scattering, straining, and mode-coupling. SIAM Journal on Applied Mathematics, 61(5): 1545–1577, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/33839.

Terman:2001:PAP

[163] D. H. Terman, G. B. Ermentrout, and A. C. Yew. Propagating activity patterns in thalamic neuronal networks. SIAM Journal on Applied Mathematics, 61(5):1578–1604, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36509.

Kim:2001:VSM

[164] Sun-Chul Kim and Stephen Childress. Vorticity selection with multiple eddies in two-dimensional steady flow at high Reynolds number. SIAM Journal on Applied Mathematics, 61(5):1605– 1617, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/34981.

Simon:2001:RDE

[165] Peter L. Simon. A retarded differential equation model of wave propagation in a thin ring. SIAM Journal on Applied Mathematics, 61(5):1618– 1627, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/31555.

Cheney:2001:OAM

[166] Margaret Cheney, David Isaacson, and Matti Lassas. Optimal acoustic measurements. SIAM Journal on Applied Mathematics, 61(5):1628–1647, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/36911.

Elmer:2001:TWS

[167] Christopher E. Elmer and Erik S. Van Vleck. Traveling wave solutions for bistable differential-difference equations with periodic diffusion. SIAM Journal on Applied Mathematics, 61(5):1648– 1679, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/35711.

Fibich:2001:SFD

[168] G. Fibich. Self-focusing in the damped nonlinear Schrödinger equation. SIAM Journal on Applied Mathematics, 61(5): 1680–1705, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/36260.

Poulton:2001:NLE

[169] C. G. Poulton, L. C. Botten, R. C.

McPhedran, N. A. Nicorovici, and A. B. Movchan. Noncommuting limits in electromagnetic scattering: Asymptotic analysis for an array of highly conducting inclusions. SIAM Journal on Applied Mathematics, 61(5):1706–1730, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35226.

Owen:2001:MLT

[170] Markus R. Owen and Mark A. Lewis.

The mechanics of lung tissue under highfrequency ventilation. SIAM Journal on Applied Mathematics, 61(5):1731– 1761, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36365.

Medvedev:2001:STD

[171] Georgi S. Medvedev and Nancy Kopell.

Synchronization and transient dynamics in the chains of electrically coupled Fitzhugh–Nagumo oscillators. SIAM Journal on Applied Mathematics, 61(5): 1762–1801, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/36880.

Georgieva:2001:RMS

[172] Anna Georgieva, Thomas Kriecherbauer, and Stephanos Venakides. 1: 2 resonance mediated second harmonic generation in a 1-D nonlinear discrete periodic medium. SIAM Journal on Applied Mathematics, 61(5):1802–1815, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/36534.

Smith:2001:AAB

[173] Gregory D. Smith, Longxiang Dai, Robert M. Miura, and Arthur Sherman. Asymptotic analysis of buffered calcium diffusion near a point source. SIAM Journal on Applied Mathematics, 61(5): 1816–1838, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/36899.

Craster:2001:FPC

[174] R. V. Craster and Yu. V. Obnosov. Fourphase checkerboard composites. SIAM Journal on Applied Mathematics, 61(6): 1839–1856, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/37182.

Tartakovsky:2001:DFS

[175] Daniel M. Tartakovsky and C. L. Winter. Dynamics of free surfaces in random porous media. SIAM Journal on Applied Mathematics, 61(6):1857– 1876, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/35818.

Antoine:2001:MDS

[176] Xavier Antoine. Microlocal diagonalization of strictly hyperbolic pseudodifferential systems and application to the design of radiation conditions in electromagnetism. SIAM Journal on Applied Mathematics, 61(6):1877–1905, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35382.

Harlen:2001:LFP

[177] O. G. Harlen, M. J. Holmes, M. J. W.

Povey, Y. Qiu, and B. D. Sleeman. A low frequency potential scattering description of acoustic propagation in dispersions. SIAM Journal on Applied Mathematics, 61(6):1906– 1931, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35984.

Cercignani:2001:DCB

[178] C. Cercignani, I. M. Gamba, and C. D. Levermore. A drift-collision balance for a Boltzmann–Poisson system in bounded domains. SIAM Journal on Applied Mathematics, 61(6):1932– 1958, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36046.

Figotin:2001:SPT

[179] Alex Figotin and Yuri A. Godin.

Spectral properties of thin-film photonic crystals. SIAM Journal on Applied Mathematics, 61(6):1959–1979, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/37283.

Carasso:2001:DBD

[180] Alfred S. Carasso. Direct blind deconvolution. SIAM Journal on Applied Mathematics, 61(6):1980–2007, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/36259.

Figotin:2001:SPN

[181] Alex Figotin and Ilya Vitebsky. Spectra of periodic nonreciprocal electric circuits. SIAM Journal on Applied Mathematics, 61(6):2008–2035, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/37058.

Doelman:2001:SMTb

[182] Arjen Doelman, Wiktor Eckhaus, and Tasso J. Kaper. Slowly modulated twopulse solutions in the Gray–Scott model II: Geometric theory, bifurcations, and splitting dynamics. SIAM Journal on Applied Mathematics, 61(6):2036–2062, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/37242.

Kwasniok:2001:LDM

[183] Frank Kwasniok. Low-dimensional models of the Ginzburg–Landau equation. SIAM Journal on Applied Mathematics, 61(6):2063–2079, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/36821.

Wall:2001:EPV

[184] O. Wall and J. Holst. Estimation of parameters in viscoplastic and creep material models. SIAM Journal on Applied Mathematics, 61(6):2080–2103, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36722.

DeLillo:2001:DSA

[185] Thomas DeLillo, Victor Isakov, Nicolas Valdivia, and Lianju Wang. The detection of the source of acoustical noise in two dimensions. SIAM Journal on Applied Mathematics, 61(6):2104– 2121, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36715.

Smith:2001:MME

[186] Warren R. Smith, John R. King, and Brian Tuck. Mathematical modelling of electrical-optical effects in semiconductor laser operation. SIAM Journal on Applied Mathematics, 61(6):2122– 2147, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36692.

Schreiber:2001:UMR

[187] Sebastian J. Schreiber. Urn models, replicator processes, and random genetic drift. SIAM Journal on Applied Mathematics, 61(6):2148–2167, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35285.

Carpio:2001:PSS

[188] A. Carpio, S. J. Chapman, and J. L. L. Velázquez. Pile-up solutions for some systems of conservation laws modelling dislocation interaction in crystals. SIAM Journal on Applied Mathematics, 61(6): 2168–2199, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/36434.

Stevens:2001:EDC

[189] Angela Stevens. Erratum: “The Derivation of Chemotaxis Equations as Limit Dynamics of Moderately Interacting Stochastic Many-Particle Systems”. SIAM Journal on Applied Mathematics, 61(6):2200, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/37749. See

[104].

Gremaud:2001:PCP

[190] Pierre A. Gremaud, C. T. Kelley, T. A. Royal, and Kristy A. Coffey. On a powder consolidation problem. SIAM Journal on Applied Mathematics, 62(1):1– 20, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36847.

Babich:2001:RDA

[191] Vassily M. Babich, Aleksei P. Kiselev, James M. H. Lawry, and Alexander S. Starkov. A ray description of all wavefields generated by a high-frequency point source near an interface. SIAM Journal on Applied Mathematics, 62(1): 21–40, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/33871.

Galper:2001:MSB

[192] A. R. Galper and T. Miloh. Modified slender body theory for a structure of variable cross-section. SIAM Journal on Applied Mathematics, 62(1):41– 57, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/37066.

Li:2001:GDS

[193] Michael Y. Li, Hal L. Smith, and Liancheng Wang. Global dynamics of an SEIR epidemic model with vertical transmission. SIAM Journal on Applied Mathematics, 62(1):58– 69, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35986.

Batcho:2001:EON

[194] Paul F. Batcho. The energy operator and new scaling relations for the incompressible Navier–Stokes equations. SIAM Journal on Applied Mathematics, 62(1):70–93, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/35412.

Alves:2001:BIF

[195] Carlos Alves and Habib Ammari.

Boundary integral formulae for the reconstruction of imperfections of small diameter in an elastic medium. SIAM Journal on Applied Mathematics, 62(1): 94–106, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36926.

Warne:2001:AFD

[196] Debra Polignone Warne and Paul G. Warne. An asymptotic finite deformation analysis for an isotropic incompressible hyperelastic half-space subjected to a tensile point load. SIAM Journal on Applied Mathematics, 62(1):107– 128, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/37699.

Heinze:2001:VPP

[197] S. Heinze, G. Papanicolaou, and A. Stevens. Variational principles for propagation speeds in inhomogeneous media. SIAM Journal on Applied Mathematics, 62(1):129–148, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36114.

Bowen:2001:AED

[198] M. Bowen, J. Hulshof, and J. R. King. Anomalous exponents and dipole solutions for the thin film equation. SIAM Journal on Applied Mathematics, 62 (1):149–179, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/36693.

Promislow:2001:ARC

[199] Keith Promislow and John M. Stockie. Adiabatic relaxation of convectivediffusive gas transport in a porous fuel cell electrode. SIAM Journal on Applied Mathematics, 62(1):180– 205, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36248.

Pinto:2001:SSAa

[200] David J. Pinto and G. Bard Ermentrout.

Spatially structured activity in synaptically coupled neuronal networks: I. traveling fronts and pulses. SIAM Journal on Applied Mathematics, 62(1):206– 225, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/34645.

Pinto:2001:SSAb

[201] David J. Pinto and G. Bard Ermentrout. Spatially structured activity in synaptically coupled neuronal networks: II. lateral inhibition and standing pulses. SIAM Journal on Applied Mathematics, 62(1):226– 243, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/34646.

Hariharan:2001:CAS

[202] S. I. Hariharan and G. W. Young.

Comparison of asymptotic solutions of a phase-field model to a sharpinterface model. SIAM Journal on Applied Mathematics, 62(1):244–263, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/37490.

Rotstein:2001:PFE

[203] Horacio G. Rotstein, Simon Brandon, Amy Novick-Cohen, and Alexander Nepomnyashchy. Phase field equations with memory: The hyperbolic case. SIAM Journal on Applied Mathematics, 62(1):264–282, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36910.

Wearing:2001:NAJ

[204] Helen J. Wearing and Jonathan A. Sherratt. Nonlinear analysis of juxtacrine patterns. SIAM Journal on Applied Mathematics, 62(1):283–309, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/37220.

Knessl:2001:TDC

[205] Charles Knessl. On two-dimensional convection-diffusion past a circle. SIAM Journal on Applied Mathematics, 62 (1):310–335, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/37515.

Li:2001:HMS

[206] Bingtuan Li and Hal Smith. How many species can two essential resources support? SIAM Journal on Applied Mathematics, 62(1):336–366, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35319.

Dinov:2001:AFD

[207] Ivo D. Dinov and De Witt L. Sumners. Applications of frequency dependent wavelet shrinkage to analyzing quality of image registration. SIAM Journal on Applied Mathematics, 62(2):367– 384, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35854.

Nie:2001:SHS

[208] Qing Nie and Fei-Ran Tian. Singularities in Hele–Shaw flows driven by a multipole. SIAM Journal on Applied Mathematics, 62(2):385–406, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/37419.

Schweizer:2001:BAS

[209] Ben Schweizer. Bifurcation analysis for surface waves generated by wind. SIAM Journal on Applied Mathematics, 62 (2):407–423, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/37256.

Caginalp:2001:RGC

[210] G. Caginalp. Renormalization group calculation of late stage interface dynamics. SIAM Journal on Applied Mathematics, 62(2):424–432, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/38327.

Nadler:2001:SAP

[211] B. Nadler, T. Naeh, and Z. Schuss. The stationary arrival process of independent diffusers from a continuum to an absorbing boundary is poissonian. SIAM Journal on Applied Mathematics, 62(2):433– 447, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/37236.

Nolan:2001:PRM

[212] Clifford J. Nolan. Permittivity recovery from multiple characteristic electric waves. SIAM Journal on Applied Mathematics, 62(2):448–461, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/38403.

Kuznetsov:2001:BHB

[213] Yu. A. Kuznetsov, O. De Feo, and S. Rinaldi. Belyakov homoclinic bifurcations in a tritrophic food chain model. SIAM Journal on Applied Mathematics, 62 (2):462–487, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/37854.

Haberman:2001:SPT

[214] Richard Haberman. Slow passage through the nonhyperbolic homoclinic orbit associated with a subcritical pitchfork bifurcation for Hamiltonian systems and the change in action. SIAM Journal on Applied Mathematics, 62(2):488– 513, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/37383.

Daripa:2001:GCS

[215] Prabir Daripa and D. Palaniappan. Generalized circle and sphere theorems for inviscid and viscous flows. SIAM Journal on Applied Mathematics, 62(2):514– 540, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/37459.

Grandits:2001:FHU

[216] Peter Grandits. Frequent hedging under transaction costs and a nonlinear Fokker–Planck PDE. SIAM Journal on Applied Mathematics, 62(2):541– 562, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36720.

Loncaric:2001:ASC

[217] J. Loncaric, V. S. Ryaben’kii, and S. V. Tsynkov. Active shielding and control of noise. SIAM Journal on Applied Mathematics, 62(2):563– 596, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36758.

Bourgeat:2001:MMN

[218] Alain Bourgeat, Olivier Gipouloux, and Eduard Marusic-Paloka. Mathematical modelling and numerical simulation of a non-Newtonian viscous flow through a thin filter. SIAM Journal on Applied Mathematics, 62(2):597– 626, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35474.

Margolis:2001:SPA

[219] Stephen B. Margolis and Melvin R.

Baer. A singular-perturbation analysis of the burning-rate eigenvalue for a two-temperature model of deflagrations in confined porous energetic materials. SIAM Journal on Applied Mathematics, 62(2):627–663, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/37778.

Gryazin:2001:NSS

[220] Yuriy A. Gryazin, Michael V. Klibanov, and Thomas R. Lucas. Numerical solution of a subsurface imaging inverse problem. SIAM Journal on Applied Mathematics, 62(2):664– 683, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/37736.

Clemons:2001:AET

[221] C. B. Clemons, S. I. Hariharan, and D. D. Quinn. Amplitude equations for time-dependent solutions of the McKendrick equations. SIAM Journal on Applied Mathematics, 62(2):684– 705, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/37813.

Bose:2001:BOA

[222] Amitabha Bose, Yair Manor, and Farzan Nadim. Bistable oscillations arising from synaptic depression. SIAM Journal on Applied Mathematics, 62(2):706– 727, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/37805.

Greenberg:2001:EAT

[223] J. M. Greenberg. Extensions and amplifications of a traffic model of Aw and Rascle. SIAM Journal on Applied Mathematics, 62(3):729–745, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/37865.

Goodman:2001:MLE

[224] R. H. Goodman, A. J. Majda, and D. W.

McLaughlin. Modulations in the leading edges of midlatitude storm tracks. SIAM Journal on Applied Mathematics, 62 (3):746–776, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/38297.

Piterbarg:2001:TLE

[225] Leonid I. Piterbarg. The top Lyapunov exponent for a stochastic flow modeling the upper ocean turbulence. SIAM Journal on Applied Mathematics, 62(3):777– 800, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36640.

Budd:2001:AMB

[226] C. J. Budd. Asymptotics of multibump blow-up self-similar solutions of the nonlinear Schrödinger equation. SIAM Journal on Applied Mathematics, 62 (3):801–830, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/38239.

Li:2001:TDG

[227] Jiequan Li. On the two-dimensional gas expansion for compressible Euler equations. SIAM Journal on Applied Mathematics, 62(3):831–852, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36134.

Dockery:2001:FFB

[228] J. Dockery and I. Klapper. Finger formation in biofilm layers. SIAM Journal on Applied Mathematics, 62(3):853– 869, 2001. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/37170.

Henry:2002:ETI

[229] B. I. Henry and S. L. Wearne. Existence of Turing instabilities in a two-species fractional reaction-diffusion system. SIAM Journal on Applied Mathematics, 62(3):870–887, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/37522.

Pelesko:2002:MME

[230] John A. Pelesko. Mathematical modeling of electrostatic MEMS with tailored dielectric properties. SIAM Journal on Applied Mathematics, 62(3):888– 908, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/38107.

Fannjiang:2002:DNL

[231] Albert Fannjiang and Tomasz Komorowski. Diffusive and nondiffusive limits of transport in nonmixing flows. SIAM Journal on Applied Mathematics, 62(3):909–923, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/37943.

Vanneste:2002:NDA

[232] J. Vanneste. Nonlinear dynamics of anisotropic disturbances in plane Couette flow. SIAM Journal on Applied Mathematics, 62(3):924–944, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/38142.

Crowdy:2002:CGD

[233] Darren Crowdy. On a class of geometrydriven free boundary problems. SIAM Journal on Applied Mathematics, 62 (3):945–964, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/35798.

Cox:2002:NSR

[234] E. A. Cox, M. P. Mortell, and S. Reck. Nonlinear standing and resonantly forced oscillations in a tube with slowly changing length. SIAM Journal on Applied Mathematics, 62(3):965– 989, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/38182.

Smith:2002:ABM

[235] N. P. Smith, A. J. Pullan, and P. J. Hunter. An anatomically based model of transient coronary blood flow in the heart. SIAM Journal on Applied Mathematics, 62(3):990–1018, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35519.

Chan:2002:MML

[236] Tony F. Chan and Jianhong Shen. Mathematical models for local nontexture inpaintings. SIAM Journal on Applied Mathematics, 62(3):1019–1043, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36884.

vanderHofstad:2002:LDC

[237] Remco van der Hofstad, Gerard Hooghiemstra, and Marten J. Klok. Large deviations for code division multiple access systems. SIAM Journal on Applied Mathematics, 62(3):1044– 1065, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36372.

Fannjiang:2002:SGE

[238] Albert C. Fannjiang, Youn sha Chan, and Glaucio H. Paulino. Strain gradient elasticity for antiplane shear cracks: a hypersingular integrodifferential equation approach. SIAM Journal on Applied Mathematics, 62(3):1066–1091, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/38048.

Bonnetier:2002:CEC

[239] E. Bonnetier and A. Chambolle. Computing the equilibrium configuration of epitaxially strained crystalline films. SIAM Journal on Applied Mathematics, 62(4):1093–1121, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36857.

Antipov:2002:DPW

[240] Y. A. Antipov. Diffraction of a plane wave by a circular cone with an impedance boundary condition. SIAM Journal on Applied Mathematics, 62(4): 1122–1152, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/36332.

Knowles:2002:IIT

[241] James K. Knowles. Impact-induced tensile waves in a rubberlike material. SIAM Journal on Applied Mathematics, 62(4):1153–1175, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/38823.

Nikolopoulos:2002:MMP

[242] C. Nikolopoulos. A model for melting of a pure material during modulated temperature differential scanning calorimetry. SIAM Journal on Applied Mathematics, 62(4):1176–1196, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/38195.

Osan:2002:RTW

[243] Remus Osan, Jonathan Rubin, and Bard Ermentrout. Regular traveling waves in a one-dimensional network of theta neurons. SIAM Journal on Applied Mathematics, 62(4):1197– 1221, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/38725.

Othmer:2002:DLT

[244] Hans G. Othmer and Thomas Hillen. The diffusion limit of transport equations II: Chemotaxis equations. SIAM Journal on Applied Mathematics, 62(4): 1222–1250, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/38277.

Banks:2002:ROM

[245] H. T. Banks, S. C. Beeler, G. M. Kepler, and H. T. Tran. Reduced order modeling and control of thin film growth in an HPCVD reactor. SIAM Journal on Applied Mathematics, 62(4):1251– 1280, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/38328.

Forger:2002:RMM

[246] Daniel B. Forger and Richard E. Kronauer. Reconciling mathematical models of biological clocks by averaging on approximate manifolds. SIAM Journal on Applied Mathematics, 62(4):1281– 1296, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/37358.

Ward:2002:DPS

[247] Michael J. Ward, Darragh McInerney, Paul Houston, David Gavaghan, and Philip Maini. The dynamics and pinning of a spike for a reactiondiffusion system. SIAM Journal on Applied Mathematics, 62(4):1297–1328, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/37511.

Benedetto:2002:PWT

[248] John J. Benedetto and Götz E. Pfander. Periodic wavelet transforms and periodicity detection. SIAM Journal on Applied Mathematics, 62(4):1329– 1368, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/37963.

Ammari:2002:ISP

[249] Habib Ammari, Gang Bao, and John L.

Fleming. An inverse source problem for Maxwell’s equations in magnetoencephalography. SIAM Journal on Applied Mathematics, 62(4):1369– 1382, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/37392.

Kim:2002:LST

[250] Sungwhan Kim, Ohin Kwon, and Jin Keun Seo. Location search techniques for a grounded conductor. SIAM Journal on Applied Mathematics, 62(4): 1383–1393, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/37090.

Russell:2002:ENB

[251] David L. Russell and Luther W. White.

An elementary nonlinear beam theory with finite buckling deformation properties. SIAM Journal on Applied Mathematics, 62(4):1394–1413, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/30913.

Deserable:2002:VTD

[252] Dominique Désérable. A versatile twodimensional cellular automata network for granular flow. SIAM Journal on Applied Mathematics, 62(4):1414–1436, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/35520.

Fibich:2002:SFF

[253] Gadi Fibich, Boaz Ilan, and George Papanicolaou. Self-focusing with fourthorder dispersion. SIAM Journal on Applied Mathematics, 62(4):1437–1462, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/38724.

Muratov:2002:SSS

[254] C. B. Muratov and V. V. Osipov. Stability of the static spike autosolitons in the Gray–Scott model. SIAM Journal on Applied Mathematics, 62(5):1463– 1487, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/38428.

Collet:2002:BDS

[255] Jean-François Collet, Thierry Goudon, Frédéric Poupaud, and Alexis Vasseur. The Beker–Döring system and its Lifshitz–Slyozov limit. SIAM Journal on Applied Mathematics, 62(5):1488– 1500, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/37885.

Degond:2002:VFM

[256] P. Degond, M. Lemou, and M. Picasso.

Viscoelastic fluid models derived from kinetic equations for polymers. SIAM Journal on Applied Mathematics, 62(5): 1501–1519, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/37440.

Chadwick:2002:AIT

[257] R. S. Chadwick. Axisymmetric indentation of a thin incompressible elastic layer. SIAM Journal on Applied Mathematics, 62(5):1520–1530, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/38822.

vanDuijn:2002:EET

[258] C. J. van Duijn, A. Mikelic, and I. S. Pop. Effective equations for two-phase flow with trapping on the micro scale. SIAM Journal on Applied Mathematics, 62(5):1531–1568, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/38556.

OBrien:2002:ASP

[259] Stephen B. G. O’Brien. Asymptotics of a sequence of pendant drops. SIAM Journal on Applied Mathematics, 62(5): 1569–1580, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/38478.

Velazquez:2002:SSM

[260] J. J. L. Velázquez. Stability of some mechanisms of chemotactic aggregation. SIAM Journal on Applied Mathematics, 62(5):1581–1633, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/38004.

Feng:2002:TST

[261] Z. Feng, M. Iannelli, and F. A. Milner. A two-strain tuberculosis model with age of infection. SIAM Journal on Applied Mathematics, 62(5):1634– 1656, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/38205.

Cook:2002:RUP

[262] L. Pamela Cook, Emeka Nwankwo, Gilberto Schleiniger, and Bryan Wood. Rapid, unambiguous polymer characterization by flow-referenced capillary viscometry. SIAM Journal on Applied Mathematics, 62(5):1657–1676, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/39007.

Bal:2002:TTD

[263] Guillaume Bal. Transport through diffusive and nondiffusive regions, embedded objects, and clear layers. SIAM Journal on Applied Mathematics, 62(5):1677– 1697, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/37260.

Schuss:2002:EDS

[264] Z. Schuss and A. Spivak. The exit distribution on the stochastic separatrix in Kramers’ exit problem. SIAM Journal on Applied Mathematics, 62(5):1698– 1711, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35905.

Horgan:2002:TPB

[265] Cornelius O. Horgan, Giuseppe Saccomandi, and Ivonne Sgura. A two-point boundary-value problem for the axial shear of hardening isotropic incompressible nonlinearly elastic materials. SIAM Journal on Applied Mathematics, 62(5): 1712–1727, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/39196.

Jones:2002:BWA

[266] Don Jones, Hristo V. Kojouharov, Dung Le, and Hal Smith. Bacterial wall attachment in a flow reactor. SIAM Journal on Applied Mathematics, 62(5): 1728–1771, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/39041.

Preusser:2002:LSM

[267] Tobias Preusser and Martin Rumpf.

A level set method for anisotropic geometric diffusion in 3D image processing. SIAM Journal on Applied Mathematics, 62(5):1772–1793, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/38466.

Korman:2002:BDH

[268] Can E. Korman, Richard Barry, and Ulas C. Kozat. On binary detection with hysteresis. SIAM Journal on Applied Mathematics, 62(5):1794– 1809, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/37516.

Varas:2002:LSP

[269] Fernando Varas and José M. Vega. Linear stability of a plane front in solid combustion at large heat of reaction. SIAM Journal on Applied Mathematics, 62(5): 1810–1822, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/38641.

Goodman:2002:EEB

[270] Jonathan Goodman and Daniel N. Ostrov. On the early exercise boundary of the American put option. SIAM Journal on Applied Mathematics, 62(5):1823– 1835, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/37829.

Krechetnikov:2002:HIP

[271] Rouslan Krechetnikov and Igor Lipatov. Hidden invariances in problems of two-dimensional and threedimensional wall jets for Newtonian and non-Newtonian fluids. SIAM Journal on Applied Mathematics, 62(6):1837– 1855, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/37890.

Fu:2002:WMR

[272] Y. B. Fu and M. Sanjarani Pour. WKB method with repeated roots and its application to the buckling analysis of an everted cylindrical tube. SIAM Journal on Applied Mathematics, 62(6):1856– 1871, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/38964.

Chapman:2002:EAC

[273] S. J. Chapman and J.-M. Vanden-Broeck. Exponential asymptotics and capillary waves. SIAM Journal on Applied Mathematics, 62(6):1872–1898, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/38116.

Smith:2002:MSS

[274] Warren R. Smith. Models for solidification and splashing in laser percussion drilling. SIAM Journal on Applied Mathematics, 62(6):1899–1923, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/37234.

Iron:2002:DMS

[275] David Iron and Michael J. Ward.

The dynamics of multispike solutions to the one-dimensional Gierer– Meinhardt model. SIAM Journal on Applied Mathematics, 62(6):1924–1951, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/39367.

Clemons:2002:ASP

[276] C. B. Clemons, S. I. Hariharan, and G. W. Young. Asymptotic solutions of a phase-field model for alloy solidification. SIAM Journal on Applied Mathematics, 62(6):1952–1979, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/39153.

Figotin:2002:ONT

[277] A. Figotin, A. Gordon, S. Molchanov, J. Quinn, and N. Stavrakas. Occupancy numbers in testing random number generators. SIAM Journal on Applied Mathematics, 62(6):1980– 2011, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/36686.

Katsevich:2002:TEF

[278] Alexander Katsevich. Theoretically exact filtered backprojection-type inversion algorithm for spiral CT. SIAM Journal on Applied Mathematics, 62(6): 2012–2026, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/38718.

Borrelli:2002:SVP

[279] A. Borrelli, C. O. Horgan, and M. C. Patria. Saint-Venant’s principle for antiplane shear deformations of linear piezoelectric materials. SIAM Journal on Applied Mathematics, 62(6):2027– 2044, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/39250.

Munch:2002:TMD

[280] Andreas Münch. The thickness of a marangoni-driven thin liquid film emerging from a meniscus. SIAM Journal on Applied Mathematics, 62(6):2045– 2063, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/38269.

Milton:2002:CMM

[281] Graeme W. Milton and Knut Sølna. Can mixing materials make electromagnetic signals travel faster? SIAM Journal on Applied Mathematics, 62(6):2064–2091 (electronic), 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/38508.

Dierkes:2002:FDS

[282] Thomas Dierkes, Oliver Dorn, Frank Natterer, Victor Palamodov, and Helmut Sielschott. Fréchet derivatives for some bilinear inverse problems. SIAM Journal on Applied Mathematics, 62(6): 2092–2113, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/38637.

Rossi:2002:ETD

[283] Louis F. Rossi and James Graham-Eagle. On the existence of twodimensional, localized, rotating, selfsimilar vortical structures. SIAM Journal on Applied Mathematics, 62(6): 2114–2128, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/38657.

Haider:2002:BIC

[284] Mansoor A. Haider, Stephen P.

Shipman, and Stephanos Venakides. Boundary-integral calculations of twodimensional electromagnetic scattering in infinite photonic crystal slabs: Channel defects and resonances. SIAM Journal on Applied Mathematics, 62(6): 2129–2148, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/38531.

Fang:2002:PIS

[285] Weifu Fang, Kazufumi Ito, and David A. Redfern. Parameter identification for semiconductor diodes by LBIC imaging. SIAM Journal on Applied Mathematics, 62(6):2149–2174, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/39249.

daMota:2002:CFP

[286] J. da Mota, W. Dantas, and D. Marchesin. Combustion fronts in porous media. SIAM Journal on Applied Mathematics, 62(6):2175–2198, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/34781.

Indeikina:2002:ETE

[287] Alexandra Indeikina and Hsueh-Chia Chang. Estimate of turbulent eddy diffusion by exact renormalization. SIAM Journal on Applied Mathematics, 63(1): 1–41, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/37959.

Tesdall:2002:SSS

[288] Allen M. Tesdall and John K. Hunter. Self-similar solutions for weak shock reflection. SIAM Journal on Applied Mathematics, 63(1):42–61, 2002.

CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/38382.

Laing:2002:MBN

[289] Carlo R. Laing, William C. Troy, Boris Gutkin, and G. Bard Ermentrout. Multiple bumps in a neuronal model of working memory. SIAM Journal on Applied Mathematics, 63(1):62– 97, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/38949.

Edwards:2002:DOP

[290] David A. Edwards and Richard A.

Cairncross. Desorption overshoot in polymer-penetrant systems: Asymptotic and computational results. SIAM Journal on Applied Mathematics, 63(1): 98–115, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/39042.

Piterbarg:2002:SPA

[291] Leonid I. Piterbarg and Tamay M.

Özgökmen. A simple prediction algorithm for the Lagrangian motion in two-dimensional turbulent flows. SIAM Journal on Applied Mathematics, 63 (1):116–148, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/39194.

Argall:2002:RTF

[292] Brenna Argall, Eugene Cheleshkin, J. M. Greenberg, Colin Hinde, and Pei-Jen Lin. A rigorous treatment of a follow-the-leader traffic model with traffic lights present. SIAM Journal on Applied Mathematics, 63(1):149– 168, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/39121.

Warne:2002:AFD

[293] Debra Polignone Warne and Paul G. Warne. An asymptotic finite deformation analysis for an isotropic compressible hyperelastic half-space subjected to a tensile point load. SIAM Journal on Applied Mathematics, 63(1):169– 194, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/39495.

Zelig:2002:HCP

[294] Daphne Zelig and Shimon Haber. Hydrodynamic cleansing of pulmonary alveoli. SIAM Journal on Applied Mathematics, 63(1):195–221, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/38609.

Morrison:2002:ASI

[295] John A. Morrison and K. G. Ramakrishnan. Asymptotic solution to an inverse problem for a shared unbuffered resource. SIAM Journal on Applied Mathematics, 63(1):222–240, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/38879.

Xiang:2002:DCM

[296] Yang Xiang. Derivation of a continuum model for epitaxial growth with elasticity on vicinal surface. SIAM Journal on Applied Mathematics, 63(1):241– 258, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/39828.

Aw:2002:DCT

[297] A. Aw, A. Klar, T. Materne, and M. Rascle. Derivation of continuum traffic flow models from microscopic follow-the-leader models. SIAM Journal on Applied Mathematics, 63(1):259– 278, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/38095.

Degond:2002:DAS

[298] Pierre Degond and Kaijun Zhang.

Diffusion approximation of a scattering matrix model of a semiconductor superlattice. SIAM Journal on Applied Mathematics, 63(1):279–298, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/36001.

Fowler:2002:ROC

[299] A. C. Fowler and Michael C. Mackey.

Relaxation oscillations in a class of delay differential equations. SIAM Journal on Applied Mathematics, 63(1):299– 323, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/39351.

Tomlinson:2002:FEM

[300] Karl A. Tomlinson, Peter J. Hunter, and Andrew J. Pullan. A finite element method for an eikonal equation model of myocardial excitation wavefront propagation. SIAM Journal on Applied Mathematics, 63(1):324– 350, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/38951.

Dumont:2002:ITE

[301] T. Dumont, S. Génieys, M. Massot, and V. A. Volpert. Interaction of thermal explosion and natural convection: Critical conditions and new oscillating regimes. SIAM Journal on Applied Mathematics, 63(1):351– 372, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/38950.

Mudavanhu:2002:NRM

[302] B. Mudavanhu and R. E. O’Malley, Jr. A new renormalization method for the asymptotic solution of weakly nonlinear vector systems. SIAM Journal on Applied Mathematics, 63(2):373– 397, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/39431.

Knessl:2002:DPQ

[303] Charles Knessl, Doo Il Choi, and Charles Tier. A dynamic priority queue model for simultaneous service of two traffic types. SIAM Journal on Applied Mathematics, 63(2):398– 422, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/39084.

Mei:2002:MDT

[304] Z. Mei, A. J. Roberts, and Zhenquan Li. Modelling the dynamics of turbulent floods. SIAM Journal on Applied Mathematics, 63(2):423–458, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35886.

Tonnelier:2002:MCF

[305] Arnaud Tonnelier. The McKean’s caricature of the Fitzhugh–Nagumo model I. the space-clamped system. SIAM Journal on Applied Mathematics, 63(2):459– 484, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/39350.

Yang:2002:WPS

[306] Jianbo Yang, Serafim Kalliadasis, John H. Merkin, and Stephen K. Scott. Wave propagation in spatially distributed excitable media. SIAM Journal on Applied Mathematics, 63(2):485– 509, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/39140.

Ruderman:2002:TME

[307] Gregory A. Ruderman, D. Scott Stewart, and Jack Jai ick Yoh. A thermomechanical model for energetic materials with phase transformations. SIAM Journal on Applied Mathematics, 63 (2):510–537, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/39025.

Yoh:2002:TME

[308] Jack Jai ick Yoh, D. Scott Stewart, and Gregory A. Ruderman. A thermomechanical model for energetic materials with phase transformations: Analysis of simple motions. SIAM Journal on Applied Mathematics, 63(2):538– 563, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/39026.

Chan:2002:EEC

[309] Tony F. Chan, Sung Ha Kang, and Jianhong Shen. Euler’s elastica and curvature-based inpainting. SIAM Journal on Applied Mathematics, 63(2):564– 592, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/39008.

Carasso:2002:AMI

[310] Alfred S. Carasso. The APEX method in image sharpening and the use of low exponent Lévy stable laws. SIAM Journal on Applied Mathematics, 63(2):593– 618, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/38931.

Carpio:2002:PPD

[311] A. Carpio and L. L. Bonilla. Pulse propagation in discrete systems of coupled excitable cells. SIAM Journal on Applied Mathematics, 63(2):619–635, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/39173.

Zhu:2002:BAP

[312] Huaiping Zhu, Sue Ann Campbell, and Gail S. K. Wolkowicz. Bifurcation analysis of a predator-prey system with nonmonotonic functional response. SIAM Journal on Applied Mathematics, 63 (2):636–682, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/39728.

Calvo:2002:DNP

[313] N. Calvo, J. I. Díaz, J. Durany, E. Schiavi, and C. Vázquez. On a doubly nonlinear parabolic obstacle problem modelling ice sheet dynamics. SIAM Journal on Applied Mathematics, 63(2):683– 707, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/38534.

Colombo:2002:HPT

[314] Rinaldo M. Colombo. Hyperbolic phase transitions in traffic flow. SIAM Journal on Applied Mathematics, 63(2):708– 721, 2002. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/39318.

Buckingham:2003:TFT

[315] Robert Buckingham, Michael Shearer, and Andrea Bertozzi. Thin film traveling waves and the Navier slip condition. SIAM Journal on Applied Mathematics, 63(2):722–744, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/40140.

Pelinovsky:2003:ADM

[316] Dmitry E. Pelinovsky and Vadim Zharnitsky. Averaging of dispersion-managed solitons: Existence and stability. SIAM Journal on Applied Mathematics, 63 (3):745–776, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/40047.

deHoop:2003:UAE

[317] Maarten V. de Hoop and A. K. Gautesen. Uniform asymptotic expansion of the square-root Helmholtz operator and the one-way wave propagator. SIAM Journal on Applied Mathematics, 63 (3):777–800, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/40230.

Liefvendahl:2003:ANI

[318] Mattias Liefvendahl and Gunilla Kreiss. Analytical and numerical investigation of the resolvent for plane Couette flow. SIAM Journal on Applied Mathematics, 63(3):801–817, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/39675.

Greenberg:2003:CMH

[319] J. M. Greenberg, A. Klar, and M. Rascle. Congestion on multilane highways. SIAM Journal on Applied Mathematics, 63(3):818–833, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/39630.

Harlen:2003:APD

[320] O. G. Harlen, M. J. Holmes, M. J. W.

Povey, and B. D. Sleeman. Acoustic propagation in dispersions and the geometric theory of diffraction. SIAM Journal on Applied Mathematics, 63(3):834– 849, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/40467.

Nadler:2003:CDI

[321] B. Nadler, T. Naeh, and Z. Schuss.

Connecting a discrete ionic simulation to a continuum. SIAM Journal on Applied Mathematics, 63(3):850– 873, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/39368.

Nguyen:2003:DKS

[322] Hiêú D. Nguyêñ. Decay of KdV solitons. SIAM Journal on Applied Mathematics, 63(3):874–888, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/40269.

Braza:2003:BSH

[323] Peter A. Braza. The bifurcation structure of the Holling–Tanner model for predator-prey interactions using two-timing. SIAM Journal on Applied Mathematics, 63(3):889–904, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/39349.

Nachbin:2003:TFB

[324] André Nachbin. A terrain-following Boussinesq system. SIAM Journal on Applied Mathematics, 63(3):905–922, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/39758.

Alvarez:2003:CRS

[325] Luis H. R. Alvarez. On the convexity and risk-sensitivity of the price of American interest rate derivatives. SIAM Journal on Applied Mathematics, 63 (3):923–936, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/38467.

Sanchez:2003:CMC

[326] Antonio L. Sánchez, Jose L. Jimenez-Alvarez, and Amable Linan. The coupling of motion and conductive heating of a gas by localized energy sources. SIAM Journal on Applied Mathematics, 63(3):937–961, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/40389.

Dorfman:2003:GTA

[327] K. D. Dorfman and H. Brenner. Generalized Taylor–Aris dispersion in spatially periodic microfluidic networks. chemical reactions. SIAM Journal on Applied Mathematics, 63(3):962–986, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/40187.

Movchan:2003:RAS

[328] A. B. Movchan, G. W. Milton, and S. K. Serkov. Realizable (average stress, average strain) pairs in a plate with holes. SIAM Journal on Applied Mathematics, 63(3):987–1028, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/39571.

Lewis:2003:DHB

[329] Gregory M. Lewis and Wayne Nagata.

Double Hopf bifurcations in the differentially heated rotating annulus. SIAM Journal on Applied Mathematics, 63(3): 1029–1055, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/38640.

Carpio:2003:DTD

[330] A. Carpio and L. L. Bonilla. Depinning transitions in discrete reactiondiffusion equations. SIAM Journal on Applied Mathematics, 63(3):1056– 1082, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/39006.

Margolis:2003:AMN

[331] Stephen B. Margolis and Timothy J.

Gardner. An asymptotic model of nonadiabatic catalytic flames in stagnationpoint flow. SIAM Journal on Applied Mathematics, 63(3):1083–1103, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/40245.

Cai:2003:RST

[332] Hongxue Cai and Richard Chadwick. Radial structure of traveling waves in the inner ear. SIAM Journal on Applied Mathematics, 63(4):1105–1120, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/38895.

Berger:2003:SWI

[333] Kurt M. Berger and Paul A. Milewski. Simulation of wave interactions and turbulence in one-dimensional water waves. SIAM Journal on Applied Mathematics, 63(4):1121–1140, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/40206.

Loncaric:2003:OAS

[334] J. Loncaric and S. V. Tsynkov. Optimization of acoustic source strength in the problems of active noise control. SIAM Journal on Applied Mathematics, 63(4):1141–1183, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/40422.

Fowler:2003:ITT

[335] A. C. Fowler and P. D. Howell. Intermittency in the transition to turbulence. SIAM Journal on Applied Mathematics, 63(4):1184–1207, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/36889.

Nykamp:2003:WNA

[336] Duane Q. Nykamp. White noise analysis of coupled linear-nonlinear systems. SIAM Journal on Applied Mathematics, 63(4):1208–1230, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/39757.

Balakotaiah:2003:HHM

[337] Vemuri Balakotaiah and Hsueh-Chia Chang. Hyperbolic homogenized models for thermal and solutal dispersion. SIAM Journal on Applied Mathematics, 63(4):1231–1258, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/36886.

Gubernov:2003:EFS

[338] V. Gubernov, G. N. Mercer, H. S. Sidhu, and R. O. Weber. Evans function stability of combustion waves. SIAM Journal on Applied Mathematics, 63(4):1259– 1275, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/40024.

Kang:2003:APT

[339] Hyeonbae Kang, EunJoo Kim, and Kyoungsun Kim. Anisotropic polarization tensors and detection of an anisotropic inclusion. SIAM Journal on Applied Mathematics, 63(4):1276– 1291, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/40619.

Luke:2003:NRT

[340] D. Russell Luke and Roland Potthast. The no response test—A sampling method for inverse scattering problems. SIAM Journal on Applied Mathematics, 63(4):1292–1312, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/40688.

DeLeenheer:2003:VDG

[341] Patrick De Leenheer and Hal L. Smith. Virus dynamics: a global analysis. SIAM Journal on Applied Mathematics, 63(4):1313–1327, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/40690.

Fannjiang:2003:HFB

[342] Albert Fannjiang, Shi Jin, and George Papanicolaou. High frequency behavior of the focusing nonlinear Schrödinger equation with random inhomogeneities. SIAM Journal on Applied Mathematics, 63(4):1328–1358, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/35559.

Uecker:2003:AIB

[343] Hannes Uecker. Approximation of the integral boundary layer equation by the Kuramoto–Sivashinsky equation. SIAM Journal on Applied Mathematics, 63(4): 1359–1377, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/40590.

Dercole:2003:BAP

[344] Fabio Dercole, Jean-Olivier Irisson, and Sergio Rinaldi. Bifurcation analysis of a prey-predator coevolution model. SIAM Journal on Applied Mathematics, 63(4): 1378–1391, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/41161.

Vynnycky:2003:AMM

[345] M. Vynnycky and E. Birgersson. Analysis of a model for multicomponent mass transfer in the cathode of a polymer electrolyte fuel cell. SIAM Journal on Applied Mathematics, 63(4):1392– 1423, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/39369.

Wang:2003:LSI

[346] Dehua Wang. Large solutions to the initial-boundary value problem for planar magnetohydrodynamics. SIAM Journal on Applied Mathematics, 63(4): 1424–1441, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/40928.

Daniele:2003:WHT

[347] V. G. Daniele. The Wiener–Hopf technique for impenetrable wedges having arbitrary aperture angle. SIAM Journal on Applied Mathematics, 63(4):1442– 1460, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/40023.

Sivaloganathan:2003:MRC

[348] Jeyabal Sivaloganathan and Scott J. Spector. Myriad radial cavitating equilibria in nonlinear elasticity. SIAM Journal on Applied Mathematics, 63(4): 1461–1473, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/39700.

Bal:2003:TRR

[349] Guillaume Bal and Leonid Ryzhik. Time reversal and refocusing in Random media. SIAM Journal on Applied Mathematics, 63(5):1475–1498, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/40108.

Huang:2003:HIS

[350] Huaxiong Huang, Robert M. Miura, William P. Ireland, and Ernest Puil. Heat-induced stretching of a glass tube under tension: Application to glass microelectrodes. SIAM Journal on Applied Mathematics, 63(5):1499–1519, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/39346.

Sherratt:2003:PTW

[351] Jonathan A. Sherratt. Periodic travelling wave selection by Dirichlet boundary conditions in oscillatory reactiondiffusion systems. SIAM Journal on Applied Mathematics, 63(5):1520–1538, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/39248.

LeVeque:2003:SWL

[352] Randall J. LeVeque and Darryl H.

Yong. Solitary waves in layered nonlinear media. SIAM Journal on Applied Mathematics, 63(5):1539–1560, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/40815.

Sopasakis:2003:FAM

[353] Alexandros Sopasakis. Formal asymptotic models of vehicular traffic. model closures. SIAM Journal on Applied Mathematics, 63(5):1561–1584, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/40302.

Noonburg:2003:PFW

[354] V. W. Noonburg, D. Benardete, and B. Pollina. A periodically forced Wilson–Cowan system. SIAM Journal on Applied Mathematics, 63(5):1585– 1603, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/40814.

Blyth:2003:OFN

[355] M. G. Blyth and P. Hall. Oscillatory flow near a stagnation point. SIAM Journal on Applied Mathematics, 63(5):1604– 1614, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/40817.

Nishikawa:2003:SPO

[356] Takashi Nishikawa and Frank C. Hoppensteadt. System of phase oscillators with diagonalizable interaction. SIAM Journal on Applied Mathematics, 63(5): 1615–1626, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/41113.

Drover:2003:NCN

[357] Jonathan D. Drover and Bard Ermentrout. Nonlinear coupling near a degenerate Hopf (Bautin) bifurcation. SIAM Journal on Applied Mathematics, 63(5): 1627–1647, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/41261.

Fouque:2003:SPO

[358] J.-P. Fouque, G. Papanicolaou, R. Sircar, and K. Sølna. Singular perturbations in option pricing. SIAM Journal on Applied Mathematics, 63(5):1648– 1665, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/40155.

Landman:2003:CCM

[359] K. A. Landman, G. J. Pettet, and D. F. Newgreen. Chemotactic cellular migration: Smooth and discontinuous travelling wave solutions. SIAM Journal on Applied Mathematics, 63(5):1666– 1681, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/40469.

vandenBerg:2003:FAB

[360] Jan Bouwe van den Berg, Josephus Hulshof, and John R. King. Formal asymptotics of bubbling in the harmonic map heat flow. SIAM Journal on Applied Mathematics, 63(5):1682– 1717, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/40887.

Fibich:2003:BNA

[361] G. Fibich, B. Ilan, and S. Tsynkov.

Backscattering and nonparaxiality arrest collapse of damped nonlinear waves. SIAM Journal on Applied Mathematics, 63(5):1718–1736, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/41185.

Slemrod:2003:RFD

[362] M. Slemrod. The radio-frequency driven plasma sheath: Asymptotics and analysis. SIAM Journal on Applied Mathematics, 63(5):1737–1763, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/41183.

Solna:2003:APS

[363] Knut Sølna. Acoustic pulse spreading in a Random fractal. SIAM Journal on Applied Mathematics, 63(5):1764– 1788, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/40465.

Frohlich:2003:OPW

[364] J. Fröhlich and M. Uhlmann. Orthonormal polynomial wavelets on the interval and applications to the analysis of turbulent flow fields. SIAM Journal on Applied Mathematics, 63(5):1789– 1830, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/40411.

Ritter:2003:IFR

[365] L. R. Ritter, W. E. Olmstead, and V. A. Volpert. Initiation of free-radical polymerization waves. SIAM Journal on Applied Mathematics, 63(5):1831– 1848, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/41324.

Wang:2003:EES

[366] Nianqing Wang and Peter Smereka. Effective equations for sound and void wave propagation in bubbly fluids. SIAM Journal on Applied Mathematics, 63(6):1849–1888, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/41305.

Liu:2003:CTR

[367] Hailiang Liu and Eitan Tadmor. Critical thresholds in 2D restricted Euler– Poisson equations. SIAM Journal on Applied Mathematics, 63(6):1889–1910, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/41698.

Frigaard:2003:EST

[368] I. A. Frigaard, G. Ngwa, and O. Scherzer. On effective stopping time selection for visco-plastic nonlinear BV diffusion filters used in image denoising. SIAM Journal on Applied Mathematics, 63(6):1911–1934, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/40046.

Izhikevich:2003:SCO

[369] Eugene M. Izhikevich and Frank C.

Hoppensteadt. Slowly coupled oscillators: Phase dynamics and synchronization. SIAM Journal on Applied Mathematics, 63(6):1935–1953, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/40094.

Fister:2003:OCA

[370] K. Renee Fister and John Carl Panetta.

Optimal control applied to competing chemotherapeutic cell-kill strategies. SIAM Journal on Applied Mathematics, 63(6):1954–1971, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/41348.

Ragazzo:2003:MST

[371] Clodoaldo Grotta Ragazzo. On the motion of solids through an ideal liquid: Approximated equations for many body systems. SIAM Journal on Applied Mathematics, 63(6):1972–1997, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/39427.

Rotstein:2003:CML

[372] Horacio G. Rotstein, Nancy Kopell, Anatol M. Zhabotinsky, and Irving R. Epstein. A Canard mechanism for localization in systems of globally coupled oscillators. SIAM Journal on Applied Mathematics, 63(6):1998–2019, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/41184.

Ali:2003:LMN

[373] G. Alì. Low Mach number flows in time-dependent domains. SIAM Journal on Applied Mathematics, 63(6):2020– 2041, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/40073.

Fannjiang:2003:FPA

[374] Albert Fannjiang and Tomasz Komorowski. Frozen path approximation for turbulent diffusion and fractional Brownian motion in random flows. SIAM Journal on Applied Mathematics, 63(6):2042–2062, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/33529.

Al-Omari:2003:STF

[375] J. F. M. Al-Omari and S. A. Gourley. Stability and traveling fronts in Lotka–Volterra competition models with stage structure. SIAM Journal on Applied Mathematics, 63(6):2063– 2086, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/41650.

Chapman:2003:HMS

[376] S. J. Chapman and D. R. Heron. A hierarchy of models for superconducting thin films. SIAM Journal on Applied Mathematics, 63(6):2087–2127, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/41033.

Aubert:2003:ISU

[377] Gilles Aubert, Michel Barlaud, Olivier Faugeras, and Stéphanie Jehan-Besson. Image segmentation using active contours: Calculus of variations or shape gradients? SIAM Journal on Applied Mathematics, 63(6):2128–2154, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/40892.

Morrison:2003:ASEb

[378] John A. Morrison. Asymptotic shape of the Erlang capacity region of a critically loaded multiservice shared resource. SIAM Journal on Applied Mathematics, 64(1):1–17, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/40109.

Kriegsmann:2003:SMD

[379] James J. Kriegsmann and Michael J.

Miksis. Steady motion of a drop along a liquid interface. SIAM Journal on Applied Mathematics, 64(1):18– 40, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/40021.

Berres:2003:SDP

[380] Stefan Berres, Raimund Bürger, Kenneth H. Karlsen, and Elmer M. Tory. Strongly degenerate parabolichyperbolic systems modeling polydisperse sedimentation with compression.

SIAM Journal on Applied Mathematics, 64(1):41–80, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/40816.

Pilyugin:2003:DCG

[381] Sergei S. Pilyugin and Paul Waltman. Divergence criterion for generic planar systems. SIAM Journal on Applied Mathematics, 64(1):81–93, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/41841.

Leise:2003:MSD

[382] Tanya L. Leise and Jay R. Walton. A method for solving dynamically accelerating crack problems in linear viscoelasticity. SIAM Journal on Applied Mathematics, 64(1):94–107, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/41953.

Burger:2003:FOD

[383] Martin Burger and René Pinnau. Fast optimal design of semiconductor devices. SIAM Journal on Applied Mathematics, 64(1):108–126, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/42056.

Morrison:2003:ASE

[384] John A. Morrison and Debasis Mitra. Asymptotic shape of the Erlang capacity region of a multiservice shared resource. SIAM Journal on Applied Mathematics, 64(1):127– 151, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/39941.

Andrews:2003:MAC

[385] Kevin T. Andrews, L. Chapman, J. R.

Fernández, M. Fisackerly, M. Shillor, L. Vanerian, and T. VanHouten. A membrane in adhesive contact. SIAM Journal on Applied Mathematics, 64 (1):152–169, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/40620.

Wellander:2003:HME

[386] Niklas Wellander and Gerhard Kristensson. Homogenization of the Maxwell equations at fixed frequency. SIAM Journal on Applied Mathematics, 64 (1):170–195, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/40336.

Gobbert:2003:HTB

[387] Matthias K. Gobbert and Christian Ringhofer. A homogenization technique for the Boltzmann equation for low pressure chemical vapor deposition. SIAM Journal on Applied Mathematics, 64 (1):196–215, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/39347.

Boatto:2003:NSL

[388] Stefanella Boatto and Hildeberto E.

Cabral. Nonlinear stability of a latitudinal ring of point-vortices on a nonrotating sphere. SIAM Journal on Applied Mathematics, 64(1):216–230, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/39996.

Edelman:2003:GPD

[389] P. H. Edelman, S. W. Peterson, V. Reiner, and J. H. Stout. Geochemical phase diagrams and Gale diagrams. SIAM Journal on Applied Mathematics, 64(1):231–259, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/41182.

Arino:2003:GRE

[390] Julien Arino, C. Connell McCluskey, and P. van den Driessche. Global results for an epidemic model with vaccination that exhibits backward bifurcation. SIAM Journal on Applied Mathematics, 64(1):260–276, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/41382.

Wang:2003:AVA

[391] Jing Wang, Robert Gulliver, and Fadil Santosa. Analysis of a variational approach to progressive lens design. SIAM Journal on Applied Mathematics, 64 (1):277–296, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/40894.

Martin:2003:ASI

[392] P. A. Martin. Acoustic scattering by inhomogeneous obstacles. SIAM Journal on Applied Mathematics, 64(1):297– 308, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/41437.

Reif:2003:LIS

[393] B. A. Pettersson Reif and O. Andreassen. On local isotropy in stratified homogeneous turbulence. SIAM Journal on Applied Mathematics, 64(1):309– 321, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/42155.

Shipman:2003:RBS

[394] Stephen P. Shipman and Stephanos Venakides. Resonance and bound states in photonic crystal slabs. SIAM Journal on Applied Mathematics, 64(1):322– 342, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/41112.

Chen:2003:HBQ

[395] Zhi-Min Chen, Michael Ghil, Eric Simonnet, and Shouhong Wang. Hopf bifurcation in quasi-geostrophic channel flow. SIAM Journal on Applied Mathematics, 64(1):343–368, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/40616.

Kramer:2003:SMRa

[396] Peter R. Kramer and Andrew J. Majda.

Stochastic mode reduction for the immersed boundary method. SIAM Journal on Applied Mathematics, 64(2):369– 400, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/42213.

Kramer:2003:SMRb

[397] Peter R. Kramer and Andrew J. Majda. Stochastic mode reduction for particlebased simulation methods for complex microfluid systems. SIAM Journal on Applied Mathematics, 64(2):401– 422, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/42214.

Yariv:2003:EMC

[398] Ehud Yariv and Howard Brenner. The electrophoretic mobility of a closely fitting sphere in a cylindrical pore. SIAM Journal on Applied Mathematics, 64 (2):423–441, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/41111.

Hintermuller:2003:SOS

[399] Michael Hintermüller and Wolfgang Ring. A second order shape optimization approach for image segmentation. SIAM Journal on Applied Mathematics, 64 (2):442–467, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/40390.

Gunther:2003:MFD

[400] Marco Günther, Axel Klar, Thorsten Materne, and Raimund Wegener. Multivalued fundamental diagrams and stop and go waves for continuum traffic flow equations. SIAM Journal on Applied Mathematics, 64(2):468–483, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/40470.

Hagerty:2003:RII

[401] Patrick Hagerty, Anthony M. Bloch, and Michael I. Weinstein. Radiation induced instability. SIAM Journal on Applied Mathematics, 64(2):484– 524, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/41871.

Elschner:2003:GPR

[402] J. Elschner, G. C. Hsiao, and A. Rathsfeld. Grating profile reconstruction based on finite elements and optimization techniques. SIAM Journal on Applied Mathematics, 64(2):525–545, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/42001.

Flyer:2003:NIB

[403] Natasha Flyer and Bengt Fornberg. On the nature of initial-boundary value solutions for dispersive equations. SIAM Journal on Applied Mathematics, 64 (2):546–564, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/41585.

Larsen:2003:NFA

[404] Edward W. Larsen, Guido Thömmes, and Axel Klar. New frequency-averaged approximations to the equations of radiative heat transfer. SIAM Journal on Applied Mathematics, 64(2):565– 582, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/41384.

Gremaud:2003:SCG

[405] Pierre A. Gremaud, John V. Matthews, and David G. Schaeffer. Secondary circulation in granular flow through nonaxisymmetric hoppers. SIAM Journal on Applied Mathematics, 64(2):583– 600, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/41512.

Ammari:2003:RMT

[406] Habib Ammari and Faouzi Triki. Resonances for microstrip transmission lines. SIAM Journal on Applied Mathematics, 64(2):601–636, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/41839.

Ruan:2003:HSE

[407] Weihua Ruan, M. E. Clark, Meide Zhao, and Anthony Curcio. A hyperbolic system of equations of blood flow in an arterial network. SIAM Journal on Applied Mathematics, 64(2):637– 667, 2003. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/41529.

Droske:2004:VAN

[408] M. Droske and M. Rumpf. A variational approach to nonrigid morphological image registration. SIAM Journal on Applied Mathematics, 64(2):668– 687, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/41952.

Van:2004:ATE

[409] Tri Van and Aihua Wood. Analysis of transient electromagnetic scattering from overfilled cavities. SIAM Journal on Applied Mathematics, 64(2):688– 708, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/41925.

Cakoni:2004:DSI

[410] Fioralba Cakoni and David Colton. The determination of the surface impedance of a partially coated obstacle from far field data. SIAM Journal on Applied Mathematics, 64(2):709–723, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/42425.

Zhao:2004:ARM

[411] Hongkai Zhao. Analysis of the response matrix for an extended target. SIAM Journal on Applied Mathematics, 64 (3):725–745, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/41528.

Abramov:2004:ABN

[412] Vyacheslav M. Abramov. Asymptotic behavior of the number of lost messages. SIAM Journal on Applied Mathematics, 64(3):746–761, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/40525.

Dobson:2004:OLE

[413] David C. Dobson and Fadil Santosa. Optimal localization of eigenfunctions in an inhomogeneous medium. SIAM Journal on Applied Mathematics, 64(3):762– 774, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/42616.

Garcke:2004:DIM

[414] Harald Garcke, Britta Nestler, and Bjorn Stinner. A diffuse interface model for alloys with multiple components and phases. SIAM Journal on Applied Mathematics, 64(3):775– 799, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/41314.

Barashenkov:2004:TSD

[415] I. V. Barashenkov and E. V. Zemlyanaya. Traveling solitons in the damped-driven nonlinear Schrödinger equation. SIAM Journal on Applied Mathematics, 64 (3):800–818, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/42483.

Hong:2004:BTV

[416] John Hong and Blake Temple. A bound on the total variation of the conserved quantities for solutions of a general resonant nonlinear balance law. SIAM Journal on Applied Mathematics, 64(3):819– 857, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/40524.

Kugler:2004:PIP

[417] Philipp Kügler. A parameter identification problem of mixed type related to the manufacture of car windshields. SIAM Journal on Applied Mathematics, 64(3):858–877, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/42333.

Andrianov:2004:SRP

[418] Nikolai Andrianov and Gerald Warnecke. On the solution to the Riemann problem for the compressible duct flow. SIAM Journal on Applied Mathematics, 64(3):878–901, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/42423.

Hyvonen:2004:CEM

[419] Nuutti Hyvönen. Complete electrode model of electrical impedance tomography: Approximation properties and characterization of inclusions. SIAM Journal on Applied Mathematics, 64 (3):902–931, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/42330.

Keenan:2004:SDM

[420] Daniel M. Keenan and Paula A. Shorter.

Simulation by diffusion on a manifold with boundary: Applications to ultrasonic prenatal medical imaging. SIAM Journal on Applied Mathematics, 64 (3):932–960, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/40890.

Haber:2004:ESS

[421] Shimon Haber. Exact similarity solutions of coupled nonsteady Navier– Stokes and energy equations in liquids. SIAM Journal on Applied Mathematics, 64(3):961–976, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/42045.

Grajales:2004:DWA

[422] Juan Carlos Muñoz Grajales and André Nachbin. Dispersive wave attenuation due to orographic forcing. SIAM Journal on Applied Mathematics, 64(3):977– 1001, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/41276.

Berlyand:2004:HNN

[423] L. Berlyand and E. Khruslov. Homogenized non-Newtonian viscoelastic rheology of a suspension of interacting particles in a viscous Newtonian fluid. SIAM Journal on Applied Mathematics, 64(3): 1002–1034, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/40391.

Linton:2004:SIA

[424] C. M. Linton and P. A. Martin. Semiinfinite arrays of isotropic point scatterers. A unified approach. SIAM Journal on Applied Mathematics, 64(3):1035– 1056, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/42789.

Hazard:2004:SAF

[425] Christophe Hazard and Karim Ramdani. Selective acoustic focusing using time-harmonic reversal mirrors. SIAM Journal on Applied Mathematics, 64(3): 1057–1076, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/42873.

Hamdache:2004:IEC

[426] Kamel Hamdache and Mouhcine Tilioua.

Interlayer exchange coupling for ferromagnets through spacers. SIAM Journal on Applied Mathematics, 64(3):1077– 1097, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/39891.

Yariv:2004:CID

[427] Ehud Yariv, Howard Brenner, and Sangtae Kim. Curvature-induced dispersion in electro-osmotic serpentine flows. SIAM Journal on Applied Mathematics, 64(4):1099–1124, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/42284.

Bass:2004:GOA

[428] F. Bass, V. Freilikher, A. A. Maradudin, and V. Prosentsov. Geometrical optics approximation for nonlinear equations. SIAM Journal on Applied Mathematics, 64(4):1125–1132, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/42661.

Papanicolaou:2004:SST

[429] George Papanicolaou, Leonid Ryzhik, and Knut Sølna. Statistical stability in time reversal. SIAM Journal on Applied Mathematics, 64(4):1133– 1155, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/41110.

Bazin:2004:SMN

[430] Pierre-Louis Bazin and Mireille Boutin. Structure from motion: a new look from the point of view of invariant theory. SIAM Journal on Applied Mathematics, 64(4):1156–1174, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/40246.

Greenberg:2004:CR

[431] J. M. Greenberg. Congestion redux. SIAM Journal on Applied Mathematics, 64(4):1175–1185, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/43173.

Chadwick:2004:SGF

[432] Richard S. Chadwick, Brett Shoelson, and Hongxue Cai. Surface Green’s functions for an incompressible, transversely isotropic elastic half-space. SIAM Journal on Applied Mathematics, 64(4): 1186–1197, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/42533.

Velazquez:2004:PDSa

[433] J. J. L. Velázquez. Point dynamics in a singular limit of the Keller– Segel model 1: Motion of the concentration regions. SIAM Journal on Applied Mathematics, 64(4):1198–1223, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/43388.

Velazquez:2004:PDSb

[434] J. J. L. Velázquez. Point dynamics in a singular limit of the Keller– Segel model 2: Formation of the concentration regions. SIAM Journal on Applied Mathematics, 64(4):1224–1248, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/43389.

Wang:2004:PEH

[435] Guan Jun Wang and Jacques Beaumont. Parameter estimation of the Hodgkin–Huxley gating model: An inversion procedure. SIAM Journal on Applied Mathematics, 64(4):1249–1267, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/41982.

Ilhan:2004:SPB

[436] Aytac Ilhan, Mattias Jonsson, and Ronnie Sircar. Singular perturbations for boundary value problems arising from exotic options. SIAM Journal on Applied Mathematics, 64(4):1268– 1293, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/42004.

Gleeson:2004:MAM

[437] James P. Gleeson, Olivia M. Roche, Jonathan West, and Anne Gelb. Modelling annular micromixers. SIAM Journal on Applied Mathematics, 64(4): 1294–1310, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/42040.

Hintermuller:2004:TBV

[438] M. Hintermüller and K. Kunisch. Total bounded variation regularization as a bilaterally constrained optimization problem. SIAM Journal on Applied Mathematics, 64(4):1311–1333, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/42278.

Crowdy:2004:CMC

[439] Darren Crowdy and Jonathan Marshall. Constructing multiply connected quadrature domains. SIAM Journal on Applied Mathematics, 64(4):1334– 1359, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/43854.

Pelinovsky:2004:PRR

[440] Dmitry E. Pelinovsky and Jianke Yang. Parametric resonance and radiative decay of dispersion-managed solitons. SIAM Journal on Applied Mathematics, 64(4):1360–1382, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/42235.

Kirillov:2004:CKC

[441] Oleg N. Kirillov and Alexander P.

Seyranian. Collapse of the Keldysh chains and stability of continuous nonconservative systems. SIAM Journal on Applied Mathematics, 64(4):1383– 1407, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/41472.

Olyslager:2004:DCS

[442] Frank Olyslager. Discretization of continuous spectra based on perfectly matched layers. SIAM Journal on Applied Mathematics, 64(4):1408–1433, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/43019.

Antoniou:2004:DTT

[443] Eliana S. Antoniou, John K. Bechtold, and Moshe Matalon. A diffusionalthermal theory of near-stoichiometric premixed flames. SIAM Journal on Applied Mathematics, 64(4):1434–1456, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/42428.

Auer:2004:SFB

[444] Ch. Auer, F. Schürrer, and W. Koller.

A semicontinuous formulation of the Bloch–Boltzmann–Peierls equations. SIAM Journal on Applied Mathematics, 64(4):1457–1475, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://

epubs.siam.org/sam-bin/dbq/article/
42610.

Sochen:2004:SCP

[445] Nir A. Sochen, Chen Sagiv, and Ron Kimmel. Stereographic combing a porcupine or studies on direction diffusion in image processing. SIAM Journal on Applied Mathematics, 64(5):1477– 1508, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/41551.

Fellner:2004:BPN

[446] Klemens Fellner and Christian Schmeiser. Burgers–Poisson: a nonlinear dispersive model equation. SIAM Journal on Applied Mathematics, 64(5):1509– 1525, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/41034.

Goudon:2004:LFL

[447] Thierry Goudon, Oscar Sánchez, Juan Soler, and Luis L. Bonilla. Lowfield limit for a nonlinear discrete driftdiffusion model arising in semiconductor superlattices theory. SIAM Journal on Applied Mathematics, 64(5):1526– 1549, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/41730.

Faugeras:2004:WPT

[448] Olivier Faugeras and Gerardo Hermosillo. Well-posedness of two nonrigid multimodal image registration methods. SIAM Journal on Applied Mathematics, 64(5):1550–1587, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/42490.

Haider:2004:RBM

[449] Mansoor A. Haider. A radial biphasic model for local cell-matrix mechanics in articular cartilage. SIAM Journal on Applied Mathematics, 64(5):1588– 1608, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/41770.

Rubin:2004:SSP

[450] Jonathan E. Rubin and William C.

Troy. Sustained spatial patterns of activity in neuronal populations without recurrent excitation. SIAM Journal on Applied Mathematics, 64(5):1609– 1635, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/42580.

Aregba-Driollet:2004:MMS

[451] D. Aregba-Driollet, F. Diele, and R. Natalini. A mathematical model for the sulphur dioxide aggression to calcium carbonate stones: Numerical approximation and asymptotic analysis. SIAM Journal on Applied Mathematics, 64(5): 1636–1667, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/42829.

Bressloff:2004:EST

[452] Paul C. Bressloff. Euclidean shifttwist symmetry in population models of self-aligning objects. SIAM Journal on Applied Mathematics, 64(5):1668– 1690, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/43601.

Shen:2004:BVD

[453] Jianhong Shen. Bayesian video dejittering by the BV image model. SIAM Journal on Applied Mathematics, 64(5): 1691–1708, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/41869.

Abdallah:2004:MPH

[454] N. Ben Abdallah, I. M. Gamba, and Axel Klar. The Milne problem for high field kinetic equations. SIAM Journal on Applied Mathematics, 64(5):1709– 1736, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/40889.

Yong:2004:DRL

[455] Wen-An Yong. Diffusive relaxation limit of multidimensional isentropic hydrodynamical models for semiconductors. SIAM Journal on Applied Mathematics, 64(5):1737–1748, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/42740.

Carasso:2004:SII

[456] Alfred S. Carasso. Singular integrals, image smoothness, and the recovery of texture in image deblurring. SIAM Journal on Applied Mathematics, 64(5): 1749–1774, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/42830.

Budd:2004:SSB

[457] C. J. Budd, V. A. Galaktionov, and J. F. Williams. Self-similar blowup in higher-order semilinear parabolic equations. SIAM Journal on Applied Mathematics, 64(5):1775–1809, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/41552.

Fouque:2004:TRD

[458] Jean-Pierre Fouque, Josselin Garnier, and André Nachbin. Time reversal for dispersive waves in random media. SIAM Journal on Applied Mathematics, 64(5):1810–1838, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/42237.

Wilkening:2004:ASDa

[459] Jon Wilkening, Len Borucki, and J. A.

Sethian. Analysis of stress-driven grain boundary diffusion. Part I. SIAM Journal on Applied Mathematics, 64(6): 1839–1863, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/43823.

Wilkening:2004:ASDb

[460] Jon Wilkening, Len Borucki, and J. A. Sethian. Analysis of stressdriven grain boundary diffusion. Part II: Degeneracy. SIAM Journal on Applied Mathematics, 64(6):1864–1886, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/43824.

Podio-Guidugli:2004:EDW

[461] Paolo Podio-Guidugli and Giuseppe Tomassetti. On the evolution of domain walls in hard ferromagnets. SIAM Journal on Applied Mathematics, 64(6): 1887–1906, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/43402.

Ikehata:2004:NSC

[462] Masaru Ikehata and Samuli Siltanen. Numerical solution of the Cauchy problem for the stationary Schrödinger equation using Faddeev’s Green function. SIAM Journal on Applied Mathematics, 64(6):1907–1932, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/42491.

Ringhofer:2004:SAE

[463] C. Ringhofer, M. Nedjalkov, H. Kosina, and S. Selberherr. Semiclassical approximation of electron-phonon scattering beyond Fermi’s Golden Rule. SIAM Journal on Applied Mathematics, 64(6): 1933–1953, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/42891.

Jonsson:2004:RAW

[464] B. Lars G. Jonsson, Mats Gustafsson, Vaughan H. Weston, and Maarten V. de Hoop. Retrofocusing of acoustic wave fields by iterated time reversal. SIAM Journal on Applied Mathematics, 64(6): 1954–1986, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/42696.

Destuynder:2004:NCP

[465] Philippe Destuynder and Jocelyne Vétillard. A noise control problem arising in a flow duct. SIAM Journal on Applied Mathematics, 64(6):1987– 2017, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/41891.

Ammari:2004:GWP

[466] Habib Ammari and Fadil Santosa.

Guided waves in a photonic bandgap structure with a line defect. SIAM Journal on Applied Mathematics, 64(6): 2018–2033, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/40402.

Guo:2004:CFS

[467] X. Guo and Q. Zhang. Closed-form solutions for perpetual American put options with regime switching. SIAM Journal on Applied Mathematics, 64(6): 2034–2049, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/42608.

Miller:2004:MRI

[468] Gregory H. Miller. Minimal rotationally invariant bases for hyperelasticity. SIAM Journal on Applied Mathematics, 64(6): 2050–2075, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/43877.

Saucier:2004:ASG

[469] Antoine Saucier, Jean-Yves Degorce, and Michel Meunier. Analytical solutions of a growth model for a melt region induced by a focused laser beam. SIAM Journal on Applied Mathematics, 64(6): 2076–2095, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/41301.

Bridges:2004:NCW

[470] Thomas J. Bridges and Fiona E. Laine-Pearson. Nonlinear counterpropagating waves, multisymplectic geometry, and the instability of standing waves. SIAM Journal on Applied Mathematics, 64(6): 2096–2120, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/42375.

Armano:2004:NMD

[471] Tiziana Armano and Paolo Cermelli. Noncrystallographic motion of a dislocation as a fine mixture of rectilinear paths. SIAM Journal on Applied Mathematics, 64(6):2121–2143, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/43063.

Matar:2004:RSC

[472] Omar K. Matar and Satish Kumar. Rupture of a surfactant-covered thin liquid film on a flexible wall. SIAM Journal on Applied Mathematics, 64(6):2144– 2166, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/42002.

Nita:2004:FSS

[473] Bogdan G. Nita, Kenneth H. Matson, and Arthur B. Weglein. Forward scattering series and seismic events:

Far field approximations, critical and postcritical events. SIAM Journal on Applied Mathematics, 64(6):2167–2185, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/43561.

Rubinstein:2004:GLM

[474] Jacob Rubinstein, Michelle Schatzman, and Peter Sternberg. Ginzburg– Landau model in thin loops with narrow constrictions. SIAM Journal on Applied Mathematics, 64(6):2186–2204, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/43445.

Li:2004:EMM

[475] Jia Li, Yican Zhou, Zhien Ma, and James M. Hyman. Epidemiological models for mutating pathogens. SIAM Journal on Applied Mathematics, 65(1):1– 23, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/43018.

Kriegsmann:2004:CTT

[476] G. A. Kriegsmann. Complete transmission through a two-dimensional diffraction grating. SIAM Journal on Applied Mathematics, 65(1):24–42, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/42739.

Efendiev:2004:NHC

[477] Y. Efendiev and A. Pankov. Numerical homogenization and correctors for nonlinear elliptic equations. SIAM Journal on Applied Mathematics, 65(1):43– 68, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/42488.

Drover:2004:ACM

[478] Jonathan Drover, Jonathan Rubin, Jianzhong Su, and Bard Ermentrout. Analysis of a Canard mechanism by which excitatory synaptic coupling can synchronize neurons at low firing frequencies. SIAM Journal on Applied Mathematics, 65(1):69– 92, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/43123.

Komorowski:2004:CED

[479] T. Komorowski and P. Widelski. Computation of the effective diffusivity tensor for transport of a passive scalar in a turbulent incompressible flow. SIAM Journal on Applied Mathematics, 65(1): 93–112, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/43519.

Bila:2004:ASA

[480] Nicoleta Bîla. Application of symmetry analysis to a PDE arising in the car windshield design. SIAM Journal on Applied Mathematics, 65(1):113– 130, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/43403.

Bressloff:2004:FBE

[481] Paul C. Bressloff and Stefanos E. Folias. Front bifurcations in an excitatory neural network. SIAM Journal on Applied Mathematics, 65(1):131– 151, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/43448.

Topaz:2004:SPT

[482] Chad M. Topaz and Andrea L.

Bertozzi. Swarming patterns in a twodimensional kinematic model for biological groups. SIAM Journal on Applied Mathematics, 65(1):152–174, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/43742.

Hald:2004:SSS

[483] Ole H. Hald and Aleksei I. Shestakov. Stability of stationary solutions of the multifrequency radiation diffusion equations. SIAM Journal on Applied Mathematics, 65(1):175–193, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/40730.

Haddar:2004:FDO

[484] Houssem Haddar and Rainer Kress.

On the Fréchet derivative for obstacle scattering with an impedance boundary condition. SIAM Journal on Applied Mathematics, 65(1):194–208, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/43541.

Wu:2004:MMC

[485] Jianhua Wu, Hua Nie, and Gail S. K.

Wolkowicz. A mathematical model of competition for two essential resources in the unstirred chemostat. SIAM Journal on Applied Mathematics, 65(1):209– 229, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/42328.

Glasman:2004:IWP

[486] E. A. Glasman, A. A. Golovin, and A. A. Nepomnyashchy. Instabilities of wavy patterns governed by coupled Burgers equations. SIAM Journal on Applied Mathematics, 65(1):230– 251, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/43291.

Ammari:2004:SEI

[487] Habib Ammari, Ohin Kwon, Jin Keun Seo, and Eung Je Woo. T-scan electrical impedance imaging system for anomaly detection. SIAM Journal on Applied Mathematics, 65(1):252–266, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/43375.

Champneys:2004:RTI

[488] A. R. Champneys and W. B. Fraser. Resonance tongue interaction in the parametrically excited column. SIAM Journal on Applied Mathematics, 65 (1):267–298, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/41827.

Lipton:2004:BEC

[489] Robert Lipton and Tungyang Chen. Bounds and extremal configurations for the torsional rigidity of coated fiber reinforced shafts. SIAM Journal on Applied Mathematics, 65(1):299– 315, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/42422.

Campbell:2004:DCB

[490] Sue Ann Campbell, R. Edwards, and P. van den Driessche. Delayed coupling between two neural network loops. SIAM Journal on Applied Mathematics, 65(1):316–335, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/43483.

Levine:2004:SFC

[491] Howard A. Levine and Joanna Renclawowicz. Singularity formation in chemotaxis—A conjecture of Nagai. SIAM Journal on Applied Mathematics, 65(1):336–360, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/43172. See erratum

[590].

Erban:2004:ICB

[492] Radek Erban and Hans G. Othmer.

From individual to collective behavior in bacterial chemotaxis. SIAM Journal on Applied Mathematics, 65(2):361– 391, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/43323.

Kuznetsov:2004:SPH

[493] Alexey Kuznetsov, Mads Kærn, and Nancy Kopell. Synchrony in a population of hysteresis-based genetic oscillators. SIAM Journal on Applied Mathematics, 65(2):392–425, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/43602.

Corli:2004:RPR

[494] Andrea Corli and Haitao Fan. The Riemann problem for reversible reactive flows with metastability. SIAM Journal on Applied Mathematics, 65(2):426– 457, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/42967.

Lin:2004:SDB

[495] Changhao Lin and Lawrence E. Payne. Spatial decay bounds in time dependent pipe flow of an incompressible viscous fluid. SIAM Journal on Applied Mathematics, 65(2):458–474, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/60632.

Lipton:2004:HTA

[496] Robert Lipton. Homogenization theory and the assessment of extreme field values in composites with random microstructure. SIAM Journal on Applied Mathematics, 65(2):475–493, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/42697.

Cortez:2004:PRI

[497] Ricardo Cortez, Charles S. Peskin, John M. Stockie, and Douglas Varela. Parametric resonance in immersed elastic boundaries. SIAM Journal on Applied Mathematics, 65(2):494–520, 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/42534.

Park:2005:DRM

[498] J. H. Park, A. Bayliss, and B. J.

Matkowsky. Dynamics in a rod model of solid flame waves. SIAM Journal on Applied Mathematics, 65(2):521– 549, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/60378.

Gourley:2005:DRD

[499] Stephen A. Gourley and Yang Kuang.

A delay reaction-diffusion model of the spread of bacteriophage infection. SIAM Journal on Applied Mathematics, 65 (2):550–566, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/43661.

Yin:2005:TTS

[500] G. Yin and Hanqin Zhang. Two-time-scale Markov chains and applications to quasi-birth-death queues. SIAM Journal on Applied Mathematics, 65(2):567– 586, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/39756.

Staab:2005:ARI

[501] P. L. Staab, M. J. Rempe, and D. R. Kassoy. Acoustic-rotational internal flow caused by transient sidewall mass addition. SIAM Journal on Applied Mathematics, 65(2):587–617, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/42156.

Kuwamura:2005:TPO

[502] Masataka Kuwamura. On the Turing patterns in one-dimensional gradient/skew-gradient dissipative systems. SIAM Journal on Applied Mathematics, 65(2):618–643, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/42489.

Atay:2005:SBN

[503] Fatihcan M. Atay and Axel Hutt. Stability and bifurcations in neural fields with finite propagation speed and general connectivity. SIAM Journal on Applied Mathematics, 65(2):644–666, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/43088.

BrazeSilva:2005:REP

[504] Pablo Braz e Silva. Resolvent estimates for plane Couette flow. SIAM Journal on Applied Mathematics, 65(2):667– 683, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/42694.

Basener:2005:BCP

[505] Bill Basener and David S. Ross. Booming and crashing populations and Easter Island. SIAM Journal on Applied Mathematics, 65(2):684–701, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/42695.

Laurie:2005:PGH

[506] H. Laurie and R. Penne. Projective geometry of human motion, with an application to injury risk. SIAM Journal on Applied Mathematics, 65(2):702– 719, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/42965.

Lin:2005:GFK

[507] Chun-Chi Lin and Hartmut R. Schwetlick.

On the geometric flow of Kirchhoff elastic rods. SIAM Journal on Applied Mathematics, 65(2):720–736, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/43171.

Xiao:2005:BRD

[508] Dongmei Xiao and Leslie Stephen Jennings. Bifurcations of a ratio-dependent predator-prey system with constant rate harvesting. SIAM Journal on Applied Mathematics, 65(3):737–753, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/42871.

Liu:2005:GSP

[509] Weishi Liu. Geometric singular perturbation approach to steady-state Poisson–Nernst–Planck systems. SIAM Journal on Applied Mathematics, 65 (3):754–766, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/42093.

Cakoni:2005:DSC

[510] Fioralba Cakoni, David Colton, and Peter Monk. The determination of the surface conductivity of a partially coated dielectric. SIAM Journal on Applied Mathematics, 65(3):767– 789, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/60422.

Sleeman:2005:ESS

[511] B. D. Sleeman, Michael J. Ward, and J. C. Wei. The existence and stability of spike patterns in a chemotaxis model. SIAM Journal on Applied Mathematics, 65(3):790–817, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/41511.

Andersson:2005:FIR

[512] Fredrik Andersson. Fast inversion of the Radon transform using logpolar coordinates and partial backprojections. SIAM Journal on Applied Mathematics, 65(3):818–837, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/43600.

Baginski:2005:DAI

[513] Frank Baginski. On the design and analysis of inflated membranes: Natural and pumpkin shaped balloons. SIAM Journal on Applied Mathematics, 65(3):838– 857, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/43847.

Delfour:2005:MDC

[514] Michel C. Delfour. Modeling and design of coated stents to optimize the effect of the dose. SIAM Journal on Applied Mathematics, 65(3):858– 881, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/41160.

Burger:2005:MCS

[515] Raimund Bürger, Kenneth H. Karlsen, and John D. Towers. A model of continuous sedimentation of flocculated suspensions in clarifier-thickener units. SIAM Journal on Applied Mathematics, 65 (3):882–940, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/60620.

Crowdy:2005:ESE

[516] Darren Crowdy and Michael Siegel.

Exact solutions for the evolution of a bubble in Stokes flow: a Cauchy transform approach. SIAM Journal on Applied Mathematics, 65(3):941–963, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/43084.

Nuno:2005:DTS

[517] M. Nuño, Z. Feng, M. Martcheva, and C. Castillo-Chavez. Dynamics of two-strain influenza with isolation and partial cross-immunity. SIAM Journal on Applied Mathematics, 65(3):964– 982, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/43882.

Hoffmann:2005:HIB

[518] K.-H. Hoffmann, N. D. Botkin, and V. N. Starovoitov. Homogenization of interfaces between rapidly oscillating fine elastic structures and fluids. SIAM Journal on Applied Mathematics, 65(3): 983–1005, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/42157.

Holcman:2005:MCD

[519] D. Holcman and Z. Schuss. Modeling calcium dynamics in dendritic spines. SIAM Journal on Applied Mathematics, 65(3):1006–1026, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/42894.

Izen:2005:ESF

[520] Steven H. Izen, David P. Rohler, and Sastry K. L. A. Exploiting symmetry in fan beam CT: Overcoming third generation undersampling. SIAM Journal on Applied Mathematics, 65(3):1027– 1052, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/41700.

Bonilla:2005:PGP

[521] L. L. Bonilla, M. Kindelan, and J. B.

Keller. Periodically generated propagating pulses. SIAM Journal on Applied Mathematics, 65(3):1053–1079, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/43494.

Kirkinis:2005:GAS

[522] Eleftherios Kirkinis and Hungyu Tsai.

Generalized azimuthal shear deformations in compressible isotropic elastic materials. SIAM Journal on Applied Mathematics, 65(3):1080–1099, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/43807.

Agueev:2005:MWR

[523] Dmitri Agueev and Dmitry Pelinovsky.

Modeling of wave resonances in lowcontrast photonic crystals. SIAM Journal on Applied Mathematics, 65(4): 1101–1129, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/60605.

Gutierrez-Romero:2005:UPN

[524] Susana Gutiérrez-Romero, Jesús F.

Palacián, and Patricia Yanguas. A universal procedure for normalizing ndegree-of-freedom polynomial Hamiltonian systems. SIAM Journal on Applied Mathematics, 65(4):1130–1152, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/43439.

Elmer:2005:SDF

[525] Christopher E. Elmer and Erik S. Van Vleck. Spatially discrete FitzHugh– Nagumo equations. SIAM Journal on Applied Mathematics, 65(4):1153– 1174, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/43687.

Dorfman:2005:CSF

[526] Abram Dorfman. Combustion stabilization by forced oscillations in a duct. SIAM Journal on Applied Mathematics, 65(4):1175–1199, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/41557.

Chung:2005:HFI

[527] Soon-Yeong Chung and Carlos A. Berenstein. ω-harmonic functions and inverse conductivity problems on networks. SIAM Journal on Applied Mathematics, 65(4):1200–1226, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/43274.

Memoli:2005:DFG

[528] Facundo Mémoli and Guillermo Sapiro. Distance functions and geodesics on submanifolds of R^{d} and point clouds. SIAM Journal on Applied Mathematics, 65(4): 1227–1260, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/42877.

Araujo:2005:MTGa

[529] Robyn P. Araujo and D. L. Sean McElwain. A mixture theory for the genesis of residual stresses in growing tissues I: a general formulation. SIAM Journal on Applied Mathematics, 65(4):1261– 1284, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/60711.

Caflisch:2005:MDO

[530] R. E. Caflisch, Mark F. Gyure, Hans D.

Robinson, and Eli Yablonovitch. Modeling, design, and optimization of a solid state electron spin qubit. SIAM Journal on Applied Mathematics, 65(4):1285– 1304, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/60618.

Lutscher:2005:EDP

[531] Frithjof Lutscher, Elizaveta Pachepsky, and Mark A. Lewis. The effect of dispersal patterns on stream populations. SIAM Journal on Applied Mathematics, 65(4):1305–1327, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/44040.

Adimy:2005:MSH

[532] Mostafa Adimy, Fabien Crauste, and Shigui Ruan. A mathematical study of the hematopoiesis process with applications to chronic myelogenous leukemia. SIAM Journal on Applied Mathematics, 65(4):1328–1352, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/60469.

Campos:2005:SUD

[533] L. M. B. C. Campos and P. G. T. A. Serrão. On the sound in unbounded and ducted vortex flows. SIAM Journal on Applied Mathematics, 65(4):1353– 1368, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/42707.

Schuster:2005:RSL

[534] Thomas Schuster and Eric Todd Quinto. On a regularization scheme for linear operators in distribution spaces with an application to the spherical Radon transform. SIAM Journal on Applied Mathematics, 65(4):1369–1387, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/43879.

Stolk:2005:MSD

[535] Christiaan C. Stolk and Maarten V. de Hoop. Modeling of seismic data in the downward continuation approach. SIAM Journal on Applied Mathematics, 65(4): 1388–1406, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/43954.

Chiera:2005:SPC

[536] B. A. Chiera, A. E. Krzesinski, and P. G. Taylor. Some properties of the capacity value function. SIAM Journal on Applied Mathematics, 65(4):1407– 1419, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/43085.

Landman:2005:DCC

[537] K. A. Landman, M. J. Simpson, J. L.

Slater, and D. F. Newgreen. Diffusive and chemotactic cellular migration: Smooth and discontinuous traveling wave solutions. SIAM Journal on Applied Mathematics, 65(4):1420– 1442, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/60406.

White:2005:ATE

[538] Benjamin S. White. Asymptotic theory of electroseismic prospecting. SIAM Journal on Applied Mathematics, 65(4): 1443–1462, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/60410.

Bazant:2005:CVR

[539] Martin Z. Bazant, Kevin T. Chu, and B. J. Bayly. Current-voltage relations for electrochemical thin films. SIAM Journal on Applied Mathematics, 65(5): 1463–1484, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/60993.

Chu:2005:ETF

[540] Kevin T. Chu and Martin Z. Bazant.

Electrochemical thin films at and above the classical limiting current. SIAM Journal on Applied Mathematics, 65(5): 1485–1505, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/60992.

Jaffe:2005:TDP

[541] Jules S. Jaffe. Three-dimensional probability density functions via tomographic inversion. SIAM Journal on Applied Mathematics, 65(5):1506– 1525, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/42390.

Craciun:2005:MEC

[542] Gheorghe Craciun and Martin Feinberg. Multiple equilibria in complex chemical reaction networks: I. The injectivity property. SIAM Journal on Applied Mathematics, 65(5):1526–1546, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/44027.

Diaz:2005:AHO

[543] Julien Diaz and Patrick Joly. An analysis of higher order boundary conditions for the wave equation. SIAM Journal on Applied Mathematics, 65(5):1547– 1575, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/43614.

Ghil:2005:SBD

[544] Michael Ghil, Tian Ma, and Shouhong Wang. Structural bifurcation of 2-D nondivergent flows with Dirichlet boundary conditions: Applications to boundary-layer separation. SIAM Journal on Applied Mathematics, 65(5): 1576–1596, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/43881.

Wang:2005:ASE

[545] Wendi Wang and Xiao-Qiang Zhao.

An age-structured epidemic model in a patchy environment. SIAM Journal on Applied Mathematics, 65(5):1597– 1614, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/43124.

Dascal:2005:MPB

[546] Lorina Dascal and Nir A. Sochen.

A maximum principle for Beltrami color flow. SIAM Journal on Applied Mathematics, 65(5):1615–1632, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/43083.

Hoppe:2005:PSF

[547] R. H. W. Hoppe, W. G. Litvinov, and T. Rahman. Problems of stationary flow of electrorheological fluids in a cylindrical coordinate system. SIAM Journal on Applied Mathematics, 65(5):1633– 1656, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/43288.

Protopopescu:2005:DAL

[548] V. Protopopescu and L. Thevenot. Diffusion approximation for linear transport with multiplying boundary conditions. SIAM Journal on Applied Mathematics, 65(5):1657–1676, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/42424.

King:2005:RKU

[549] J. R. King and J. D. Evans. Regularization by kinetic undercooling of blow-up in the ill-posed Stefan problem. SIAM Journal on Applied Mathematics, 65(5): 1677–1707, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/60528.

Gross:2005:WNN

[550] L. K. Gross and J. Yu. Weakly nonlinear and numerical analyses of dynamics in a solid combustion model. SIAM Journal on Applied Mathematics, 65(5):1708– 1725, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/43950.

Garcia-Cervera:2005:NWL

[551] Carlos J. Garcia-Cervera. Néel walls in low anisotropy symmetric double layers. SIAM Journal on Applied Mathematics, 65(5):1726–1747, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/43776.

Lou:2005:ESD

[552] Yuan Lou, Qing Nie, and Frederic Y. M. Wan. Effects of sog on dppreceptor binding. SIAM Journal on Applied Mathematics, 65(5):1748–1771, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/43321.

Ouahsine:2005:ABI

[553] A. Ouahsine and P. A. Bois. Asymptotic behavior of internal Rossby waves with a sharp density interface. SIAM Journal on Applied Mathematics, 65(5):1772– 1793, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/41473.

Alexander:2005:BAS

[554] M. E. Alexander and S. M. Moghadas.

Bifurcation analysis of an SIRS epidemic model with generalized incidence. SIAM Journal on Applied Mathematics, 65(5): 1794–1816, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/60494.

Chan:2005:ATV

[555] Tony F. Chan and Selim Esedoglu. Aspects of total variation regularized L^{1} function approximation. SIAM Journal on Applied Mathematics, 65(5):1817– 1837, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/60429.

Cogan:2005:CFG

[556] N. G. Cogan and James P. Keener.

Channel formation in gels. SIAM Journal on Applied Mathematics, 65(6): 1839–1854, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/60551.

Youngquist:2005:FDF

[557] Robert C. Youngquist, Philip T. Metzger, and Kelly N. Kilts. Force density function relationships in 2-D granular media. SIAM Journal on Applied Mathematics, 65(6):1855–1869, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/61221.

Andreasen:2005:SOA

[558] Viggo Andreasen and Thomas Frommelt. A school-oriented, age-structured epidemic model. SIAM Journal on Applied Mathematics, 65(6):1870–1887, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/61068.

Maier:2005:WMR

[559] Thorsten Maier. Wavelet Mie representations for solenoidal vector fields with applications to ionospheric geomagnetic data. SIAM Journal on Applied Mathematics, 65(6):1888–1912, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/60379.

Du:2005:RTI

[560] Qiang Du, Chun Liu, and Xiaoqiang Wang. Retrieving topological information for phase field models. SIAM Journal on Applied Mathematics, 65(6): 1913–1932, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/60641.

Freire:2005:FCL

[561] Emilio Freire. The focus-center-limit cycle bifurcation in symmetric 3D piecewise linear systems. SIAM Journal on Applied Mathematics, 65(6):1933– 1951, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/60610.

Ibragimov:2005:MMP

[562] Akif Ibragimov, Dinara Khalmanova, Peter P. Valko, and Jay R. Walton. On a mathematical model of the productivity index of a well from reservoir engineering. SIAM Journal on Applied Mathematics, 65(6):1952–1980, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/60765.

Rienstra:2005:WHE

[563] Sjoerd W. Rienstra. Webster’s Horn equation revisited. SIAM Journal on Applied Mathematics, 65(6):1981–2004, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/41304.

Nykamp:2005:RPC

[564] Duane Q. Nykamp. Revealing pairwise coupling in linear-nonlinear networks. SIAM Journal on Applied Mathematics, 65(6):2005–2032, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/43707.

Kartchevski:2005:MAG

[565] E. M. Kartchevski, A. I. Nosich, and G. W. Hanson. Mathematical analysis of the generalized natural modes of an inhomogeneous optical fiber. SIAM Journal on Applied Mathematics, 65(6): 2033–2048, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/60437.

Bao:2005:IMS

[566] Gang Bao and Peijun Li. Inverse medium scattering problems for electromagnetic waves. SIAM Journal on Applied Mathematics, 65(6):2049–2066, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/60743.

Folias:2005:SLT

[567] Stefanos E. Folias and Paul C. Bressloff. Stimulus-locked traveling waves and breathers in an excitatory neural network. SIAM Journal on Applied Mathematics, 65(6):2067–2092, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/61517.

Lim:2005:EMN

[568] Chjan C. Lim. Energy maximizers, negative temperatures, and robust symmetry breaking in vortex dynamics on a nonrotating sphere. SIAM Journal on Applied Mathematics, 65(6):2093– 2106, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/60591.

Ammari:2005:RSI

[569] Habib Ammari. Reconstruction of a small inclusion in a two-dimensional open waveguide. SIAM Journal on Applied Mathematics, 65(6):2107–2127, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/61538.

Goldsztein:2005:TNB

[570] Guillermo H. Goldsztein. Transport of nutrients in bones. SIAM Journal on Applied Mathematics, 65(6):2128– 2140, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/61663.

Biscari:2005:FIM

[571] Paolo Biscari and Timothy J. Sluckin.

Field-induced motion of nematic disclinations. SIAM Journal on Applied Mathematics, 65(6):2141–2157, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/61889.

Glebov:2005:SPT

[572] Sergei Glebov, Oleg Kiselev, and Vladimir Lazarev. Slow passage through resonance for a weakly nonlinear dispersive wave. SIAM Journal on Applied Mathematics, 65(6):2158–2177, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/61808.

Besse:2005:IIP

[573] Christophe Besse, Jean Claudel, Pierre Degond, Fabrice Deluzet, Gérard Gallice, and Christian Tessieras. Instability of the ionospheric plasma: Modeling and analysis. SIAM Journal on Applied Mathematics, 65(6):2178– 2198, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/60658.

Camassa:2005:TRS

[574] Roberto Camassa, Alp Findikoglu, and Grant Lythe. Transmission, reflection, and second-harmonic generation in a nonlinear waveguide. SIAM Journal on Applied Mathematics, 66(1):1– 28, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/61518.

Carles:2005:SCL

[575] Rémi Carles, Norbert J. Mauser, and Hans Peter Stimming. (semi)classical limit of the Hartree equation with harmonic potential. SIAM Journal on Applied Mathematics, 66(1):29–56, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/60973.

Bressloff:2005:WIP

[576] Paul C. Bressloff. Weakly interacting pulses in synaptically coupled neural media. SIAM Journal on Applied Mathematics, 66(1):57–81, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/61637.

Knessl:2005:BPU

[577] Charles Knessl and John A. Morrison. Blocking probabilities for an underloaded or overloaded link with trunk reservation. SIAM Journal on Applied Mathematics, 66(1):82–97, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/42659.

Andersson:2005:NCH

[578] Sean B. Andersson. Nonadiabatic corrections to the Hannay–Berry phase. SIAM Journal on Applied Mathematics, 66(1):98–121, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/60660.

Pierre:2005:CGV

[579] C. Pierre, F. Plouraboue, and M. Quintard. Convergence of the generalized volume averaging method on a convection-diffusion problem: a spectral perspective. SIAM Journal on Applied Mathematics, 66(1):122–152, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/61001.

Figueiredo:2005:SAA

[580] Isabel N. Figueiredo, Carlos F. Leal, and Cecília S. Pinto. Shape analysis of an adaptive elastic rod model. SIAM Journal on Applied Mathematics, 66(1):153– 173, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/60444.

Blyth:2005:MSL

[581] M. G. Blyth and J.-M. Vanden-Broeck.

Magnetic shaping of a liquid metal column and deformation of a bubble in a vortex flow. SIAM Journal on Applied Mathematics, 66(1):174– 186, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/62235.

Moskow:2005:AMS

[582] S. Moskow, F. Santosa, and J. Zhang. An approximate method for scattering by thin structures. SIAM Journal on Applied Mathematics, 66(1):187– 205, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/61738.

Dai:2005:AOL

[583] Min Dai and Yue Kuen Kwok. American options with lookback payoff. SIAM Journal on Applied Mathematics, 66 (1):206–227, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

                                        (electronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/43734.

Goldsztein:2005:VSF

[584] Guillermo H. Goldsztein. Volume of suspension that flows through a small orifice before it clogs. SIAM Journal on Applied Mathematics, 66(1):228– 236, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/61616.

Tsai:2005:EST

[585] Je-Chiang Tsai and James Sneyd. Existence and stability of traveling waves in buffered systems. SIAM Journal on Applied Mathematics, 66(1):237– 265, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/61829.

Acebron:2005:SHE

[586] Juan A. Acebrón and Renato Spigler.

Second harmonics effects in random Duffing oscillators. SIAM Journal on Applied Mathematics, 66(1):266–285, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/43708.

Dolak:2005:KSM

[587] Yasmin Dolak and Christian Schmeiser.

The Keller–Segel model with logistic sensitivity function and small diffusivity. SIAM Journal on Applied Mathematics, 66(1):286–308, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/61284.

Guo:2005:TPV

[588] Yujin Guo, Zhenguo Pan, and M. J.

Ward. Touchdown and pull-in voltage behavior of a MEMS device with varying dielectric properties. SIAM Journal on Applied Mathematics, 66(1):309– 338, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/61339.

Schuss:2005:RSD

[589] Z. Schuss and A. Spivak. On recovering the shape of a domain from the trace of the heat kernel. SIAM Journal on Applied Mathematics, 66(1):339– 360, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/42492.

Halverson:2005:ESF

[590] Matthew A. Halverson, Howard A. Levine, and Joanna Renclawowicz. Erratum: “Singularity Formation in Chemotaxis—A Conjecture of Nagai”. SIAM Journal on Applied Mathematics, 66(1):361–364, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/63155. See [491].

Kolehmainen:2005:ICP

[591] Ville Kolehmainen, Matti Lassas, and Petri Ola. The inverse conductivity problem with an imperfectly known boundary. SIAM Journal on Applied Mathematics, 66(2):365–383, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/61273.

Stewart:2005:TDE

[592] D. Scott Stewart and Aslan R. Kasimov. Theory of detonation with an embedded sonic locus. SIAM Journal on Applied Mathematics, 66(2):384– 407, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/61693.

vandenBerg:2005:SLH

[593] Jan Bouwe van den Berg, Claude-Michel Brauner, Josephus Hulshof, and Alessandra Lunardi. The speed law for highly radiative flames in a gaseous mixture with large activation energy. SIAM Journal on Applied Mathematics, 66 (2):408–432, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/62031.

Yu:2005:PMU

[594] Hwa-Lung Yu and George Christakos. Porous media upscaling in terms of mathematical epistemic cognition. SIAM Journal on Applied Mathematics, 66(2):433–446, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/61443.

Araujo:2005:MTGb

[595] Robyn P. Araujo and D. L. Sean McElwain. A mixture theory for the genesis of residual stresses in growing tissues II: Solutions to the biphasic equations for a multicell spheroid. SIAM Journal on Applied Mathematics, 66(2):447– 467, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/60712.

Cox:2005:CDA

[596] Stephen M. Cox and Bruce H. Candy.

Class-D audio amplifiers with negative feedback. SIAM Journal on Applied Mathematics, 66(2):468–488, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/61746.

Pinfield:2005:APD

[597] V. J. Pinfield, O. G. Harlen, M. J. W.

Povey, and B. D. Sleeman. Acoustic propagation in dispersions in the long wavelength limit. SIAM Journal on Applied Mathematics, 66(2):489– 509, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/61698.

Billingham:2005:IST

[598] J. Billingham. The initial surface Tension–Driven flow of a wedge of viscous fluid. SIAM Journal on Applied Mathematics, 66(2):510–532, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/62241.

Truskinovsky:2005:KMP

[599] Lev Truskinovsky and Anna Vainchtein. Kinetics of Martensitic phase transitions: Lattice model. SIAM Journal on Applied Mathematics, 66(2):533– 553, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL http://epubs.siam.org/
                                        sam-bin/dbq/article/61694.

Ayyadi:2005:SSU

[600] Asma El Ayyadi. Semiconductor simulations using a coupled quantum drift-diffusion Schrödinger–Poisson model. SIAM Journal on Applied Mathematics, 66(2):554–572, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/61080.

Sedaghat:2005:CCO

[601] H. Sedaghat, C. M. Kent, and M. A. Wood. Criteria for the convergence, oscillation, and bistability of pulse circulation in a ring of excitable media. SIAM Journal on Applied Mathematics, 66 (2):573–590, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X

(electronic). URL http://epubs.siam.
org/sam-bin/dbq/article/61707.

Haddar:2005:CBS

[602] Houssem Haddar, Steven Kusiak, and John Sylvester. The convex backscattering support. SIAM Journal on Applied Mathematics, 66(2):591–615, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/61623.

Kon:2005:NSO

[603] Ryusuke Kon. Nonexistence of synchronous orbits and class coexistence in matrix population models. SIAM Journal on Applied Mathematics, 66(2):616– 626, 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL http://epubs.siam.org/
sam-bin/dbq/article/62353.

Hsu:2005:MIM

[604] Sze-Bi Hsu and Ying-Hen Hsieh. Modeling intervention measures and severitydependent public response during severe acute respiratory syndrome outbreak. SIAM Journal on Applied Mathematics, 66(2):627–647, January 2005. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kerschen:2005:IPE

[605] Gaetan Kerschen, Young Sup Lee, Alexander F. Vakakis, D. Michael Mc-Farland, and Lawrence A. Bergman. Irreversible passive energy transfer in coupled oscillators with essential nonlinearity. SIAM Journal on Applied Mathematics, 66(2):648–679, January 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Alber:2005:SMN

[606] Hans-Dieter Alber and Peicheng Zhu. Solutions to a model with nonuniformly parabolic terms for phase evolution driven by configurational forces. SIAM Journal on Applied Mathematics, 66(2):680–699, January 2005. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zaks:2005:PSP

[607] Michael A. Zaks, Alla Podolny, Alexander A. Nepomnyashchy, and Alexander A. Golovin. Periodic stationary patterns governed by a convective Cahn–Hilliard equation. SIAM Journal on Applied Mathematics, 66 (2):700–720, January 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yang:2005:CEC

[608] Hongtao Yang. Calibration of the extended CIR model. SIAM Journal on Applied Mathematics, 66(2):721–735, January 2005. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sivaloganathan:2006:CRM

[609] Jeyabal Sivaloganathan, Scott J. Spector, and Viveka Tilakraj. The convergence of regularized minimizers for cavitation problems in nonlinear elasticity. SIAM Journal on Applied Mathematics, 66(3):736–757, January 2006. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bao:2006:DRB

[610] Weizhu Bao, Qiang Du, and Yanzhi Zhang. Dynamics of rotating Bose– Einstein condensates and its efficient and accurate numerical computation. SIAM Journal on Applied Mathematics, 66(3):758–786, January 2006. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mohan:2006:TMS

[611] Aruna Mohan and Howard Brenner.

Thermophoretic motion of a slightly deformed sphere through a viscous fluid. SIAM Journal on Applied Mathematics, 66(3):787–801, January 2006. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Xiao:2006:MFH

[612] Dongmei Xiao and Huaiping Zhu. Multiple focus and Hopf bifurcations in a predator-prey system with nonmonotonic functional response. SIAM Journal on Applied Mathematics, 66(3):802–819, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sparber:2006:EMT

[613] Christof Sparber. Effective mass theorems for nonlinear Schrödinger equations. SIAM Journal on Applied Mathematics, 66(3):820–842, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Martcheva:2006:RCM

[614] Maia Martcheva and Sergei S. Pilyugin.

The role of coinfection in multidisease dynamics. SIAM Journal on Applied Mathematics, 66(3):843–872, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Park:2006:REN

[615] J. H. Park, A. Bayliss, B. J. Matkowsky, and A. A. Nepomnyashchy. On the route to extinction in nonadiabatic solid flames. SIAM Journal on Applied Mathematics, 66(3):873–895, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Armbruster:2006:MDL

[616] D. Armbruster, P. Degond, and C. Ringhofer. A model for the dynamics of large queuing networks and supply chains. SIAM Journal on Applied Mathematics, 66(3):896–920, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sopasakis:2006:SMS

[617] Alexandros Sopasakis and Markos A. Katsoulakis. Stochastic modeling and simulation of traffic flow: Asymmetric single exclusion process with Arrhenius look-ahead dynamics. SIAM Journal on Applied Mathematics, 66 (3):921–944, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bateman:2006:TWS

[618] Michael D. Bateman and Erik S. Van Vleck. Traveling wave solutions to a coupled system of spatially discrete Nagumo equations. SIAM Journal on Applied Mathematics, 66(3):945–976, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Quintanilla:2006:QAD

[619] Ramón Quintanilla and Reinhard Racke. Qualitative aspects in dual-phase-lag thermoelasticity. SIAM Journal on Applied Mathematics, 66(3):977–1001, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chandler-Wilde:2006:ASM

[620] Simon N. Chandler-Wilde, Eric Heinemeyer, and Roland Potthast. Acoustic scattering by mildly rough unbounded surfaces in three dimensions. SIAM Journal on Applied Mathematics, 66 (3):1002–1026, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Banks:2006:MSD

[621] H. T. Banks, K. Ito, G. M. Kepler, and J. A. Toivanen. Material surface design to counter electromagnetic interrogation of targets. SIAM Journal on Applied Mathematics, 66(3):1027–1049, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cooke:2006:PPI

[622] Kenneth L. Cooke, Richard H. Elderkin, and Wenzhang Huang. Predator-prey interactions with delays due to juvenile maturation. SIAM Journal on Applied Mathematics, 66(3):1050–1079, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ren:2006:ESS

[623] Xiaofeng Ren and Juncheng Wei. Existence and stability of spherically layered solutions of the diblock copolymer equation. SIAM Journal on Applied Mathematics, 66(3):1080–1099, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Liu:2006:SSP

[624] Shengqiang Liu and Edoardo Beretta. A stage-structured predator-prey model of Beddington–DeAngelis type. SIAM Journal on Applied Mathematics, 66 (4):1101–1129, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fried:2006:RBS

[625] Eliot Fried. On the relationship between supplemental balances in two theories for pure interface motion. SIAM Journal on Applied Mathematics, 66 (4):1130–1149, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Siebel:2006:FDT

[626] Florian Siebel and Wolfram Mauser. On the fundamental diagram of traffic flow. SIAM Journal on Applied Mathematics, 66(4):1150–1162, January 2006. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gurski:2006:ECL

[627] K. F. Gurski, G. B. McFadden, and M. J. Miksis. The effect of contact lines on the Rayleigh instability with anisotropic surface energy. SIAM Journal on Applied Mathematics, 66 (4):1163–1187, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Landis:2006:EPN

[628] S. Landis, B. Z. Bobrovsky, and Z. Schuss. The exit problem in a nonlinear system driven by 1/f noise: The delay locked loop. SIAM Journal on Applied Mathematics, 66(4):

1188–1208, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Brugger:2006:MIS

[629] S. C. Brugger, A. Schenk, and W. Fichtner. Moments of the inverse scattering operator of the Boltzmann equation: Theory and applications. SIAM Journal on Applied Mathematics, 66 (4):1209–1226, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cui:2006:WPC

[630] Zhenlu Cui, M. Gregory Forest, Qi Wang, and Hong Zhou. On weak plane Couette and Poiseuille flows of rigid rod and platelet ensembles. SIAM Journal on Applied Mathematics, 66 (4):1227–1260, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Renardy:2006:DRR

[631] Michael Renardy. Draw resonance revisited. SIAM Journal on Applied Mathematics, 66(4):1261–1269, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zabarankin:2006:HFA

[632] Michael Zabarankin and Andrei F.

Ulitko. Hilbert formulas for r-analytic functions in the domain exterior to spindle. SIAM Journal on Applied Mathematics, 66(4):1270–1300, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cheng:2006:MRN

[633] Chang-Yuan Cheng, Kuang-Hui Lin, and Chih-Wen Shih. Multistability in recurrent neural networks. SIAM Journal on Applied Mathematics, 66 (4):1301–1320, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Craciun:2006:MEC

[634] Gheorghe Craciun and Martin Feinberg. Multiple equilibria in complex chemical reaction networks: II. The species-reaction graph. SIAM Journal on Applied Mathematics, 66(4): 1321–1338, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Baer:2006:MBA

[635] S. M. Baer, B. W. Kooi, Yu. A. Kuznetsov, and H. R. Thieme. Multiparametric bifurcation analysis of a basic two-stage population model. SIAM Journal on Applied Mathematics, 66 (4):1339–1365, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kirkland:2006:EDP

[636] Stephen Kirkland, Chi-Kwong Li, and Sebastian J. Schreiber. On the evolution of dispersal in patchy landscapes. SIAM Journal on Applied Mathematics, 66(4):1366–1382, January 2006. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chen:2006:VEL

[637] Yunmei Chen, Stacey Levine, and Murali Rao. Variable exponent, linear growth functionals in image restoration. SIAM Journal on Applied Mathematics, 66(4):1383–1406, January 2006. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fister:2006:ITD

[638] K. Renee Fister, Maeve L. McCarthy, and Seth F. Oppenheimer. The identification of a time dependent sorption parameter from soil column experiments. SIAM Journal on Applied Mathematics, 66(4):1407–1423, January 2006. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Boldin:2006:IPS

[639] Barbara Boldin. Introducing a population into a steady community: The critical case, the center manifold, and the direction of bifurcation. SIAM Journal on Applied Mathematics, 66 (4):1424–1453, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wang:2006:EMP

[640] Wendi Wang and Xiao-Qiang Zhao. An epidemic model with population dispersal and infection period. SIAM Journal on Applied Mathematics, 66 (4):1454–1472, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dkhil:2006:TFP

[641] Fathi Dkhil and K. P. Hadeler. Traveling fronts in pressure-driven combustion. SIAM Journal on Applied Mathematics, 66(5):1473–1481, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cuesta:2006:SWT

[642] C. M. Cuesta and J. R. King. Small- and waiting-time behavior of a Darcy flow model with a dynamic pressure saturation relation. SIAM Journal on Applied Mathematics, 66(5):1482–1511, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schaerer:2006:PHG

[643] C. E. Schaerer, D. Marchesin, M. Sarkis, and P. Bedrikovetsky. Permeability hysteresis in gravity counterflow segregation. SIAM Journal on Applied Mathematics, 66(5):1512–1532, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bohun:2006:STS

[644] C. Sean Bohun, Huaxiong Huang, Ian Frigaard, and Shuqing Liang. A semianalytical thermal stress model for the Czochralski growth of Type III-V compounds. SIAM Journal on Applied Mathematics, 66(5):1533–1562, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yakubu:2006:DTS

[645] Abdul-Aziz Yakubu and John E. Franke. Discrete-time SIS epidemic model in a seasonal environment. SIAM Journal on Applied Mathematics, 66(5): 1563–1587, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shearer:2006:MTL

[646] M. Shearer and R. Levy. The motion of a thin liquid film driven by surfactant and gravity. SIAM Journal on Applied Mathematics, 66(5):1588–1609, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Munch:2006:IAF

[647] Andreas Münch and P. L. Evans. Interaction of advancing fronts and meniscus profiles formed by surface-tension-gradient-driven liquid films. SIAM Journal on Applied Mathematics, 66 (5):1610–1631, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nikolova:2006:AFG

[648] Mila Nikolova, Selim Esedoglu, and Tony F. Chan. Algorithms for finding global minimizers of image segmentation and denoising models. SIAM Journal on Applied Mathematics, 66 (5):1632–1648, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Linton:2006:MSM

[649] C. M. Linton and P. A. Martin. Multiple scattering by multiple spheres: a new proof of the Lloyd–Berry formula for the effective wavenumber. SIAM Journal on Applied Mathematics, 66 (5):1649–1668, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gleeson:2006:NLS

[650] J. P. Gleeson and F. O’Doherty. Non-Lorentzian spectral lineshapes near a Hopf bifurcation. SIAM Journal on Applied Mathematics, 66(5):1669–1688, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Saxton:2006:PTC

[651] Katarzyna Saxton and Ralph Saxton.

Phase transitions and change of type in low-temperature heat propagation. SIAM Journal on Applied Mathematics, 66(5):1689–1702, January 2006. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Marheineke:2006:FDT

[652] Nicole Marheineke and Raimund Wegener. Fiber dynamics in turbulent flows: General modeling framework. SIAM Journal on Applied Mathematics, 66(5):1703–1726, January 2006. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bowen:2006:LLD

[653] Mark Bowen and Thomas P. Witelski. The linear limit of the dipole problem for the thin film equation. SIAM Journal on Applied Mathematics, 66 (5):1727–1748, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lee:2006:EAB

[654] Sunmi Lee, Russel E. Caflisch, and Young-Ju Lee. Exact artificial boundary conditions for continuum and discrete elasticity. SIAM Journal on Applied Mathematics, 66(5):1749–1775, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cain:2006:DBP

[655] John W. Cain. Dynamic behavior of a paced cardiac fiber. SIAM Journal on Applied Mathematics, 66(5): 1776–1792, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Horak:2006:CBM

[656] Jiří Horák, Gabriel J. Lord, and Mark A. Peletier. Cylinder buckling: The mountain pass as an organizing center. SIAM Journal on Applied Mathematics, 66 (5):1793–1824, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Arora:2006:CSS

[657] Rajan Arora and V. D. Sharma. Convergence of strong shock in a Van der Waals gas. SIAM Journal on Applied Mathematics, 66(5):1825–1837, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Humi:2006:SLE

[658] Mayer Humi. On the solution of Long’s Equation with shear. SIAM Journal on Applied Mathematics, 66 (6):1839–1852, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Simons:2006:ECS

[659] R. R. L. Simons and S. A. Gourley. Extinction criteria in stage-structured population models with impulsive culling. SIAM Journal on Applied Mathematics, 66(6):1853–1870, January 2006. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dominici:2006:RSS

[660] Diego Dominici and Charles Knessl. Ray solution of a singularly perturbed elliptic PDE with applications to communications networks. SIAM Journal on Applied Mathematics, 66(6): 1871–1894, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rosencrans:2006:SDE

[661] Steve Rosencrans and Xuefeng Wang.

Suppression of the Dirichlet eigenvalues of a coated body. SIAM Journal on Applied Mathematics, 66(6): 1895–1916, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See corrigendum [816].

Griffiths:2006:RMC

[662] Roger E. Griffiths and Mark Pernarowski. Return map characterizations for a model of bursting with two slow variables. SIAM Journal on Applied Mathematics, 66(6):1917–1948, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Billingham:2006:STD

[663] J. Billingham. Surface tension-driven flow in a slender wedge. SIAM Journal on Applied Mathematics, 66(6): 1949–1977, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Edwards:2006:CET

[664] David A. Edwards. Convection effects in thin reaction zones: Applications to BI-Acore. SIAM Journal on Applied Mathematics, 66(6):1978–1997, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Smith:2006:MDA

[665] Ralph C. Smith, Andrew G. Hatch, Tathagata De, Murti V. Salapaka, Ricardo C. H. del Rosario, and Julie K. Raye. Model development for atomic force microscope stage mechanisms. SIAM Journal on Applied Mathematics, 66(6):1998–2026, January 2006. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Drobnjak:2006:OMM

[666] Ivana Drobnjak, A. C. Fowler, and Michael C. Mackey. Oscillations in a maturation model of blood cell production. SIAM Journal on Applied Mathematics, 66(6):2027–2048, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ambrosi:2006:CTI

[667] D. Ambrosi. Cellular traction as an inverse problem. SIAM Journal on Applied Mathematics, 66(6):2049–2060, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bouchitte:2006:HWP

[668] Guy Bouchitté and Didier Felbacq. Homogenization of a wire photonic crystal: The case of small volume fraction. SIAM Journal on Applied Mathematics, 66(6):2061–2084, January 2006. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pardo:2006:TDH

[669] D. Pardo, L. Demkowicz, C. Torres-Verdín, and M. Paszynski. Twodimensional high-accuracy simulation of resistivity logging-while-drilling (LWD) measurements using a self-adaptive goaloriented hp finite element method. SIAM Journal on Applied Mathematics, 66(6):2085–2106, January 2006. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Park:2006:ANE

[670] Jinhae Park and M. Carme Calderer.

Analysis of nonlocal electrostatic effects in chiral smectic C liquid crystals.

SIAM Journal on Applied Mathematics, 66(6):2107–2126, January 2006. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Droske:2006:MSL

[671] Marc Droske and Wolfgang Ring. A Mumford–Shah level-set approach for geometric image registration. SIAM Journal on Applied Mathematics, 66 (6):2127–2148, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Knackfuss:2006:TJP

[672] Rosenei Felippe Knackfuss and Liliane Basso Barichello. On the temperature-jump problem in rarefied gas dynamics: The effect of the Cercignani–Lampis boundary condition. SIAM Journal on Applied Mathematics, 66(6):2149–2186, January 2006. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gasser:2006:AOD

[673] Ingenuin Gasser and Herbert Steinrück. An asymptotic one-dimensional model to describe fires in tunnels III: The transient problem. SIAM Journal on Applied Mathematics, 66(6):2187–2203, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ruan:2006:EGS

[674] Shigui Ruan and Xue-Zhong He. Erratum: Global stability in chemostattype competition models with nutrient recycling. SIAM Journal on Applied Mathematics, 66(6):2204–2205, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [1].

Appelo:2006:PML

[675] Daniel Appelö, Thomas Hagstrom, and Gunilla Kreiss. Perfectly matched layers for hyperbolic systems: General formulation, well-posedness, and stability. SIAM Journal on Applied Mathematics, 67(1):1–23, January 2006. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chitnis:2006:BAM

[676] Nakul Chitnis, J. M. Cushing, and J. M.

Hyman. Bifurcation analysis of a mathematical model for malaria transmission. SIAM Journal on Applied Mathematics, 67(1):24–45, January 2006. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jungel:2006:DNQ

[677] Ansgar Jüngel, Daniel Matthes, and Josipa Pina Mili i. Derivation of new quantum hydrodynamic equations using entropy minimization. SIAM Journal on Applied Mathematics, 67 (1):46–68, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

White:2006:EPL

[678] Benjamin S. White and Minyao Zhou.

Electroseismic prospecting in layered media. SIAM Journal on Applied Mathematics, 67(1):69–98, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gong:2006:QLC

[679] Qiang Gong and Rajan Batta. A queue-length cutoff model for a preemptive two-priority M/M/1 system.

SIAM Journal on Applied Mathematics, 67(1):99–115, January 2006. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

vandenBerg:2006:FBF

[680] Jan Bouwe van den Berg, Vincent Guyonne, and Josephus Hulshof. Flame balls for a free boundary combustion model with radiative transfer. SIAM Journal on Applied Mathematics, 67 (1):116–137, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ou:2006:SSR

[681] Chunhua Ou and Jianhong Wu. Spatial spread of rabies revisited: Influence of age-dependent diffusion on nonlinear dynamics. SIAM Journal on Applied Mathematics, 67(1):138–163, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Canic:2006:MVB

[682] Sunčica Çanić, Josip Tambača, Giovanna Guidoboni, Andro Mikelić, Craig J. Hartley, and Doreen Rosenstrauch. Modeling viscoelastic behavior of arterial walls and their interaction with pulsatile blood flow. SIAM Journal on Applied Mathematics, 67 (1):164–193, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Engstrom:2006:IBB

[683] Christian Engström. Inverse bounds and bulk properties of complex-valued twocomponent composites. SIAM Journal on Applied Mathematics, 67(1):194–213, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hwang:2006:CMT

[684] Dae Kun Hwang and Alejandro D. Rey.

Computational modeling of texture formation and optical performance of liquid crystal films on patterned surfaces. SIAM Journal on Applied Mathematics, 67(1):214–234, January 2006. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kamotski:2006:BBA

[685] V. V. Kamotski, L. Ju. Fradkin, B. A.

Samokish, V. A. Borovikov, and V. M. Babich. On Budaev and Bogy’s approach to diffraction by the 2D tractionfree elastic wedge. SIAM Journal on Applied Mathematics, 67(1):235–259, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Iggidr:2006:GAN

[686] Abderrhaman Iggidr, Jean-Claude Kamgang, Gauthier Sallet, and Jean-Jules Tewa. Global analysis of new malaria intrahost models with a competitive exclusion principle. SIAM Journal on Applied Mathematics, 67 (1):260–278, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Haber:2006:SPM

[687] Shimon Haber. A spherical particle moving slowly in a fluid with a radially varying viscosity. SIAM Journal on Applied Mathematics, 67(1):279–304, January 2006. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Erneux:2007:RQS

[688] Thomas Erneux and Albert Goldbeter. Rescue of the quasi-steady-state approximation in a model for oscillations in an enzymatic cascade. SIAM Journal on Applied Mathematics, 67 (2):305–320, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tesdall:2007:TPP

[689] Allen M. Tesdall, Richard Sanders, and Barbara L. Keyfitz. The triple point paradox for the nonlinear wave system. SIAM Journal on Applied Mathematics, 67(2):321–336, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Georgescu:2007:GSV

[690] Paul Georgescu and Ying-Hen Hsieh. Global stability for a virus dynamics model with nonlinear incidence of infection and removal. SIAM Journal on Applied Mathematics, 67(2):337–353, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Prusa:2007:SCM

[691] Vít Průša. Sufficient conditions for monotone linear stability of steady and oscillatory Hagen–Poiseuille flow. SIAM Journal on Applied Mathematics, 67 (2):354–363, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mitchell:2007:AFF

[692] S. L. Mitchell, R. Kuske, and A. P. Peirce. An asymptotic framework for finite hydraulic fractures including leakoff. SIAM Journal on Applied Mathematics, 67(2):364–386, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lei:2007:SDD

[693] Jinzhi Lei and Michael C. Mackey.

Stochastic differential delay equation, moment stability, and application to hematopoietic stem cell regulation system. SIAM Journal on Applied Mathematics, 67(2):387–407, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gourley:2007:SVB

[694] Stephen A. Gourley, Rongsong Liu, and Jianhong Wu. Some vector borne diseases with structured host populations: Extinction and spatial spread. SIAM Journal on Applied Mathematics, 67(2):408–433, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Flores:2007:ADT

[695] G. Flores, G. Mercado, J. A. Pelesko, and N. Smyth. Analysis of the dynamics and touchdown in a model of electrostatic MEMS. SIAM Journal on Applied Mathematics, 67(2):434–446, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Biscari:2007:BRE

[696] Paolo Biscari and Stefano Turzi.

Boundary-roughness effects in nematic liquid crystals. SIAM Journal on Applied Mathematics, 67(2):447–463, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Balasuriya:2007:HLN

[697] Sanjeeva Balasuriya, Georg Gottwald, John Hornibrook, and Stéphane Lafortune. High Lewis number combustion wavefronts: a perturbative Melnikov analysis. SIAM Journal on Applied Mathematics, 67(2):464–486, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Evje:2007:WST

[698] Steinar Evje and Tore Flåtten. On the wave structure of two-phase flow models. SIAM Journal on Applied Mathematics, 67(2):487–511, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kazanci:2007:PFA

[699] Fatma Gurel Kazanci and Bard Ermentrout. Pattern formation in an array of oscillators with electrical and chemical coupling. SIAM Journal on Applied Mathematics, 67(2):512–529, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pedersen:2007:ENC

[700] Morten Gram Pedersen and Mads Peter Sørensen. The effect of noise on βcell burst period. SIAM Journal on Applied Mathematics, 67(2):530–542, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Huntley:2007:OAMa

[701] Helga S. Huntley, Esteban G. Tabak, and Edward H. Suh. An optimization approach to modeling sea ice dynamics. part 1: Lagrangian framework. SIAM Journal on Applied Mathematics, 67(2):543–560, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Huntley:2007:OAMb

[702] Helga S. Huntley and Esteban G. Tabak.

An optimization approach to modeling sea ice dynamics, part 2: Finite ice strength effects. SIAM Journal on Applied Mathematics, 67(2):561–581, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Frank:2007:FAM

[703] Martin Frank, Hartmut Hensel, and Axel Klar. A fast and accurate moment method for the Fokker–Planck equation and applications to electron radiotherapy. SIAM Journal on Applied Mathematics, 67(2):582–603, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ou:2007:NAA

[704] Zhong-Hui Ou, Shi-Qiang Dai, Peng Zhang, and Li-Yun Dong. Nonlinear analysis in the Aw–Rascle anticipation model of traffic flow. SIAM Journal on Applied Mathematics, 67 (3):605–618, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Artiles:2007:EEF

[705] William Artiles and Roberto A.

Kraenkel. An exact equation for the free surface of a fluid in a porous medium. SIAM Journal on Applied Mathematics, 67(3):619–629, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Huang:2007:FGM

[706] Huaxiong Huang, Jonathan J. Wylie, Robert M. Miura, and Peter D. Howell. On the formation of glass microelectrodes. SIAM Journal on Applied Mathematics, 67(3):630–666, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Valadkhan:2007:PMM

[707] Sina Valadkhan, Kirsten Morris, and Amir Khajepour. Passivity of magnetostrictive materials. SIAM Journal on Applied Mathematics, 67(3):667–686, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shipman:2007:GMP

[708] Stephen Shipman and Darko Volkov. Guided modes in periodic slabs: Existence and nonexistence. SIAM Journal on Applied Mathematics, 67(3):687–713, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yun:2007:EEF

[709] KiHyun Yun. Estimates for electric fields blown up between closely adjacent conductors with arbitrary shape. SIAM Journal on Applied Mathematics, 67(3):714–730, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rong:2007:MAA

[710] Libin Rong, Zhilan Feng, and Alan S. Perelson. Mathematical analysis of agestructured HIV-1 dynamics with combination antiretroviral therapy. SIAM Journal on Applied Mathematics, 67 (3):731–756, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Li:2007:AMG

[711] Jiaxu Li and Yang Kuang. Analysis of a model of the glucose-insulin regulatory system with two delays. SIAM Journal on Applied Mathematics, 67 (3):757–776, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Amar:2007:NSA

[712] Chokri Ben Amar, Nabil Gmati, Christophe Hazard, and Karim Ramdani. Numerical simulation of acoustic time reversal mirrors. SIAM Journal on Applied Mathematics, 67(3):777–791, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Golovaty:2007:SFR

[713] Dmitry Golovaty. On step-function reaction kinetics model in the absence of material diffusion. SIAM Journal on Applied Mathematics, 67(3):792–809, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Caputo:2007:SPJ

[714] J. G. Caputo and L. Loukitch. Statics of point Josephson junctions in a microstrip line. SIAM Journal on Applied Mathematics, 67(3):810–836, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Aletti:2007:FOC

[715] Giacomo Aletti, Giovanni Naldi, and Giuseppe Toscani. First-order continuous models of opinion formation.

SIAM Journal on Applied Mathematics, 67(3):837–853, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Guyonne:2007:MFB

[716] Vincent Guyonne and Pascal Noble. On a model of flame ball with radiative transfer. SIAM Journal on Applied Mathematics, 67(3):854–868, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Deseri:2007:MFE

[717] L. Deseri and J. M. Golden. The minimum free energy for continuous spectrum materials. SIAM Journal on Applied Mathematics, 67(3):869–892, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Knudsen:2007:BME

[718] Kim Knudsen, Matti Lassas, Jennifer L. Mueller, and Samuli Siltanen. D-bar method for electrical impedance tomography with discontinuous conductivities. SIAM Journal on Applied Mathematics, 67(3):893–913, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mahaffy:2007:MCW

[719] Joseph M. Mahaffy and Leah Edelstein-Keshet. Modeling cyclic waves of circulating T cells in autoimmune diabetes. SIAM Journal on Applied Mathematics, 67(4):915–937, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Williams:2007:WSS

[720] Timothy D. Williams and Vernon A. Squire. Wave scattering at the sea-ice/ ice-shelf transition with other applications. SIAM Journal on Applied Mathematics, 67(4):938–959, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Burger:2007:IPR

[721] Martin Burger, Robert S. Eisenberg, and Heinz W. Engl. Inverse problems related to ion channel selectivity. SIAM Journal on Applied Mathematics, 67(4):960–989, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Perez-Garcia:2007:MMN

[722] Víctor M. Pérez-García, Pedro J. Torres, and Gaspar D. Montesinos. The method of moments for nonlinear Schrödinger equations: Theory and applications. SIAM Journal on Applied Mathematics, 67(4):990–1015, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fowler:2007:FRC

[723] A. C. Fowler, Natalia Kopteva, and Charles Oakley. The formation of river channels. SIAM Journal on Applied Mathematics, 67(4):1016–1040, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Garcia:2007:CFB

[724] R. D. M. Garcia and C. E. Siewert. Channel flow of a binary mixture of rigid spheres described by the linearized Boltzmann equation and driven by temperature, pressure, and concentration gradients. SIAM Journal on Applied Mathematics, 67(4):1041–1063, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Skeldon:2007:PSF

[725] A. C. Skeldon and G. Guidoboni. Pattern selection for Faraday waves in an incompressible viscous fluid. SIAM Journal on Applied Mathematics, 67 (4):1064–1100, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hyvonen:2007:LTR

[726] Nuutti Hyvönen. Locating transparent regions in optical absorption and scattering tomography. SIAM Journal on Applied Mathematics, 67(4): 1101–1123, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Liu:2007:RSS

[727] J. J. Liu, G. Nakamura, and M. Sini. Reconstruction of the shape and surface impedance from acoustic scattering data for an arbitrary cylinder. SIAM Journal on Applied Mathematics, 67 (4):1124–1146, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Goldstein:2007:DLR

[728] Byron Goldstein, Harold Levine, and David Torney. Diffusion limited reactions. SIAM Journal on Applied Mathematics, 67(4):1147–1165, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bradshaw-Hajek:2007:CEF

[729] B. H. Bradshaw-Hajek, Y. M. Stokes, and E. O. Tuck. Computation of extensional fall of slender viscous drops by a one-dimensional Eulerian method. SIAM Journal on Applied Mathematics, 67(4):1166–1182, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Goudon:2007:SRH

[730] Thierry Goudon, Vera Miljanović, and Christian Schmeiser. On the Shockley– Read–Hall model: Generation-recombination in semiconductors. SIAM Journal on Applied Mathematics, 67(4):1183–1201, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Noble:2007:RWG

[731] Pascal Noble. Roll-waves in general hyperbolic systems with source terms. SIAM Journal on Applied Mathematics, 67(4):1202–1212, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jung:2007:MIS

[732] Yoon Mo Jung, Sung Ha Kang, and Jianhong Shen. Multiphase image segmentation via Modica–Mortola phase transition. SIAM Journal on Applied Mathematics, 67(5):1213–1232, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Linton:2007:SSI

[733] C. M. Linton, R. Porter, and I. Thompson. Scattering by a semi-infinite periodic array and the excitation of surface waves. SIAM Journal on Applied Mathematics, 67(5):1233–1258, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Liu:2007:CHA

[734] J. J. Liu, J. K. Seo, M. Sini, and E. J. Woo. On the convergence of the harmonic B_{z} algorithm in magnetic resonance electrical impedance tomography. SIAM Journal on Applied Mathematics, 67(5):1259–1282, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Allen:2007:APS

[735] L. J. S. Allen, B. M. Bolker, Y. Lou, and A. L. Nevai. Asymptotic profiles of the steady states for an SIS epidemic patch model. SIAM Journal on Applied Mathematics, 67(5):1283–1309, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tseluiko:2007:NDE

[736] Dmitri Tseluiko and Demetrios T. Papageorgiou. Nonlinear dynamics of electrified thin liquid films. SIAM Journal on Applied Mathematics, 67 (5):1310–1329, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fata:2007:ESR

[737] S. Nintcheu Fata and B. B. Guzina. Elastic scatterer reconstruction via the adjoint sampling method. SIAM Journal on Applied Mathematics, 67 (5):1330–1352, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Devaney:2007:ISP

[738] Anthony J. Devaney, Edwin A. Marengo, and Mei Li. Inverse source problem in nonhomogeneous background media. SIAM Journal on Applied Mathematics, 67(5):1353–1378, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Georgescu:2007:GDP

[739] Paul Georgescu and Ying-Hen Hsieh.

Global dynamics of a predator-prey model with stage structure for the predator. SIAM Journal on Applied Mathematics, 67(5):1379–1395, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Golubitsky:2007:SSD

[740] Martin Golubitsky, LieJune Shiau, and Ian Stewart. Spatiotemporal symmetries in the disynaptic canal-neck projection. SIAM Journal on Applied Mathematics, 67(5):1396–1417, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Moore:2007:MCS

[741] R. O. Moore, G. Biondini, and W. L.

Kath. A method to compute statistics of large, noise-induced perturbations of nonlinear Schrödinger solitons. SIAM Journal on Applied Mathematics, 67(5):1418–1439, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kolehmainen:2007:ICP

[742] Ville Kolehmainen, Matti Lassas, and Petri Ola. The inverse conductivity problem with an imperfectly known boundary in three dimensions. SIAM Journal on Applied Mathematics, 67 (5):1440–1452, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Li:2007:HBM

[743] Bingtuan Li and Yang Kuang. Heteroclinic bifurcation in the Michaelis– Menten-type ratio-dependent predatorprey system. SIAM Journal on Applied Mathematics, 67(5):1453–1464, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chadwick:2007:LSB

[744] Edmund Chadwick and Nina Fishwick. Lift on slender bodies with elliptical cross section evaluated by using an Oseen flow model. SIAM Journal on Applied Mathematics, 67 (5):1465–1478, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Peng:2007:SPR

[745] Rui Peng, Junping Shi, and Mingxin Wang. Stationary pattern of a ratiodependent food chain model with diffusion. SIAM Journal on Applied Mathematics, 67(5):1479–1503, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See erratum [947].

Muraki:2007:SIW

[746] David J. Muraki. A simple illustration of a weak spectral cascade. SIAM Journal on Applied Mathematics, 67 (5):1504–1521, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Banaji:2007:MPI

[747] Murad Banaji, Pete Donnell, and Stephen Baigent. P matrix properties, injectivity, and stability in chemical reaction systems. SIAM Journal on Applied Mathematics, 67(6): 1523–1547, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tsynkov:2007:WLE

[748] S. V. Tsynkov. Weak lacunae of electromagnetic waves in dilute plasma. SIAM Journal on Applied Mathematics, 67(6):1548–1581, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Peterson:2007:ACS

[749] A. W. Peterson and S. V. Tsynkov. Active control of sound for composite regions. SIAM Journal on Applied Mathematics, 67(6):1582–1609, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Morrison:2007:UCS

[750] John A. Morrison and Phil Whiting. On the uplink of a cellular system with imperfect power control and multiple services. SIAM Journal on Applied Mathematics, 67(6):1610–1632, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schoof:2007:CDG

[751] Christian Schoof. Cavitation on deformable glacier beds. SIAM Journal on Applied Mathematics, 67(6): 1633–1653, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Strunin:2007:ACS

[752] D. V. Strunin. Attractors in confined source problems for coupled nonlinear diffusion. SIAM Journal on Applied Mathematics, 67(6):1654–1674, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

DeVault:2007:NIC

[753] Kristen J. DeVault, Pierre A. Gremaud, and Helge Kristian Jenssen. Numerical investigation of cavitation in multidimensional compressible flows. SIAM Journal on Applied Mathematics, 67 (6):1675–1692, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yafia:2007:HBD

[754] Radouane Yafia. Hopf bifurcation in differential equations with delay for tumor–immune system competition model. SIAM Journal on Applied Mathematics, 67(6):1693–1703, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gotz:2007:SMA

[755] T. Götz, A. Klar, N. Marheineke, and R. Wegener. A stochastic model and associated Fokker–Planck equation for the fiber lay-down process in nonwoven production processes. SIAM Journal on Applied Mathematics, 67 (6):1704–1717, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Garnier:2007:PPT

[756] Josselin Garnier and George Papanicolaou. Pulse propagation and time reversal in random waveguides. SIAM Journal on Applied Mathematics, 67 (6):1718–1739, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zhang:2007:DIQ

[757] Yanzhi Zhang, Weizhu Bao, and Qiang Du. The dynamics and interaction of quantized vortices in the Ginzburg– Landau–Schrödinger equation. SIAM Journal on Applied Mathematics, 67 (6):1740–1775, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Blowey:2007:SWT

[758] James F. Blowey, John R. King, and Stephen Langdon. Small- and waitingtime behavior of the thin-film equation. SIAM Journal on Applied Mathematics, 67(6):1776–1807, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Marheineke:2007:FDT

[759] Nicole Marheineke and Raimund Wegener. Fiber dynamics in turbulent flows: Specific Taylor drag. SIAM Journal on Applied Mathematics, 68(1):1–23, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

McLaughlin:2007:ARW

[760] Joyce R. McLaughlin, Daniel Renzi, and Jeong-Rock Yoon. Anisotropy reconstruction from wave fronts in transversely isotropic acoustic media. SIAM Journal on Applied Mathematics, 68 (1):24–42, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Larsen:2007:AMA

[761] Edward W. Larsen and Liang Liang.

The atomic mix approximation for charged particle transport. SIAM Journal on Applied Mathematics, 68(1):43– 58, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Degond:2007:SDL

[762] P. Degond and C. Ringhofer. Stochastic dynamics of long supply chains with random breakdowns. SIAM Journal on Applied Mathematics, 68(1):59–79, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Destrade:2007:CRW

[763] M. Destrade and G. Saccomandi. Creep, recovery, and waves in a nonlinear fiberreinforced viscoelastic solid. SIAM Journal on Applied Mathematics, 68 (1):80–97, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Singer:2007:ATN

[764] A. Singer and Z. Schuss. Activation through a narrow opening. SIAM Journal on Applied Mathematics, 68 (1):98–108, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Maestre:2007:STD

[765] Faustino Maestre, Arnaud Münch, and Pablo Pedregal. A spatio-temporal design problem for a damped wave equation. SIAM Journal on Applied Mathematics, 68(1):109–132, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Roques:2007:PRE

[766] Lionel Roques and Mickaël D. Chekroun. On population resilience to external perturbations. SIAM Journal on Applied Mathematics, 68(1):133–153, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Greer:2007:EGI

[767] Meredith L. Greer, P. van den Driessche, Lin Wang, and G. F. Webb. Effects of general incidence and polymer joining on nucleated polymerization in a model of prion proliferation. SIAM Journal on Applied Mathematics, 68(1):154–170, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jungel:2007:HDH

[768] Ansgar Jüngel, Stefan Krause, and Paola Pietra. A hierarchy of diffusive higher-order moment equations for semiconductors. SIAM Journal on Applied Mathematics, 68(1):171–198, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rufty:2007:DIB

[769] Alan Rufty. A Dirichlet-integral–based dual-access collocation-kernel approach to point source gravity-field modeling. SIAM Journal on Applied Mathematics, 68(1):199–221, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Huang:2007:MTD

[770] H. Huang, P. Lin, and W. Zhou.

Moisture transport and diffusive instability during bread baking. SIAM Journal on Applied Mathematics, 68 (1):222–238, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hohenegger:2007:SRV

[771] Christel Hohenegger and Peter J. Mucha. Statistical reconstruction of velocity profiles for nanoparticle image velocimetry. SIAM Journal on Applied Mathematics, 68(1):239–252, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hayakawa:2007:CFA

[772] Carole K. Hayakawa, Jerome Spanier, and Vasan Venugopalan. Coupled forward-adjoint Monte Carlo simulations of radiative transport for the study of optical probe design in heterogeneous tissues. SIAM Journal on Applied Mathematics, 68(1):253–270, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Luke:2007:ISO

[773] D. Russell Luke and Anthony J. Devaney. Identifying scattering obstacles by the construction of nonscattering waves. SIAM Journal on Applied Mathematics, 68(1):271–291, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ebert:2007:CSL

[774] Ute Ebert, Bernard Meulenbroek, and Lothar Schäfer. Convective stabilization of a Laplacian moving boundary problem with kinetic undercooling. SIAM Journal on Applied Mathematics, 68(1):292–310, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Waagaard:2007:ISM

[775] Ole Henrik Waagaard and Johannes Skaar. Inverse scattering in multimode structures. SIAM Journal on Applied Mathematics, 68(2):311–333, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Katsevich:2007:FBI

[776] Alexander Katsevich and Mikhail Kapralov. Filtered backprojection inversion of the cone beam transform for a general class of curves. SIAM Journal on Applied Mathematics, 68 (2):334–353, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nykamp:2007:EHD

[777] Duane Q. Nykamp. Exploiting historydependent effects to infer network connectivity. SIAM Journal on Applied Mathematics, 68(2):354–391, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Finch:2007:ISM

[778] David Finch, Markus Haltmeier, and Rakesh. Inversion of spherical means and the wave equation in even dimensions. SIAM Journal on Applied Mathematics, 68(2):392–412, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Moutari:2007:HLM

[779] S. Moutari and M. Rascle. A hybrid Lagrangian model based on the Aw–Rascle traffic flow model. SIAM Journal on Applied Mathematics, 68(2):413–436, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fornasier:2007:RCI

[780] Massimo Fornasier and Riccardo March.

Restoration of color images by vector valued BV functions and variational calculus. SIAM Journal on Applied Mathematics, 68(2):437–460, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zabarankin:2007:ACM

[781] M. Zabarankin. Asymmetric creeping motion of a rigid spindle-shaped body in a viscous fluid. SIAM Journal on Applied Mathematics, 68(2):461–485, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Consolini:2007:ESP

[782] Luca Consolini and Mario Tosques. On the existence of small periodic solutions for the 2-dimensional inverted pendulum on a cart. SIAM Journal on Applied Mathematics, 68(2):486–502, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wang:2007:DSB

[783] Hao Wang, Hal L. Smith, Yang Kuang, and James J. Elser. Dynamics of stoichiometric bacteria-algae interactions in the epilimnion. SIAM Journal on Applied Mathematics, 68(2):503–522, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Perrey-Debain:2007:DAA

[784] Emmanuel Perrey-Debain and I. David Abrahams. A diffusion analysis approach to TE mode propagation in randomly perturbed optical waveguides. SIAM Journal on Applied Mathematics, 68(2):523–543, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sheng:2007:GRP

[785] Wancheng Sheng, Meina Sun, and Tong Zhang. The generalized Riemann problem for a scalar nonconvex Chapman– Jouguet combustion model. SIAM Journal on Applied Mathematics, 68 (2):544–561, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Xu:2007:AWC

[786] Rui-Yue Xu, Peng Zhang, Shi-Qiang Dai, and S. C. Wong. Admissibility of a wide cluster solution in “anisotropic” higher-order traffic flow models. SIAM Journal on Applied Mathematics, 68 (2):562–573, ???? 2007. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

George:2007:FTR

[787] Ashvin George and Yoram Bresler.

Fast tomographic reconstruction via rotation-based hierarchical backprojection. SIAM Journal on Applied Mathematics, 68(2):574–597, ???? 2007. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Preusser:2008:PFM

[788] T. Preusser, M. Droske, C. S. Garbe, A. Telea, and M. Rumpf. A phase field method for joint denoising, edge detection, and motion estimation in image sequence processing. SIAM Journal on Applied Mathematics, 68 (3):599–618, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bonilla:2008:HBC

[789] Luis L. Bonilla, Manuel Carretero, and J. B. Keller. Homogeneous branchedchain explosions. SIAM Journal on Applied Mathematics, 68(3):619–628, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hazard:2008:STE

[790] Christophe Hazard and Michael H. Meylan. Spectral theory for an elastic thin plate floating on water of finite depth. SIAM Journal on Applied Mathematics, 68(3):629–647, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bonilla:2008:HLF

[791] L. L. Bonilla, T. Götz, A. Klar, N. Marheineke, and R. Wegener. Hydrodynamic limit of a Fokker–Planck equation describing fiber lay-down processes. SIAM Journal on Applied Mathematics, 68(3):648–665, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Engstrom:2008:IBT

[792] Christian Engström. Inverse bounds of two-component composites. SIAM Journal on Applied Mathematics, 68 (3):666–679, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ai:2008:TWB

[793] Shangbing Ai. Traveling waves in a bioremediation model. SIAM Journal on Applied Mathematics, 68(3):680–693, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shen:2008:CSF

[794] Jackie (Jianhong) Shen. Cucker–Smale flocking under hierarchical leadership.

SIAM Journal on Applied Mathematics, 68(3):694–719, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mitchell:2008:NTH

[795] Colleen Mitchell and Michael Reed.

Neural timing in highly convergent systems. SIAM Journal on Applied Mathematics, 68(3):720–737, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sipahi:2008:STF

[796] Rifat Sipahi, Fatihcan M. Atay, and Silviu-Iulian Niculescu. Stability of traffic flow behavior with distributed delays modeling the memory effects of the drivers. SIAM Journal on Applied Mathematics, 68(3):738–759, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cook:2008:SSP

[797] Benjamin P. Cook, Andrea L. Bertozzi, and A. E. Hosoi. Shock solutions for particle-laden thin films. SIAM Journal on Applied Mathematics, 68 (3):760–783, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kusiak:2008:ICM

[798] Steven Kusiak and John Weatherwax.

Identification and characterization of a mobile source in a general parabolic differential equation with constant coefficients. SIAM Journal on Applied Mathematics, 68(3):784–805, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zhang:2008:DSF

[799] Sirong Zhang, Laurent Younes, John Zweck, and J. Tilak Ratnanather. Diffeomorphic surface flows: a novel method of surface evolution. SIAM Journal on Applied Mathematics, 68 (3):806–824, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Maslowe:2008:NCL

[800] S. A. Maslowe and N. Nigam. The nonlinear critical layer for Kelvin modes on a vortex with a continuous velocity profile. SIAM Journal on Applied Mathematics, 68(3):825–843, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Singer:2008:PRD

[801] A. Singer, Z. Schuss, A. Osipov, and D. Holcman. Partially reflected diffusion. SIAM Journal on Applied Mathematics, 68(3):844–868, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gabriel:2008:MMS

[802] J.-P. Gabriel, L. M. Studer, D. G. Rüegg, and M.-A. Schnetzer. A mathematical model for the steady activation of a skeletal muscle. SIAM Journal on Applied Mathematics, 68 (3):869–889, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wendl:2008:RCM

[803] Michael C. Wendl. Random covering of multiple one-dimensional domains with an application to DNA sequencing. SIAM Journal on Applied Mathematics, 68(3):890–905, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hanke:2008:FME

[804] Martin Hanke and Birgit Schappel.

The factorization method for electrical impedance tomography in the halfspace. SIAM Journal on Applied Mathematics, 68(4):907–924, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Aubert:2008:VAR

[805] Gilles Aubert and Jean-François Aujol.

A variational approach to removing multiplicative noise. SIAM Journal on Applied Mathematics, 68(4):925–946, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Adams:2008:HER

[806] Fred C. Adams and Anthony M. Bloch.

Hill’s equation with random forcing terms. SIAM Journal on Applied Mathematics, 68(4):947–980, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

DApice:2008:FDM

[807] Ciro D’Apice, Rosanna Manzo, and Benedetto Piccoli. A fluid dynamic model for telecommunication networks with sources and destinations. SIAM Journal on Applied Mathematics, 68 (4):981–1003, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bueno:2008:SSM

[808] H. Bueno, G. Ercole, and A. Zumpano.

Stationary solutions of a model for the growth of tumors and a connection between the nonnecrotic and necrotic phases. SIAM Journal on Applied Mathematics, 68(4):1004–1025, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Uhlmann:2008:RDU

[809] Gunther Uhlmann and Jenn-Nan Wang. Reconstructing discontinuities using complex geometrical optics solutions. SIAM Journal on Applied Mathematics, 68(4):1026–1044, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Touboul:2008:BAG

[810] Jonathan Touboul. Bifurcation analysis of a general class of nonlinear integrate-and-fire neurons. SIAM Journal on Applied Mathematics, 68(4): 1045–1079, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dassios:2008:MSE

[811] G. Dassios and A. S. Fokas. Methods for solving elliptic PDEs in spherical coordinates. SIAM Journal on Applied Mathematics, 68(4):1080–1096, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lechleiter:2008:FMA

[812] Armin Lechleiter, Nuutti Hyvönen, and Harri Hakula. The factorization method applied to the complete electrode model of impedance tomography. SIAM Journal on Applied Mathematics, 68(4): 1097–1121, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Arrayas:2008:FIF

[813] Manuel Arrayás, Santiago Betelú, Marco A. Fontelos, and José L. Trueba. Fingering from ionization fronts in plasmas. SIAM Journal on Applied Mathematics, 68(4):1122–1145, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lagache:2008:EMV

[814] Thibault Lagache and David Holcman.

Effective motion of a virus trafficking inside a biological cell. SIAM Journal on Applied Mathematics, 68 (4):1146–1167, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Subramanian:2008:OLL

[815] Ajay Subramanian. Optimal liquidation by a large investor. SIAM Journal on Applied Mathematics, 68 (4):1168–1201, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rosencrans:2008:CSD

[816] Steve Rosencrans and Xuefeng Wang.

Corrigendum: Suppression of the Dirichlet eigenvalues of a coated body. SIAM Journal on Applied Mathematics, 68(4):1202, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [661].

Goldsztein:2008:STP

[817] Guillermo H. Goldsztein. Solute transport in porous media. media with capillaries as voids. SIAM Journal on Applied Mathematics, 68(5):1203–1222, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bressloff:2008:DPR

[818] Paul C. Bressloff, Berton A. Earnshaw, and Michael J. Ward. Diffusion of protein receptors on a cylindrical dendritic membrane with partially absorbing traps. SIAM Journal on Applied Mathematics, 68(5): 1223–1246, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

La:2008:CDA

[819] Richard J. La and Priya Ranjan. Convergence of dual algorithm with arbitrary communication delays. SIAM Journal on Applied Mathematics, 68 (5):1247–1267, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sviercoski:2008:AAG

[820] R. F. Sviercoski, C. L. Winter, and A. W. Warrick. Analytical approximation for the generalized Laplace equation with step function coefficient. SIAM Journal on Applied Mathematics, 68 (5):1268–1281, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Quinto:2008:LTE

[821] Eric Todd Quinto and Ozan Öktem. Local tomography in electron microscopy. SIAM Journal on Applied Mathematics, 68(5):1282–1303, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Liu:2008:BCM

[822] Chun Liu and Hailiang Liu. Boundary conditions for the microscopic FENE models. SIAM Journal on Applied Mathematics, 68(5):1304–1315, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Camacho:2008:TWS

[823] Victor Camacho, Robert D. Guy, and Jon Jacobsen. Traveling waves and shocks in a viscoelastic generalization of Burgers’ equation. SIAM Journal on Applied Mathematics, 68(5): 1316–1332, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Perrier:2008:CJC

[824] Vincent Perrier. The Chapman– Jouguet closure for the Riemann problem with vaporization. SIAM Journal on Applied Mathematics, 68(5): 1333–1359, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tesdall:2008:SSS

[825] Allen M. Tesdall, Richard Sanders, and Barbara L. Keyfitz. Self-similar solutions for the triple point paradox in gasdynamics. SIAM Journal on Applied Mathematics, 68(5):1360–1377, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Griesmaier:2008:AFM

[826] Roland Griesmaier. An asymptotic factorization method for inverse electromagnetic scattering in layered media. SIAM Journal on Applied Mathematics, 68(5):1378–1403, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Olagunju:2008:ANS

[827] David O. Olagunju, Anand B. Vyas, and Shangyou Zhang. Analytical and numerical solutions for torsional flow between coaxial discs with heat transfer. SIAM Journal on Applied Mathematics, 68(5):1404–1422, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cuesta:2008:SSW

[828] C. M. Cuesta and C. Schmeiser. Stability of solitary waves in a semiconductor drift-diffusion model. SIAM Journal on Applied Mathematics, 68 (5):1423–1438, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Abrahams:2008:ACD

[829] I. David Abrahams, Anthony M. J. Davis, and Stefan G. Llewellyn Smith. Asymmetric channel divider in Stokes flow. SIAM Journal on Applied Mathematics, 68(5):1439–1463, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Anderson:2008:GAS

[830] David F. Anderson. Global asymptotic stability for a class of nonlinear chemical equations. SIAM Journal on Applied Mathematics, 68(5): 1464–1476, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rodriguez:2008:VCE

[831] Ana Alonso Rodríguez and Alberto Valli. Voltage and current excitation for time-harmonic eddy-current problems.

SIAM Journal on Applied Mathematics, 68(5):1477–1494, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wang:2008:GSV

[832] Lin Wang and Guan zhong Dai. Global stability of virus spreading in complex heterogeneous networks. SIAM Journal on Applied Mathematics, 68 (5):1495–1502, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kriegsmann:2008:EDS

[833] G. A. Kriegsmann. Electric discharge sintering: a mathematical model. SIAM Journal on Applied Mathematics, 68 (6):1503–1517, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Caginalp:2008:NTP

[834] Gunduz Caginalp, Xinfu Chen, and Christof Eck. Numerical tests of a phase field model with second order accuracy. SIAM Journal on Applied Mathematics, 68(6):1518–1534, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Phung:2008:TRF

[835] Kim Dang Phung and Xu Zhang. Time reversal focusing of the initial state for Kirchhoff plate. SIAM Journal on Applied Mathematics, 68(6):1535–1556, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ammari:2008:EIT

[836] H. Ammari, E. Bonnetier, Y. Capdeboscq, M. Tanter, and M. Fink. Electrical impedance tomography by elastic deformation. SIAM Journal on Applied Mathematics, 68(6):1557–1573, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Garnier:2008:ETE

[837] Josselin Garnier and Knut Sølna. Effective transport equations and enhanced backscattering in random waveguides. SIAM Journal on Applied Mathematics, 68(6):1574–1599, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hsu:2008:CEM

[838] Sze-Bi Hsu and Ting-Hao Hsu. Competitive exclusion of microbial species for a single nutrient with internal storage. SIAM Journal on Applied Mathematics, 68(6):1600–1617, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gosse:2008:EUC

[839] Laurent Gosse and Olof Runborg. Existence, uniqueness, and a constructive solution algorithm for a class of finite Markov moment problems. SIAM Journal on Applied Mathematics, 68 (6):1618–1640, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zhang:2008:IDS

[840] Hang Zhang and M. Carme Calderer. Incipient dynamics of swelling of gels. SIAM Journal on Applied Mathematics, 68(6):1641–1664, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gourley:2008:MMC

[841] Stephen A. Gourley and Xingfu Zou.

A mathematical model for the control and eradication of a wood boring beetle infestation. SIAM Journal on Applied Mathematics, 68(6): 1665–1687, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zadorozhnii:2008:FER

[842] V. I. Zadorozhnii, T. J. Sluckin, V. Yu.

Reshetnyak, and K. S. Thomas. The Frederiks Effect and related phenomena in ferronematic materials. SIAM Journal on Applied Mathematics, 68 (6):1688–1716, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Smith:2008:MVP

[843] Hal L. Smith. Models of virulent phage growth with application to phage therapy. SIAM Journal on Applied Mathematics, 68(6):1717–1737, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bal:2008:TBI

[844] Guillaume Bal and Kui Ren. Transportbased imaging in random media. SIAM Journal on Applied Mathematics, 68 (6):1738–1762, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Casenave:2008:TLD

[845] C. Casenave and E. Montseny. Timelocal dissipative formulation and stable numerical schemes for a class of integrodifferential wave equations. SIAM Journal on Applied Mathematics, 68 (6):1763–1782, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Thompson:2008:ITS

[846] I. Thompson and C. M. Linton. An interaction theory for scattering by defects in arrays. SIAM Journal on Applied Mathematics, 68(6):1783–1806, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cox:2008:MSA

[847] Benjamin P. Cox and Warren R. Smith. On a multiple-scales analysis of multilateral phenomena in semiconductor lasers. SIAM Journal on Applied Mathematics, 69(1):1–21, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kim:2008:MTA

[848] Sungwhan Kim, Jeehyun Lee, Jin Keun Seo, Eung Je Woo, and Habib Zribi. Multifrequency trans-admittance scanner: Mathematical framework and feasibility. SIAM Journal on Applied Mathematics, 69(1):22–36, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gopalakrishnan:2008:ANT

[849] J. Gopalakrishnan, S. Moskow, and F. Santosa. Asymptotic and numerical techniques for resonances of thin photonic structures. SIAM Journal on Applied Mathematics, 69(1):37–63, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stocker:2008:LSA

[850] Christina Stöcker and Axel Voigt. A level set approach to anisotropic surface evolution with free adatoms. SIAM Journal on Applied Mathematics, 69 (1):64–80, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Marengo:2008:ISP

[851] Edwin A. Marengo, Mohamed R.

Khodja, and Abdelkader Boucherif. Inverse source problem in nonhomogeneous background media. Part II: Vector formulation and antenna substrate performance characterization. SIAM Journal on Applied Mathematics, 69 (1):81–110, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Susanto:2008:BDW

[852] Hadi Susanto. Boundary driven waveguide arrays: Supratransmission and saddle-node bifurcation. SIAM Journal on Applied Mathematics, 69(1):111–125, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fontelos:2008:SCD

[853] M. A. Fontelos and U. Kindelán. The shape of charged drops over a solid surface and symmetry-breaking instabilities. SIAM Journal on Applied Mathematics, 69(1):126–148, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Clarelli:2008:MMC

[854] Fabrizio Clarelli, Antonio Fasano, and Roberto Natalini. Mathematics and monument conservation: Free boundary models of marble sulfation. SIAM Journal on Applied Mathematics, 69 (1):149–168, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Knoblauch:2008:NAM

[855] Andreas Knoblauch. Neural associative memory and the Willshaw–Palm probability distribution. SIAM Journal on Applied Mathematics, 69(1):169–196, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Knoblauch:2008:CFE

[856] Andreas Knoblauch. Closed-form expressions for the moments of the binomial probability distribution. SIAM Journal on Applied Mathematics, 69 (1):197–204, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Faugeras:2008:ASC

[857] Olivier Faugeras, François Grimbert, and Jean-Jacques Slotine. Absolute stability and complete synchronization in a class of neural fields models. SIAM Journal on Applied Mathematics, 69 (1):205–250, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Golovin:2008:TPF

[858] A. A. Golovin, B. J. Matkowsky, and V. A. Volpert. Turing pattern formation in the Brusselator model with superdiffusion. SIAM Journal on Applied Mathematics, 69(1):251–272, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chen:2008:PLD

[859] Xinfu Chen and Yuanwei Qi. Propagation of local disturbances in reaction diffusion systems modeling quadratic autocatalysis. SIAM Journal on Applied Mathematics, 69(1):273–282, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wu:2008:TEM

[860] Jinbiao Wu, C. Sean Bohun, and Huaxiong Huang. A thermal elastic model for directional crystal growth with weak anisotropy. SIAM Journal on Applied Mathematics, 69(1):283–304, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ellis:2008:CMC

[861] A. S. Ellis and F. T. Smith. A continuum model for a chute flow of grains. SIAM Journal on Applied Mathematics, 69(2):305–329, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Huang:2008:OFD

[862] H. Huang, R. M. Miura, and J. J. Wylie.

Optical fiber drawing and dopant transport. SIAM Journal on Applied Mathematics, 69(2):330–347, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Korzec:2008:SSD

[863] M. D. Korzec, P. L. Evans, A. Münch, and B. Wagner. Stationary solutions of driven fourth- and sixth-order Cahn–Hilliard-type equations. SIAM Journal on Applied Mathematics, 69 (2):348–374, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wang:2008:MMM

[864] Zhi-An Wang, Thomas Hillen, and Michael Li. Mesenchymal motion models in one dimension. SIAM Journal on Applied Mathematics, 69(2):375–397, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Budaev:2008:DNP

[865] Bair V. Budaev and David B. Bogy. Diffraction by a nonconvex polygon. SIAM Journal on Applied Mathematics, 69(2):398–418, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wei:2008:MES

[866] Juncheng Wei and Matthias Winter. Mutually exclusive spiky pattern and segmentation modeled by the fivecomponent Meinhardt–Gierer system. SIAM Journal on Applied Mathematics, 69(2):419–452, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yariv:2008:TDS

[867] Ehud Yariv. Thermophoresis due to strong temperature gradients. SIAM Journal on Applied Mathematics, 69 (2):453–472, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Glasner:2008:ORT

[868] K. B. Glasner. Ostwald ripening in thin film equations. SIAM Journal on Applied Mathematics, 69(2):473–493, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bonilla:2008:NES

[869] L. L. Bonilla, L. Barletti, and M. Alvaro.

Nonlinear electron and spin transport in semiconductor superlattices. SIAM Journal on Applied Mathematics, 69 (2):494–513, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Olmstead:2008:TBS

[870] W. E. Olmstead and Catherine A.

Roberts. Thermal blow-up in a subdiffusive medium. SIAM Journal on Applied Mathematics, 69(2):514–523, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rosso:2008:SRC

[871] Riccardo Rosso and Marco Verani. Stabilizing role of a curvature correction to line tension. SIAM Journal on Applied Mathematics, 69(2):524–551, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stone:2008:SPM

[872] Brenton R. Stone, Gino Biondini, and William L. Kath. Statistics of polarization-mode dispersion emulators with unequal sections. SIAM Journal on Applied Mathematics, 69(2):552–564, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gebauer:2008:IAT

[873] Bastian Gebauer and Otmar Scherzer.

Impedance-acoustic tomography. SIAM Journal on Applied Mathematics, 69 (2):565–576, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kang:2008:IPS

[874] Hyeonbae Kang, Eunjoo Kim, and Graeme W. Milton. Inclusion pairs satisfying Eshelby’s uniformity property. SIAM Journal on Applied Mathematics, 69(2):577–595, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Du:2008:QAP

[875] Yihong Du, Peter Y. H. Pang, and Mingxin Wang. Qualitative analysis of a prey-predator model with stage structure for the predator. SIAM Journal on Applied Mathematics, 69 (2):596–620, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tang:2008:CLC

[876] Yilei Tang, Deqing Huang, Shigui Ruan, and Weinian Zhang. Coexistence of limit cycles and homoclinic loops in a SIRS model with a nonlinear incidence rate. SIAM Journal on Applied Mathematics, 69(2):621–639, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zhang:2008:PFM

[877] Tianyu Zhang, N. G. Cogan, and Qi Wang. Phase field models for biofilms. I. Theory and one-dimensional simulations. SIAM Journal on Applied Mathematics, 69(3):641–669, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Velo:2008:RDR

[878] Ani P. Velo, George A. Gazonas, Erwin Bruder, and Nancy Rodriguez.

Recursive dispersion relations in onedimensional periodic elastic media. SIAM Journal on Applied Mathematics, 69(3):670–689, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Horne:2008:SFW

[879] Rudy L. Horne, Christopher K. R. T.

Jones, and Tobias Schäfer. The suppression of four-wave mixing by random dispersion. SIAM Journal on Applied Mathematics, 69(3):690–703, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dai:2008:SLA

[880] Shu Dai and David G. Schaeffer. Spectrum of a linearized amplitude equation for alternans in a cardiac fiber. SIAM Journal on Applied Mathematics, 69(3):704–719, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Glimm:2008:TSF

[881] James Glimm, Xiaomei Ji, Jiequan Li, Xiaolin Li, Peng Zhang, Tong Zhang, and Yuxi Zheng. Transonic shock formation in a rarefaction Riemann problem for the 2D compressible Euler equations. SIAM Journal on Applied Mathematics, 69(3):720–742, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kampel:2008:FNC

[882] Guido Kampel and Guillermo H. Goldsztein. Filters. the number of channels that can clog in a network. SIAM Journal on Applied Mathematics, 69 (3):743–762, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Renardy:2008:SWS

[883] Michael Renardy. Short wave stability for inviscid shear flow. SIAM Journal on Applied Mathematics, 69 (3):763–768, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schaeffer:2008:SIS

[884] David G. Schaeffer and Richard M. Iverson. Steady and intermittent slipping in a model of landslide motion regulated by pore-pressure feedback. SIAM Journal on Applied Mathematics, 69 (3):769–786, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Calderer:2008:CTC

[885] Maria-Carme Calderer and Sookyung Joo. A continuum theory of chiral smectic C liquid crystals. SIAM Journal on Applied Mathematics, 69 (3):787–809, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Eloe:2008:OSR

[886] P. Eloe, R. H. Liu, M. Yatsuki, G. Yin, and Q. Zhang. Optimal selling rules in a regime-switching exponential Gaussian diffusion model. SIAM Journal on Applied Mathematics, 69(3):810–829, ???? 2008. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Antoine:2008:FFM

[887] X. Antoine, B. Pinçon, K. Ramdani, and B. Thierry. Far field modeling of electromagnetic time reversal and application to selective focusing on small scatterers. SIAM Journal on Applied Mathematics, 69(3):830–844, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zabarankin:2008:FHAa

[888] Michael Zabarankin. The framework of k-harmonically analytic functions for three-dimensional Stokes flow problems, Part I. SIAM Journal on Applied Mathematics, 69(3):845–880, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zabarankin:2008:FHAb

[889] Michael Zabarankin. The framework of k-harmonically analytic functions for three-dimensional Stokes flow problems, Part II. SIAM Journal on Applied Mathematics, 69(3):881–907, ???? 2008. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

DeHoop:2009:EGF

[890] Maarten V. De Hoop and Knut Sølna.

Estimating a Green’s function from “field-field” correlations in a random medium. SIAM Journal on Applied Mathematics, 69(4):909–932, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gonzalez:2009:SCS

[891] O. Gonzalez. On stable, complete, and singularity-free boundary integral formulations of exterior Stokes flow. SIAM Journal on Applied Mathematics, 69(4):933–958, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kon:2009:MHI

[892] Ryusuke Kon and Sebastian J. Schreiber. Multiparasitoid-host interactions with egg-limited encounter rates. SIAM Journal on Applied Mathematics, 69 (4):959–976, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shinar:2009:SRC

[893] Guy Shinar, Uri Alon, and Martin Feinberg. Sensitivity and robustness in chemical reaction networks. SIAM Journal on Applied Mathematics, 69 (4):977–998, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Reluga:2009:AHC

[894] Timothy C. Reluga, Harel Dahari, and Alan S. Perelson. Analysis of hepatitis C virus infection models with hepatocyte homeostasis. SIAM Journal on Applied Mathematics, 69(4):999–1023, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zeev:2009:ITD

[895] Noam Zeev and Fioralba Cakoni. The identification of thin dielectric objects from far field or near field scattering data. SIAM Journal on Applied Mathematics, 69(4):1024–1042, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Thiyanaratnam:2009:MSD

[896] Pradeep Thiyanaratnam, Russel Caflisch, Paulo S. Motta, and Jack W. Judy. Modeling, simulation, and design for a customizable electrodeposition process.

SIAM Journal on Applied Mathematics, 69(4):1043–1064, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Marty:2009:AWL

[897] Renaud Marty and Knut Sølna. Acoustic waves in long range random media. SIAM Journal on Applied Mathematics, 69(4):1065–1083, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Needham:2009:UFW

[898] D. J. Needham, S. Langdon, G. S. Busswell, and J. P. Gilchrist. The unsteady flow of a weakly compressible fluid in a thin porous layer I: Two-dimensional theory. SIAM Journal on Applied Mathematics, 69(4):1084–1109, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shih:2009:GDC

[899] Mau-Hsiang Shih and Feng-Sheng Tsai.

Growth dynamics of cell assemblies. SIAM Journal on Applied Mathematics, 69(4):1110–1161, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Abdallah:2009:SCS

[900] Naoufel Ben Abdallah, François Castella, Fanny Delebecque-Fendt, and Florian Méhats. The strongly confined Schrödinger–Poisson system for the transport of electrons in a nanowire. SIAM Journal on Applied Mathematics, 69(4):1162–1173, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zagaris:2009:BNC

[901] A. Zagaris, A. Doelman, N. N. Pham Thi, and B. P. Sommeijer. Blooming in a nonlocal, coupled phytoplanktonnutrient model. SIAM Journal on Applied Mathematics, 69(4):1174–1204, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jiang:2009:CCD

[902] Jifa Jiang and Zhipeng Qiu. The complete classification for dynamics in a nine-dimensional West Nile virus model. SIAM Journal on Applied Mathematics, 69(5):1205–1227, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Iron:2009:SCI

[903] David Iron, Theodore Kolokolonikov, John Rumsey, and Juncheng Wei. Stability of curved interfaces in the perturbed two-dimensional Allen–Cahn system. SIAM Journal on Applied Mathematics, 69(5):1228–1243, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Aguirre:2009:TLC

[904] Pablo Aguirre, Eduardo González-Olivares, and Eduardo Sáez. Three limit cycles in a Leslie–Gower predatorprey model with additive Allee effect. SIAM Journal on Applied Mathematics, 69(5):1244–1262, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Strong:2009:DDC

[905] Natalia Strong. Double-diffusive convection in a porous layer in the presence of vibration. SIAM Journal on Applied Mathematics, 69(5):1263–1276, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fricks:2009:TDM

[906] John Fricks, Lingxing Yao, Timothy C.

Elston, and M. Gregory Forest. Timedomain methods for diffusive transport in soft matter. SIAM Journal on Applied Mathematics, 69(5): 1277–1308, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kubzina:2009:DSI

[907] V. Kubzina, A. K. Gautesen, and L. Ju. Fradkin. Diffraction by a semiinfinite interfacial crack sandwiched between two isotropic half planes. SIAM Journal on Applied Mathematics, 69 (5):1309–1333, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Galaktionov:2009:CSB

[908] V. A. Galaktionov and P. J. Harwin. On center subspace behavior in thin film equations. SIAM Journal on Applied Mathematics, 69(5): 1334–1358, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hitczenko:2009:BOI

[909] Pawel Hitczenko and Georgi S. Medvedev.

Bursting oscillations induced by small noise. SIAM Journal on Applied Mathematics, 69(5):1359–1392, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Xu:2009:DCM

[910] Haoyun Xu and Yang Xiang. Derivation of a continuum model for the longrange elastic interaction on stepped epitaxial surfaces in 2 + 1 dimensions. SIAM Journal on Applied Mathematics, 69(5):1393–1414, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cox:2009:VCS

[911] Stephen M. Cox and Stephen C. Creagh. Voltage and current spectra for matrix power converters. SIAM Journal on Applied Mathematics, 69(5): 1415–1437, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Volkov:2009:DLS

[912] Darko Volkov. A double layer surface traction free Green’s tensor. SIAM Journal on Applied Mathematics, 69 (5):1438–1456, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Knessl:2009:TDD

[913] Charles Knessl and John A. Morrison. A two-dimensional diffusion approximation for a loss model with trunk reservation. SIAM Journal on Applied Mathematics, 69(5):1457–1476, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Monnier:2009:NME

[914] Jérôme Monnier, Patrick Witomski, Patrick Chow-Wing-Bom, and Claire Scheid. Numerical modeling of electrowetting by a shape inverse approach.

SIAM Journal on Applied Mathematics, 69(5):1477–1500, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Thanh:2009:RPN

[915] Mai Duc Thanh. The Riemann problem for a nonisentropic fluid in a nozzle with discontinuous cross-sectional area. SIAM Journal on Applied Mathematics, 69(6):1501–1519, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gayen:2009:SSW

[916] Rupanwita Gayen and B. N. Mandal.

Scattering of surface water waves by a floating elastic plate in two dimensions. SIAM Journal on Applied Mathematics, 69(6):1520–1541, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sun:2009:IEW

[917] Meina Sun. Interactions of elementary waves for the Aw–Rascle model. SIAM Journal on Applied Mathematics, 69(6):1542–1558, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lynch:2009:DDP

[918] Sean Lynch and Charles Knessl. Driftdiffusion past a circle: Shadow region asymptotics. SIAM Journal on Applied Mathematics, 69(6):1559–1579, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Xia:2009:NCE

[919] Yonghui Xia and Maoan Han. New conditions on the existence and stability of periodic solution in Lotka– Volterra’s population system. SIAM Journal on Applied Mathematics, 69 (6):1580–1597, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hoang:2009:BES

[920] Viet Ha Hoang. Bounds for the effective stress of classical and strain gradient plastic composites. SIAM Journal on Applied Mathematics, 69 (6):1598–1617, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Katsaounis:2009:EEL

[921] Theodoros Katsaounis and Athanasios E. Tzavaras. Effective equations for localization and shear band formation. SIAM Journal on Applied Mathematics, 69(6):1618–1643, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ackleh:2009:NJA

[922] Azmy S. Ackleh and Keng Deng. A nonautonomous juvenile-adult model: Well-posedness and long-time behavior via a comparison principle. SIAM Journal on Applied Mathematics, 69 (6):1644–1661, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Harrach:2009:DIE

[923] Bastian Harrach and Jin Keun Seo. Detecting inclusions in electrical impedance tomography without reference measurements. SIAM Journal on Applied Mathematics, 69(6):1662–1681, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Carmona:2009:MCM

[924] René Carmona and Lixin Wang. Monte Carlo Malliavin computation of the sensitivities of solutions of SPDEs. SIAM Journal on Applied Mathematics, 69 (6):1682–1711, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Choksi:2009:PDM

[925] Rustum Choksi, Mark A. Peletier, and J. F. Williams. On the phase diagram for microphase separation of diblock copolymers: An approach via a nonlocal Cahn–Hilliard functional. SIAM Journal on Applied Mathematics, 69 (6):1712–1738, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mulone:2009:NSD

[926] Giuseppe Mulone and Brian Straughan.

Nonlinear stability for diffusion models in biology. SIAM Journal on Applied Mathematics, 69(6):1739–1758, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kim:2009:NEM

[927] Sungwhan Kim, Jin Keun Seo, and Taeyoung Ha. A nondestructive evaluation method for concrete voids: Frequency differential electrical impedance scanning. SIAM Journal on Applied Mathematics, 69(6):1759–1771, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Clayton:2009:MSB

[928] Aubrey Clayton and Steven N. Evans.

Mutation-selection balance with recombination: Convergence to equilibrium for polynomial selection costs. SIAM Journal on Applied Mathematics, 69 (6):1772–1792, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Atkinson:2009:FBP

[929] C. Atkinson. A free boundary problem of a real option model. SIAM Journal on Applied Mathematics, 69 (6):1793–1804, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

vanLeeuwaarden:2009:AIE

[930] J. S. H. van Leeuwaarden and N. M. Temme. Asymptotic inversion of the Erlang B formula. SIAM Journal on Applied Mathematics, 70(1):1–23, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fok:2009:IGD

[931] Pak-Wing Fok and Tom Chou. Interface growth driven by surface kinetics and convection. SIAM Journal on Applied Mathematics, 70(1):24–39, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hinow:2009:AMT

[932] Peter Hinow, Frank Le Foll, Pierre Magal, and Glenn F. Webb. Analysis of a model for transfer phenomena in biological populations. SIAM Journal on Applied Mathematics, 70(1):40–62, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gibelli:2009:CTA

[933] Livio Gibelli and Stefano Turzi. A catastrophe-theoretic approach to tricritical points with application to liquid crystals. SIAM Journal on Applied Mathematics, 70(1):63–76, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Isaacson:2009:RDM

[934] Samuel A. Isaacson. The reactiondiffusion master equation as an asymptotic approximation of diffusion to a small target. SIAM Journal on Applied Mathematics, 70(1):77–111, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mommer:2009:MSU

[935] Mario S. Mommer and Dirk Lebiedz.

Modeling subdiffusion using reaction diffusion systems. SIAM Journal on Applied Mathematics, 70(1):112–132, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Xue:2009:MMT

[936] Chuan Xue and Hans G. Othmer. Multiscale models of taxis-driven patterning in bacterial populations. SIAM Journal on Applied Mathematics, 70 (1):133–167, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Elschner:2009:OMI

[937] Johannes Elschner, George C. Hsiao, and Andreas Rathsfeld. An optimization method in inverse acoustic scattering by an elastic obstacle. SIAM Journal on Applied Mathematics, 70 (1):168–187, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Thieme:2009:SBR

[938] Horst R. Thieme. Spectral bound and reproduction number for infinitedimensional population structure and time heterogeneity. SIAM Journal on Applied Mathematics, 70(1):188–211, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Moulton:2009:CEF

[939] D. E. Moulton and J. A. Pelesko. Catenoid in an electric field. SIAM Journal on Applied Mathematics, 70 (1):212–230, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Engstrom:2009:SOP

[940] Christian Engström and Markus Richter. On the spectrum of an operator pencil with applications to wave propagation in periodic and frequency dependent materials. SIAM Journal on Applied Mathematics, 70(1):231–247, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Straube:2009:AAI

[941] Ronny Straube and Michael J. Ward. An asymptotic analysis of intracellular signaling gradients arising from multiple small compartments. SIAM Journal on Applied Mathematics, 70 (1):248–269, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

deHoop:2009:ASC

[942] Adrianus T. de Hoop. Analytic solutions for a class of point-source excited transient wave motions in anisotropic, lossy media. SIAM Journal on Applied Mathematics, 70(1):270–278, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

delaHoz:2009:NSF

[943] Francisco de la Hoz. Numerical study of a flow of regular planar curves that develop singularities at finite time. SIAM Journal on Applied Mathematics, 70 (1):279–301, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Coombs:2009:DSL

[944] Daniel Coombs, Ronny Straube, and Michael Ward. Diffusion on a sphere with localized traps: Mean first passage time, eigenvalue asymptotics, and Fekete points. SIAM Journal on Applied Mathematics, 70(1):302–332, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Belhachmi:2009:HCI

[945] Zakaria Belhachmi, Dorin Bucur, Bernhard Burgeth, and Joachim Weickert. How to choose interpolation data in images. SIAM Journal on Applied Mathematics, 70(1):333–352, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zhou:2009:SDM

[946] Lin Zhou and Gregory A. Kriegsmann.

A simple derivation of microstrip transmission line equations. SIAM Journal on Applied Mathematics, 70(1):353–367, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Peng:2009:ESP

[947] Rui Peng, Junping Shi, and Mingxin Wang. Erratum: Stationary pattern of a ratio-dependent food chain model with diffusion. SIAM Journal on Applied Mathematics, 70(1):368, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [745].

Promislow:2009:PFC

[948] Keith Promislow and Brian Wetton. PEM fuel cells: a mathematical overview. SIAM Journal on Applied Mathematics, 70(2):369–409, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gurau:2009:COC

[949] Vladimir Gurau and J. Adin Mann, Jr. A critical overview of computational fluid dynamics multiphase models for proton exchange membrane fuel cells. SIAM Journal on Applied Mathematics, 70(2):410–454, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Vynnycky:2009:ARN

[950] M. Vynnycky, G. Shugai, P. Yakubenko, and N. Mellgren. Asymptotic reduction for numerical modeling of polymer electrolyte fuel cells. SIAM Journal on Applied Mathematics, 70(2):455–487, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Weber:2009:UQP

[951] Adam Z. Weber and John Newman. Using a quasi-potential transformation for modeling diffusion media in polymerelectrolyte fuel cells. SIAM Journal on Applied Mathematics, 70(2):488–509, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fehribach:2009:TPB

[952] Joseph D. Fehribach and Ryan O’Hayre.

Triple phase boundaries in solid-oxide cathodes. SIAM Journal on Applied Mathematics, 70(2):510–530, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kulikovsky:2009:EPA

[953] A. A. Kulikovsky. Efficient parallel algorithm for fuel cell stack simulation. SIAM Journal on Applied Mathematics, 70(2):531–542, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fehribach:2009:VSM

[954] Joseph D. Fehribach. Vector-space methods and Kirchhoff graphs for reaction networks. SIAM Journal on Applied Mathematics, 70(2):543–562, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Potthast:2009:UCR

[955] Roland Potthast and Martin Wannert.

Uniqueness of current reconstructions for magnetic tomography in multilayer devices. SIAM Journal on Applied Mathematics, 70(2):563–578, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bozzini:2009:MMC

[956] Benedetto Bozzini and Ivonne Sgura. A mathematical model for the corrosion of metallic bipolar plates in PEM fuel cells: Numerical and experimental issues. SIAM Journal on Applied Mathematics, 70(2):579–599, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sun:2009:NNT

[957] Pengtao Sun, Guangri Xue, Chaoyang Wang, and Jinchao Xu. New numerical techniques for a three-dimensional liquid-feed direct methanol fuel cell model. SIAM Journal on Applied Mathematics, 70(2):600–620, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chudej:2009:OCL

[958] Kurt Chudej, Hans Josef Pesch, and Kati Sternberg. Optimal control of load changes for molten carbonate fuel cell systems: a challenge in PDE constrained optimization. SIAM Journal on Applied Mathematics, 70(2):621–639, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wu:2009:SLV

[959] Fuke Wu and Yong Xu. Stochastic Lotka–Volterra population dynamics with infinite delay. SIAM Journal on Applied Mathematics, 70(3):641–657, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pierre:2009:NAN

[960] C. Pierre and F. Plouraboué. Numerical analysis of a new mixed formulation for eigenvalue convection-diffusion problems. SIAM Journal on Applied Mathematics, 70(3):658–676, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dhia:2009:DDO

[961] Anne-Sophie Bonnet-Ben Dhia, Ghania Dakhia, Christophe Hazard, and Lahcène Chorfi. Diffraction by a defect in an open waveguide: a mathematical analysis based on a modal radiation condition. SIAM Journal on Applied Mathematics, 70(3):677–693, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Azizoglu:2009:TDS

[962] S. Alp Azizoglu, S. Sencer Koc, and O. Merih Buyukdura. Time domain scattering of scalar waves by two spheres in free-space. SIAM Journal on Applied Mathematics, 70(3):694–709, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rosa:2009:EME

[963] Salvatore La Rosa, Giovanni Mascali, and Vittorio Romano. Exact maximum entropy closure of the hydrodynamical model for si semiconductors: The 8moment case. SIAM Journal on Applied Mathematics, 70(3):710–734, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Junca:2009:DOD

[964] Stéphane Junca and Bruno Lombard.

Dilatation of a one-dimensional nonlinear crack impacted by a periodic elastic wave. SIAM Journal on Applied Mathematics, 70(3):735–761, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Velo:2009:TMO

[965] Ani P. Velo, George A. Gazonas, and Takanobu Ameya. z-transform methods for the optimal design of onedimensional layered elastic media. SIAM Journal on Applied Mathematics, 70 (3):762–788, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See erratum [995].

Aronson:2009:PFM

[966] D. G. Aronson and Víctor Padrón. Pattern formation in a model of an injured nerve fiber. SIAM Journal on Applied Mathematics, 70(3):789–802, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kimura:2009:ACI

[967] Toshikazu Kimura. American continuousinstallment options: Valuation and premium decomposition. SIAM Journal on Applied Mathematics, 70(3):803–824, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Baryshnikov:2009:TEE

[968] Yuliy Baryshnikov and Robert Ghrist. Target enumeration via Euler characteristic integrals. SIAM Journal on Applied Mathematics, 70(3):825–844, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tseluiko:2009:VEF

[969] D. Tseluiko, M. G. Blyth, D. T. Papageorgiou, and J.-M. Vanden-Broeck. Viscous electrified film flow over step topography. SIAM Journal on Applied Mathematics, 70(3):845–865, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sharpe:2009:TDI

[970] Gary J. Sharpe. Thermal-diffusive instability of premixed flames for a simple chain-branching chemistry model with finite activation energy. SIAM Journal on Applied Mathematics, 70 (3):866–884, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hameed:2009:BLJ

[971] Muhammad Hameed and Jeffrey F.

Morris. Breakup of a liquid jet containing solid particles: a singularity approach. SIAM Journal on Applied Mathematics, 70(3):885–900, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jones:2009:AAB

[972] G. W. Jones, S. J. Chapman, and D. J. Allwright. Asymptotic analysis of a buckling problem for an embedded spherical shell. SIAM Journal on Applied Mathematics, 70(3):901–922, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Antipov:2009:MYS

[973] Y. A. Antipov and A. Y. Zemlyanova.

Motion of a yawed supercavitating wedge beneath a free surface. SIAM Journal on Applied Mathematics, 70 (3):923–948, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Singer:2009:PNP

[974] A. Singer and J. Norbury. A Poisson– Nernst–Planck model for biological ion channels — an asymptotic analysis in a three-dimensional narrow funnel. SIAM Journal on Applied Mathematics, 70 (3):949–968, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kobayashi:2009:IED

[975] K. Kobayashi, Yu. Shestopalov, and Yu. Smirnov. Investigation of electromagnetic diffraction by a dielectric body in a waveguide using the method of volume singular integral equation. SIAM Journal on Applied Mathematics, 70(3):969–983, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Erban:2009:ASC

[976] Radek Erban, S. Jonathan Chapman, Ioannis G. Kevrekidis, and Tomáš Vejchodský. Analysis of a stochastic chemical system close to a SNIPER bifurcation of its mean-field model. SIAM Journal on Applied Mathematics, 70 (3):984–1016, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

McHugh:2009:PIW

[977] John McHugh. Progressive internal waves of permanent form. SIAM Journal on Applied Mathematics, 70 (4):1017–1031, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pitcher:2009:OST

[978] Ashley B. Pitcher. Optimal strategies for a two-runner model of middledistance running. SIAM Journal on Applied Mathematics, 70(4):1032–1046, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

delaHoz:2009:NSS

[979] F. de la Hoz, C. J. García-Cervera, and L. Vega. A numerical study of the selfsimilar solutions of the Schrödinger map. SIAM Journal on Applied Mathematics, 70(4):1047–1077, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Goldsztein:2009:CMT

[980] Guillermo H. Goldsztein. Clogging of multigraphs as toy models of filters. SIAM Journal on Applied Mathematics, 70(4):1078–1096, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Carasso:2009:FCF

[981] Alfred S. Carasso. False characteristic functions and other pathologies in variational blind deconvolution. A method of recovery. SIAM Journal on Applied Mathematics, 70(4):1097–1119, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ren:2009:OSD

[982] Xiaofeng Ren and Juncheng Wei. Oval shaped droplet solutions in the saturation process of some pattern formation problems. SIAM Journal on Applied Mathematics, 70(4):1120–1138, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Koroteev:2009:SBL

[983] M. V. Koroteev and I. I. Lipatov. Supersonic boundary layer in regions with small temperature perturbations on the wall. SIAM Journal on Applied Mathematics, 70(4):1139–1156, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bruining:2009:FCW

[984] J. Bruining, A. A. Mailybaev, and D. Marchesin. Filtration combustion in wet porous medium. SIAM Journal on Applied Mathematics, 70(4): 1157–1177, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Xia:2009:EHT

[985] Jing Xia, Zhihua Liu, Rong Yuan, and Shigui Ruan. The effects of harvesting and time delay on predator-prey systems with Holling Type II functional response. SIAM Journal on Applied Mathematics, 70(4):1178–1200, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Shanin:2009:WDP

[986] Andrey V. Shanin. Weinstein’s diffraction problem: Embedding formula and spectral equation in parabolic approximation. SIAM Journal on Applied Mathematics, 70(4):1201–1218, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wessel-Berg:2009:LSP

[987] Dag Wessel-Berg. On a linear stability problem related to underground CO_{2} storage. SIAM Journal on Applied Mathematics, 70(4):1219–1238, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ortner:2009:CRH

[988] Norbert Ortner and Peter Wagner. On conical refraction in hexagonal and cubic media. SIAM Journal on Applied Mathematics, 70(4):1239–1259, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Alali:2009:OLB

[989] Bacim Alali and Robert Lipton. Optimal lower bounds on local stress inside random media. SIAM Journal on Applied Mathematics, 70(4):1260–1282, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Arens:2009:MES

[990] Tilo Arens, Armin Lechleiter, and D. Russell Luke. MUSIC for extended scatterers as an instance of the factorization method. SIAM Journal on Applied Mathematics, 70(4):1283–1304, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rognes:2009:MMF

[991] Marie E. Rognes, M.-Carme Calderer, and Catherine A. Micek. Modelling of and mixed finite element methods for gels in biomedical applications. SIAM Journal on Applied Mathematics, 70 (4):1305–1329, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Elling:2009:ISR

[992] Volker Elling. Instability of strong regular reflection and counterexamples to the detachment criterion. SIAM Journal on Applied Mathematics, 70 (4):1330–1340, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Noethen:2009:QSS

[993] Lena Noethen and Sebastian Walcher. Quasi-steady state and nearly invariant sets. SIAM Journal on Applied Mathematics, 70(4):1341–1363, ???? 2009. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Babak:2009:EWP

[994] Petro Babak, Anne Bourlioux, and Thomas Hillen. The effect of wind on the propagation of an idealized forest fire. SIAM Journal on Applied Mathematics, 70(4):1364–1388, ???? 2009. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Velo:2010:ETM

[995] Ani P. Velo, George A. Gazonas, and Takanobu Ameya. Erratum: ztransform methods for the optimal design of one-dimensional layered elastic media. SIAM Journal on Applied Mathematics, 70(5):1736–1737, ???? 2010. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [965].